A Complete Bibliography of Publications in Annals of Statistics (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

14 October 2024 Version 1.10

Title word cross-reference

β [CHK^{+}24]. D [BMY20, HdBCAM21]. ℓ_{1} [Hua20, LS22b]. ℓ_{2,∞} [CLC^{+}21a]. ℓ_{p} [AFW22, WSB^{+}23]. F [PLD21]. k [CGZ22b, KKZ21]. L^{2} [Sta24]. M [GLN20, LY20, RW20b]. n [Ché24, SHSS^{+}21]. p [BCL^{+}23, Ché24]. S [LGRG23]. t [HL24]. U [BDH24, BKS21, HXWP21]. X [NCDK24].

-boosting [Sta24]. -estimation [LGRG23]. -estimator [GLN20]. -estimators [LY20, RW20b]. -means [KKZ21]. -model [CHK^{+}24]. -norma [LS22b]. -optimal [BMY20]. -penalized [Hua20]. -regularized [WSB^{+}23]. -statistic [BKS21]. -statistics [HL24, BDH24, HXWP21]. -tests [PLD21]. -values [BCL^{+}23].

2876 [BGvZ21]. 2D [NT24]. 2D-Navier [NT24].

40 [BGvZ21]. acceleration [SF24]. acceptance [STG21]. accuracies [JS22]. accuracy [Goo22, KR24, KRSW21, Zhi20]. across [HT23]. action [WQW23]. activation [Kut20, SH20a, SHV24]. acyclic [RERS23]. adaptable [LAP^{+}20]. Adaptation [FGKS21, AC24, CW22a, CW22b, PS22]. Adaptive [ALMM22, BBNW20, BNX22, CG21b, CCNT21, DW23a, GLCS20, HS21a, MLMR24, OL24, RCS21, SvZ20, AWS22, BR24, CW21, CW22a, CP22a, CKP24, DP21, FGS21b, LH23, MM21, PHL20, RCSN22, TSY20, WGS^{+}22]. Adaptively [BNR23]. adaptivity [CCZ24, JKL21]. addition [SM23]. Additive [JP20, JPV21, CP22a, GMW21, LPMP24]. Adjusted [ZA23, ZWH^{+}20, ZGV22]. admissibility [DLY22, Str21]. Admissible [MS20, VWW22, DR21a, GHS21]. adversarial [BJ23, HJ24, JS22, TY23]. advisor [vdGK21]. Affine [DP22]. Affine-equivariant [DP22]. after [Tia20]. alarm [Pol21]. Algorithm [ZCW22, CP22a, FGS21b, GZ22b]. algorithmic [ZJ23]. algorithms [BB21, BV20, BGL^{+}20, CKP24, DC23, Fan22, SWJ21, STG21, SL23]. alignment [DFZ24]. All-in-one [DM22]. Almost [Cox20]. alpha [DDP21]. alpha-divergence [DDP21]. alternatives [PS20]. AMP [CFM23]. analyses [CCF21]. Analysis [JMM20, NLS21, SWJ21, AFWZ20, BSS20, BW23, CZ21, CLP21a, CvdP21, Cho20, DKA20, Dob20, HJ24, JS22, LPYZ24, PV20b, RSH20, RW20b, SO23, SHSS^{+}21, SLY22, SWZ22, WSB^{+}23, ZC23, vDD24]. Analytical [LW20]. ancestral [Ger24]. Ann [BGvZ21]. anytime [POLdHG24]. anytime-valid [POLdHG24]. application [AR21, ALPW22, BCPW23, HYF23, KJJS20, KJJS22, LUZ21, LLM20, RY21, SG20, SMZ22]. applications [BZ21, BBNW20, BBSMW22, CBS20, CMW23, CP21a, CC23, CM22b, DZ23, EJ20, EB22, GSS22, HW21a, HLS22, LEV21, LWL21, LPYZ24, OL24, VW21, ZCW22]. approach [BDH21, CZ21, Che22, DP21, DM23, DW23b, GKM21, HdBCAM21, PV20b, WWZ^{+}24]. approaches [NCDK24]. Approximate [KYZ21, SS22, BJ22, JSLH23, MV21, Fan22]. approximation [Che21, DDJ24, DZ20a, Goo22, KR24, Kol22, Lop22]. approximations [AR20, CCKK22, DZ23]. arbitrary [VWW22]. Archimax [CFN20]. ARFIMA [HCCI22]. armed [CP22a, CCL24, Cha20]. array [CT22]. array-based [CT22]. arrays [DDG21, HLS22, PWLL21, SM23]. aspect [MW24]. Assessment [KS20, CFNP21]. asset [BCPW23]. Assigning [KPST23]. assimilation [NT24]. assisted [BGL^{+}20, CJN24, DQW20, Tan20]. associated [LZ22a]. association [HLM23]. assumption [BNR23, FGRS22]. assumptions [BG21, HJ20, vDD24]. asymmetrically [CCF21]. Asymmetry [CCF21]. Asymptotic [BB21, Bha20, CVFL22, DDJ24, DR21b, GFLS21, Goo22, Hua20, HS21b, KJJS20, LHY20, QLY23, RS20b, SW07, SWZ22, ZZPZ22, KLR^{+}20, KZ21, LEV21, LS22b, SWFD21, KJJS22, ZSW22]. Asymptotically [HXWP21, FY23, KB22, Pol21]. Asymptotics [CM21, DZ24, HSS22, Han23, Li22, LW22, LSW21, YZZ22]. Augmented [HW21b]. autocovariance [BS23a]. autoregression [AF23]. autoregressions [CM21]. autoregressive [GKM21, WT23, DZ23]. auxiliary [CCZ22]. Average [SAH21, LL22, LQW^{+}22, PMM^{+}20, YNB^{+}22]. average-case [PMM^{+}20]. Averages [KLR^{+}20]. averaging [CL24, DS21, KBW20].

Backfitting [GHO22]. backward [DC23, LSV20]. balancing [WQW23]. balls [HJWW20, MM21]. Banach [DKA20]. bandable [LR20b]. bandit [Cha20]. bandits [CP22a, CCL24]. bands [DW23b, PLD21]. barriers [CM22a, SW22]. based [ALMM22, Bha20, BDH21, BP24, CFF20, Cha22b, CT22, DDP21, Dob22, GKM21, LEV21, LY20, LS22a, LSC^{+}24, MPW23, OQC22, PSD24, RGvdL22, ZD22, ZZZ24, ZXZL24, ZZYS20, ZS20]. basis [BCCP23]. Batch [LQW^{+}22]. batching [HL24]. Bayes [ADXR20, BB20, BG20, CR20, DR21a, GvdVZ20, HKSW21, MS20, MJ22, OL24]. Bayes-like [BB20]. Bayesian [AC24, AD22, BSS20, CvdP21, CR21, CLP21b, DRR20, GKM21, GR22, Kle21, Li22, MNP21, PR22, RvdV20, RSH20, RvdP20, RS20b, RS24, WG23, ZC23]. Beginning [EG21]. behavior [KLR^{+}20]. benefits [KM21a]. Berry [Zhi20]. best [WWM20]. Betti [RKP23]. between [BB21, CGK22, DDB20, HLM23, ZXZL24]. Beyond [AL21, DZ20a, PS20, BCRT23, CC23]. Bias [KZ21, DSH23]. bias-variance [DSH23]. binary [CJLZ24, FNZ22, LUZ21]. binomial [SHSS^{+}21]. biomarker [HS22]. bit [DMR22]. bivariate [Efr22, MPW20]. block [BZ23b, HZ20, KW22, XJL20, ZLN22, ZA23, ZVB21, SPY20]. blocks [ZVB21]. blockwise [YKN24, ZZZ24]. Board [Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano24a, Ano24b, Ano24c, Ano24d]. Boosted [LCI21]. boosting [LS22b, Sta24]. Bootstrap [CJN24, DZ20a, Hid21, Kol22, LY20, CZ20, CCKK22, HL20, IM21, Lop22, MPK20, ZP22, Zhi20]. Bootstrap-assisted [CJN24]. Bootstrapping [LLM20, RKP23, BDH24]. boundary [ACR24, EB22, RSH20, SM20]. bounded [CG21b]. Bounds [YNB^{+}22, BNR20, DSH23, GLCS20, JSLH23, KR20, OQC22, OvdG21, RW20a]. breaker [LO23]. Breaking [MWW23]. breaks [LLS24]. Bregman [BFL21, SWJ21]. bridge [WWM20]. Bridging [CFMY21, DDB20, FMM23]. broad [CT22]. bulk [FSW21].

calculations [BBSMW22]. calibrated [Tan20]. Calibration [VW21, FL22]. can [LR20a]. Canonical [AV23, SF22, SS20]. Carlo [AL21, DMS20, FT23, HBDD20, KJJS20, KJJS22, MMS23]. CART [CR21, CC23]. Carving [Pan23]. case [PMM^{+}20]. Causal [GMPE21, BFPM21, Ger24, IM21, KS20, KBG20, KBRW24, RvdV20, ZD22]. censored [DQW20, Efr22, HPTZ22, TYZ20]. censoring [Efr21]. Central [LWL21, Lop22, CCKK22, LPYZ24]. chain [BS23a]. chains [ALPW22, DDB20, Kol22, SPY20]. Change [SF24, XWZY24, CNYY21, DM20, FLS20, HKTY24, HKW21, LGS21, VFLRB23, WYR21, WZVS22, ZWL20]. Change-point [XWZY24, CNYY21, DM20, FLS20, HKTY24, VFLRB23]. change-points [ZWL20]. changing [Dre23]. characteristic [TCPS23]. Characterization [Ger24, KLR^{+}20, LO23]. Characterizing [BKRS23]. Charles [EG21, Str21]. charts [Pol21]. chi [ZA23]. chi-square [ZA23]. choice [MBR21]. Cholesky [LR20b]. Choosing [CGK22]. class [BSS20, CT22, GX20, WZL22]. classes [ZC23]. Classification [KRSW21, CW21, CBS20, HS22, JS22, SBC24]. classifier [CW21]. classifiers [LS22b]. classifying [KBG20]. closed [GHS21]. closure [FRB24]. CLT [LHBS23, SMZ22]. cluster [Leu22]. cluster-randomized [Leu22]. Clustering [SPY20, AFW22, AR21, ADXR20, BBNW20, BFL21, BGL^{+}20, GM22, KKZ21, LZZ21, LZ22b]. co [BJ22]. co-sufficient [BJ22]. coefficients [GW22]. Cointegration [BG22]. columns [MWY23]. combination [HS22, VW21]. Combinatorial [ZL24, MW22, WWZ^{+}24]. communication [DH24, SvZ20, SVvZ23]. Community [JLLX21, YQ21, CLM22, PC20, XJL20, YLFS22, ZZ20]. compact [KB22]. compare [DK22]. Comparison [ALPW22]. comparisons [CGZ22a, MW22, PMM^{+}20]. compatibility [BS23c]. complementary [GW22]. Completely [BS23c]. completion [WP22, XYZ21]. Complex [HW21a]. Complexity [ZC23, DR24, DH24, LSC^{+}24, MMS23, MN20, YZZ22, ZGMS24]. compliers [KBG20]. component [Nda22, PRV20, SO23]. components [BDWW22, DY21, FSW21, KLN20, KM21b, MK21a]. compound [RSH20]. computation [OL24, PS22]. Computational [HMR23, SW22, SS22, CW20, DH24, HWZ22, LZ22b, LX23, ZZ20]. computationally [XYZ21]. concave [FGKS21, HS21b]. Concentration [AR20, AR21, SS20, PV20a, RvdP20]. Concordance [SSL21]. Conditional [FT23, FL22, SL23, AC21, BJ22, BFR23, CPH23, GSUC21, HHPS20, KNBW22, LSV20, LS22a, NBW21, NCDK24, SP20, YNB^{+}22]. conditioning [HJ20]. conditions [KB22, Lah21, LY20, RCSN22]. Confidence [DHZ21, EB22, BPS20, BNR20, DW23b, FLS20, HRMS21, LY20, PLD21]. Conformal [BCRT23, BCL^{+}23]. confounders [Ger24]. confounding [GĆB22, YNB^{+}22]. conjunctions [KS20]. connected [KL21]. connections [BS23c]. connectivity [YQ21]. consequences [KMW20]. Consistency [Dob22, GS22, PR22, AY23, GKM21, HKSW21, JT21, Lah21, NT24, SS20]. Consistent [BCCG21, EJ20, HCCI22, Nic24, ZWL20, DHS20, LSWMC24, PC20, SHDH22]. constancy [SWFD21]. constant [DDB20, GHZ20]. constrained [BS23a, HS22, RW20b, dÁLM21]. constraints [HSS22, Han23, PTW21, RS20a, SvZ20, SVvZ23, XS21]. Construction [PWLL21, HLS22]. Contents [Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano24e, Ano24f, Ano24g, Ano24h]. contextual [CCL24]. Continuous [RGvdL22, DKA20, DWW^{+}24, SO23]. Continuous-time [RGvdL22, SO23]. continuum [CP22a]. continuum-armed [CP22a]. contributions [FS21, Str21]. control [CCZ22, FLS20, FL22, MM21, Pol21, RV22, ZL24]. Controlled [HBDD20].

Controlling [SXF24, GHS21]. Convergence [CCK24, LPYZ24, WN22, XYY20, ZG20, ZW21, BB21, BE22, CWZ21, CW22b, CFM23, CPT23, CM22b, DHK^{+}20, HS21b, KL21, RS20a, RS24]. convex [BZ23a, CZ21, CCZ24, CM22a, CFMY21, HSS22, Han23, LZ22a, PTW21]. convexity [CFM23, Cox20, JMM20]. copula [GKV^{+}22, ZGV22]. copulas [BP24, CFN20, ES21]. corrected [ZA23]. Correction [BGvZ21, FZ21, LBvdL21, SHV24, ZSW22]. correlated [DD23, DGR23]. correlation [GFLS21, HY22, LWL21, XY20]. correlation-free [XY20]. correlations [DHS20]. corroboration [KS20]. cost [CWZ21, CW22b]. count [CL22]. Coupled [LSV20]. coupling [JLL21]. Covariance [DMR22, AMR20, BD20, BDWW22, BSS20, CHP20, CP22b, CQZ23, Ché24, DY21, DZ24, DKH21, GLN20, HW20, LW20, LHY20, Li22, LHBS23, LSW21, LGRG23, MZ20, Mou22, QLY23, RW23, SP20, WP22, XYY20, ZW21, ZZPZ22]. covariate [KM21a, LH23, MPW23, WL22]. covariate-adaptive [LH23]. covariate-shift [KM21a]. covariates [LCI21, LH23, WL22]. Coverage [CG21a, WG23, HL24, RS20b]. Cox [ZMW22]. Cramér [GSS22]. Cramér-type [GSS22]. credible [CG21a, RS20b, WG23]. criteria [CT22, SSL21]. criterion [BDSW22]. criticism [DK22]. cross [CC23, CLC21b, FJLX22]. cross-sectional [FJLX22]. cross-validated [CLC21b]. cross-validation [CC23]. crossed [GHO22]. CrowdSourcing [PCV23]. cryo [FLS^{+}24]. cryo-EM [FLS^{+}24]. Cube [BE22]. cure [BEPV22, PV20b]. current [Efr21]. curse [HJ24, MWW23]. curved [SS20]. curves [PLD21, TCPS23]. cycles [BFPM21].

DAG [ZC23]. data [CLC^{+}21a, CP21a, CPT23, CFMY21, CP21b, Che22, CL22, DSUC23, DKA20, DQW20, DM23, ES21, EH23, FWZ21, FNL20, FJLX22, HS21a, KL20, LS22a, LM21a, LLM20, MWY23, MP23, MK21b, NT24, RW23, RCSN22, SBC24, SLY22, Tan20, TYZ20, WZVS22, WZL22, YCF24, YSK21, ZWH^{+}20, ZZZ24, dRdCMP21]. data-adaptive [RCSN22]. dataset [QTD24]. datasets [Cha22a]. debias [XZZ22]. Debiased [YSK21, GĆB22]. Debiasing [BZ23a, ZP22]. decay [LLM20]. decision [AO22a, CG21b, HW23, LQW^{+}22, SSL21, YZZ22]. decomposition [JLLX21]. Deconvolution [GLL22, BG21, RS24]. Deep [CJQY24, JSLH23, ZMW22, FGZ24, KL21, Kut20, SH20a, SHV24]. degree [SW22, ZA23]. degree-corrected [ZA23]. delayed [STG21]. delayed-acceptance [STG21]. denoised [WY20]. denoising [BDW21, CC23, FGS21a, GLCS20, SG20]. dense [XM23]. densities [KK20, NWB22, SG20]. Density [BG21, BE22, BFR23, Efr22, FGKS21, HT23, MJ22, PLD21, RCSN22, XS21, ZW21]. dependence [AC21, CLP21b, Dre23, EFdH^{+}22, FL22, LEV21, SWFD21, Vel22, VWW22, XWZY24, ZZYS20]. dependencies [BDH24]. dependent [CLP21a, DSUC23, GSS22, Ing20, LCI21, Pol21, YQ21, ZLN22, dRdCMP21]. derivatives [LL22]. descent [CLTZ20, DDP21, DDB20, ZS22]. Design [DLY22, HLS22, LFR21, PMM^{+}20, SGWH20, ZD22]. design-based [ZD22]. Designs [JPV20, BMY20, CMW23, CT22, HW21a, KYZ21, Leu22, LO23, TX21, WM21]. Detecting [WGS^{+}22]. Detection [LLS24, BM22, Bha20, CW20, CLM22, DM20, EKJ20, HKTY24, JLLX21, LGS21, MLMR24, PC20, SF24, VFLRB23, WYR21, YQ21, ZZ20]. deviation [ERV20, GSS22]. deviations [BDH21, vDD24]. difference [Cha22b, SMZ22]. difference-based [Cha22b]. differences [DK22]. differentiable [LC24]. differential [CWZ21, CGC20, LBY23, PSD24, RS20a]. Differentially [AMBL23, BRS23, CHK^{+}24, LTMC24]. diffusions [AWS22, BR24, GR22, NR20, Nic24]. Dimension [LS22a, CJQY24, DQW20, FGRS22, FNZ22, GMMM21, HdBCAM21, HS21a, KLYL20, LL22, ZZYS20]. dimensional [BDP23, BZ21, BCPW23, BV20, BW23, BFZ22, BJZ24, CZ21, CS20, CM22a, CP21a, CQZ23, CVFL22, DDP21, DZ20b, DQW20, DM23, DZ24, DWYZ23, DFZ24, DHS20, FWZ21, FLY23, FNL20, FJLX22, GZ22a, GFLS21, GW22, GLP20, GMW21, GĆB22, HW20, HSS22, HMRT22, HXWP21, Ing20, JZLM23, JO20, Kol22, LW20, LPMP24, LWL21, LLS24, LS22b, LGS21, LHBS23, NR20, SF22, Sta24, SVvZ23, Tan20, WS20, WZVS22, WT23, XS21, XWZY24, XY20, ZW21, ZC23, ZXZL24]. dimensions [BDH24, BP24, CCKK22, CLC21b, FT23, Han21, Han23, HS23, Lah21, LLHL21, LLM20, Lop22, MBR21, ZZZ24]. diminishing [WL22]. directed [RERS23]. directions [PRV20]. disclosure [CFNP21]. discovery [CCZ22, FL22, GLP20, GMPE21, GHS21, KR20, MLMR24, RV22, SXF24, ZL24]. discrete [GZ22b, HMR23, MBR21, SO23]. discrete-time [SO23]. discriminant [BW23, CZ21]. discriminators [MBR21]. Discussion [Kut20]. disjoint [BZ23b]. Dispersal [HT23]. displacement [JMM20]. Distance [ZZYS20, BBSMW22, CLP21b, DCM24, ERV20, GFLS21, NWB22]. Distance-based [ZZYS20]. distances [ZGMS24]. Distributed [CW22a, CW22b, CP21b, CJLZ24, DS21, LTMC24, SvZ20, SVvZ23]. Distribution [HdBCAM21, XY20, BDW21, BG21, BB21, BSS20, Bha20, DW23b, Efr21, GLL22, HZ20, LHY20, MM21, Mur21, RV22, SHDH22, TY23, XYY20]. distribution-free [SHDH22]. distributionally [DN21]. distributions [AO22b, AV23, BBSMW22, GFLS21, HY22, HKW21, KKZ21, Lei21, MS20, PR22, PHL20, ZG20]. divergence [DDP21]. divergent [CHP20]. diverging [LHBS23, MW24, TYZ20, WL22]. do [FGZ24]. document [BBSMW22]. documents [KPST23]. domain [AvD20, Hid21, Li22, LSW21, MPK20, RY21, YKN24]. Donoho [BKRS23]. Double [DQW20]. Double-slicing [DQW20]. Doubly [GĆB22, DWW^{+}24, NCDK24]. drift [AWS22, CGC20, NR20]. driven [CHHW21, LLSS21]. duality [ZGMS24]. dyadic [CC23]. dynamic [BJZ24, DZ24, LSWMC24, OQC22, WYR21]. dynamical [MN20].

E-statistics [POLdHG24]. E-values [VW21]. Early [Sta24]. Edge [CHK^{+}24, SXF24]. edges [AMR20]. Edgeworth [HL24, KK20, ZX22, Zhi22].

Edgeworth-type [Zhi22]. Editorial [Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano24a, Ano24b, Ano24c, Ano24d]. effect [CDRV21, DWW^{+}24, JPV20, KBRW24, LW22, MH22, SM23, YNB^{+}22]. effects [BJZ24, GHO22, KBG20, KYZ21, SAH21, Tan20]. Efficiency [FY23, STG21, KZ21]. Efficient [BS23a, BS23b, HLM23, KLN20, LPMP24, QTD24, vdVW21, Cha20, XYZ21]. Eigenvalue [CCF21, CHP20, QLY23]. eigenvalues [CHP20, HY22, LHY20, ZZPZ22]. eigenvector [AFWZ20, CCF21, XYY20]. elaborate [KS20]. elbows [BRS23]. elicitability [FZ16, FZ21]. ellipsoids [ZS22]. Empirical [BG20, DDG21, ES21, MN20, BE22, BFR23, BEV20, BP24, CR20, CLvdG22, Han21, TYZ20, XYY20, YKN24]. energy [CFM23]. enrichment [LFR21]. entire [FGS21a]. entries [XYZ21]. Entropic [ZGMS24, BB21, BNR23]. entropy [BFR23, HJWW20]. Entrywise [AFWZ20]. enumerator [TX21]. Envelope [ZS20]. Envelope-based [ZS20]. equations [CGC20, NT24, PSD24, TYZ20]. equivalence [GM22, MH20, ZC23]. equivariant [DP22, MS20]. ergodic [BR24]. ergodicity [DMS20]. Erratum [Ché24, KJJS22]. error [GZ22a, JZLM23, JSLH23, OQC22, RW20a, Tia20]. errors [BG21, HL24, JPV22, KP22]. Esseen [Zhi20]. estimate [CS20]. estimates [KL21, MJ22, ZW21]. Estimating [CDRV21, GZ22a, MK21a, BGvZ12, BGvZ21, JPV20, RCSN22, TYZ20]. Estimation [CCZ24, DZ20c, DK21, KZ21, Kol22, KM21b, MP23, MV21, PCV23, TCPS23, Vel22, ABL23, AWS22, ATW24, BCCP23, BDP23, BB20, BZ23a, BS23b, BDM22, BFR23, BBNW20, BNX22, BBSMW22, BCCG21, BRS23, CWZ21, CLC^{+}21a, CW22a, CW22b, CKP24, CNYY21, CHK^{+}24, CG21b, CJLZ24, CL24, Ché24, CL22, CCNT21, CGC20, DL22, DZ24, DMR22, DRR20, DWYZ23, DFZ24, DR24, Efr21, Efr22, EJ20, FLS^{+}24, FLS20, FGKS21, FNZ22, GHZ20, GM20, GSUC21, GLN20, Goo22, HJWW20, HWZ22, HPTZ22, HW21b, HT23, Hop20, HLM23, HR21, KBRW24, KLN20, KP22, LN22, LEV21, LW22, LLS24, LR20b, LSW21, LGRG23, LC24, LM21b, LX23, MBNWW24, MPW20, MS20, MZ20, MW24, MK21b, NWB22, OW21, PMM^{+}20, PSD24, QTD24, RY21, RGvdL22, SW22, SH20b, SGWH20, Tan20, TSY20, TY23, TV24, WT23]. estimation [WY20, XJL20, XS21, vdVW21]. estimator [DM22, GLN20, Hua20, MWW23, SG20, WWM20, YSK21, ZS22]. estimators [BZ23a, BE22, BDH21, CFF20, CJN24, Cha22b, FY23, HZ20, HS23, HS21b, LY20, MMS23, RW20b, SMZ22, ZZZ24]. Euclidean [CP21a, JPV21]. Euler [TCPS23]. evaluation [HW23]. events [EFdH^{+}22]. Evidence [ZLSK22, KS20]. Exact [BV20, DGR23, Han23, Mou22, KYZ21]. exchangeability [BCRT23]. exchangeable [DDG21, WN22]. Existence [DKH21, CS20]. expansion [BCCP23, KK20]. expansions [HL24, ZX22, Zhi22]. expected [AFWZ20, TCPS23]. expectile [GSUC21]. experiments [DLY22, LH23, ZD22]. Exploration [DCM24]. exponential [KMW20, LUZ21, LZ22a, SS20]. exponential-family [SS20]. extended [DR21a, HS21b]. Extending [MPK20]. extension [BEPV22]. extensions [FGS21a]. extent [KS20]. extraction [CP21a]. extrema [CM22b]. extremal [DSUC23]. Extreme [EH23, GSUC21, Dre23, EFdH^{+}22, LHY20]. extremes [LEV21, REZ23, ZVB21].

Factor [LLSS21, BBNW20, BW23, CL24, Dob20, FMM23, HKTY24, LR20b]. Factor-driven [LLSS21]. factorial [TX21]. factorization [PC20]. factors [KS20, YNB^{+}22, ZLSK22]. fair [AO22a]. fairness [CS22]. False [FLS20, RV22, BNR20, CCZ22, FL22, GHS21, KR20, MLMR24, Pol21, SXF24, ZL24]. families [BNR20, KMW20, LUZ21, LZ22a]. family [SS20]. far [AD22]. feasible [Cha22a]. feature [CP21a]. features [HJ24]. field [ZZ20]. fields [KB22, LSW21, NR20, TCPS23]. filling [CT22]. filter [LSV20]. filtering [CC23, GLCS20, WGS^{+}22]. filters [CSSV22]. Finding [LSWMC24]. Finetti [WN22]. Finite [MMS23, AD22, KM21b, MK21a, Zhi22]. Finite-sample [MMS23]. Fisher [YWR24, DR24, LSWMC24]. fit [BJ22]. fixed [Li22, LSW21]. fixed-domain [Li22, LSW21]. flat [BATU23]. flats [WM21]. flexible [CCK24, HLS22]. flow [YWR24]. forests [CVFL22, MGS20, RvdP20]. formal [KK20]. forms [QTD24]. Foundations [BFPM21]. fourth [BDM22]. fourth-moment [BDM22]. fractional [DK21, MP23, TX21]. Frame [dÁLM21]. Frame-constrained [dÁLM21]. framework [Cha22b, CC23, CS22, GvdVZ20, HWZ22, PHL20, vDD24]. Fréchet [BM23, CM22b, DM20, LL22, LM21a, MH22, PLD21]. free [CFM23, FLY23, HK22, KBB^{+}20, Pan23, SHDH22, XY20]. frequency [AvD20, Cho20, DK22, Hid21, MPK20, MP23, Nic24, RY21, RW23, SWZ22, YKN24]. Frequentist [Kle21, RS20b, RS24]. full [CGZ22a]. fully [KL21, SHDH22]. function [BCCP23, BGvZ12, BGvZ21, CW22b, CCK24, Kut20, SH20a, SHV24, SH20b, vDD24]. function-valued [vDD24]. Functional [DKA20, LL22, AvD20, BS23b, BRS23, CKP24, CM21, ES21, KL20, LPMP24, LS22a, LLS24, LLM20, SLY22]. functionals [KLN20, KZ21, Kol22]. functions [AMR20, CCZ24, CG21b, DKA20, DW23b, GH21, HdBCAM21, LSW21, WC20]. Fundamental [CM22a, EKJ20, MW24]. Fuse [MK21a]. fusion [WWZ^{+}24].

GANS [BCST20]. gap [DDB20]. Garza [DLY22]. Gaussian [BATU23, CW22a, CMW23, DM22, DZ20a, DL22, DWYZ23, FGRS22, GIKM22, HSS22, Hop20, JT21, KR24, Kol22, Li22, LZZ21, LSW21, LX23, MNP21, Mur21, Nda22, NT24, RY21, SG20, WY20, YWR24, ZLN22]. Gaussian-subordinated [ZLN22]. genealogies [KJJS20, KJJS22]. General [Cha22a, BG21, CMW23, Cha22b, FSW21, GvdVZ20, GSS22, Han21, LO23, MWY23, MPK20, PV20b, QLY23, QTD24, TYZ20, vDD24]. generalised [SP20]. Generalization [OQC22, HJ24, LAP^{+}20, MZ22]. generalize [LR20a]. Generalized [ZJS22, ALS23, CJN24, FGS21b, HWZ22, LHY20, SWJ21, TYZ20]. generalizing [KBG20]. genetic [ZZZ24]. Geometric [KLR^{+}20, Bha20, CP21a, DMS20]. Geometrizing [RS20a]. geometry [Eva20]. Gibbs [BCCP23, SM20]. Global [CLM22, CPT23, Cox20, Han21]. go [NLS21]. goodness [BJ22]. goodness-of-fit [BJ22]. gradient [CLTZ20, DDP21, DDB20, YWR24, ZS22]. gradient-based [DDP21]. gradually [BDH21]. gram [LPYZ24]. graph [Lei21, LW22]. Graphical [BR23, Che22, LZ22a]. graphons [ALS23]. graphs [AMR20, Bha20, BCCG21, Che22, DZ20b, DD23, Ger24, GGCvL20, MH20, RERS23, SS20, XM23]. greater [Ché24]. Grenander [CJN24]. Grenander-type [CJN24]. GRID [GLP20]. Gromov [ZGMS24]. group [BMNT23, FLS^{+}24, POLdHG24, MK21a]. Grouped [HMR23]. Groups [MW22, AO22b, JPV22]. growing [MBR21]. guarantee [MLMR24]. guarantees [CLC^{+}21a, CPT23, MNP21, ZZ20, Zhi22].

Half [BDSW22]. Half-trek [BDSW22]. Hamiltonian [DMS20]. hardness [SP20]. Hardy [FGS21a]. Hawkes [CL22, DRR20]. hazards [CLP21a, LCI21]. HC [PS20]. Heavy [AC24, BCCG21, DSUC23, FWZ21, GSUC21, GMPE21, HY22, KP22]. Heavy-tailed [AC24, BCCG21, DSUC23, FWZ21, GSUC21, GMPE21, HY22, KP22]. helps [CW22a, CCF21]. heterogeneous [EH23, KBRW24]. heteroscedastic [GSUC21, KP22]. Heteroskedastic [ZCW22, YCF24]. hidden [BNX22, GĆB22]. hierarchically [CLP21a]. hierarchy [AO22a]. High [BCPW23, BJZ24, Cho20, DZ24, DHS20, FWZ21, GĆB22, HSS22, XS21, BDP23, BDH24, BZ21, BV20, BW23, BFZ22, BP24, CZ21, CS20, CM22a, CP21a, CQZ23, CCKK22, CLC21b, CVFL22, DQW20, DM23, DWYZ23, DFZ24, FLY23, FLS^{+}24, FNL20, FGRS22, FJLX22, FNZ22, FT23, GZ22a, GFLS21, GW22, GMMM21, GLP20, GMW21, HW20, Han23, HS23, HMRT22, HXWP21, Ing20, JZLM23, JO20, KR20, Kol22, Lah21, LPMP24, LLS24, LS22b, LLHL21, LGS21, LLM20, Lop22, MP23, PV20a, PWLL21, RW23, SF22, Sta24, SVvZ23, Tan20, WS20, WZVS22, WT23, XWZY24, XY20, ZW21, ZZZ24, ZC23, ZXZL24, ZZYS20]. High-dimensional [BCPW23, BJZ24, DZ24, DHS20, FWZ21, GĆB22, HSS22, XS21, BDP23, BZ21, BV20, BW23, BFZ22, CZ21, CS20, CM22a, CP21a, CQZ23, CVFL22, DQW20, DM23, DWYZ23, DFZ24, FLY23, FNL20, FJLX22, GZ22a, GFLS21, GW22, GMW21, HMRT22, HXWP21, Ing20, JZLM23, JO20, Kol22, LPMP24, LLS24, LS22b, LGS21, SF22, Sta24, SVvZ23, Tan20, WS20, WZVS22, WT23, XWZY24, XY20, ZW21, ZC23, ZXZL24]. High-frequency [Cho20, MP23, RW23]. high-noise [FLS^{+}24]. high-probability [KR20]. Higher [DK22, FZ16, HL24, OvdG21, FZ21]. Higher-order [HL24]. Hilbert [LC24]. Hilbert-valued [LC24]. Hilbertian [JP20]. hoc [BNR20]. homogeneity [WWZ^{+}24]. Hotelling [LAP^{+}20]. HSIC [ALMM22]. Hunt [EG21]. Hurst [SH20b]. hypergraphs [YLFS22]. hypersphere [KP23]. Hypothesis [JZLM23, WS20, AV23, CMW23, FZP20, GGCvL20, KS20].

i.i.d [BNR23, CGC20]. Identifiability [ADXR20, BDSW22, GX20, PRV20]. ignorability [HW23]. image [SM20]. imbalance [WL22]. impact [FGZ24]. impacts [LH23]. importance [DP21]. Improved [CCKK22, Ché24]. imputed [Cha22a]. incomplete [CLC^{+}21a]. independence [ALMM22, BJ22, BKS21, BP24, CPH23, DHS20, FJLX22, KNBW22, LEV21, MH20, NBW21, NCDK24, SP20, SHDH22, ZZPZ22, ZXZL24]. independent [DZ20a, HXWP21, MWY23, Pol21, SO23]. index [BM23]. indices [ZXZL24]. individualized [CLM22, FNZ22]. inequalities [ALPW22, Zhi20]. Inference [CFN20, CHLZ23, DSUC23, HHPS20, LFR21, PV20a, RW23, WZVS22, XZZ22, XM23, YCF24, AMBL23, BDWW22, BCS20, BCRT21, BZ21, BS23a, BJZ24, BEV20, CCZ24, CJN24, CZ20, CLTZ20, CP21b, CJLZ24, CHHW21, CHL^{+}24, DZ20c, DZ23, Dre23, DCM24, DK21, DP22, EH23, GFLS21, GH21, GR22, GMW21, GĆB22, HYF23, HPTZ22, KR24, KLYL20, KW22, KBB^{+}20, LPYZ24, LBvdL21, NR20, Nic24, PR22, PTW21, Pan23, RvdV20, SWZ22, SL23, SM20, Tan20, WC20, XWZY24, YKN24, ZZ20, ZWH^{+}20, ZJ23, ZF24]. inferences [ZD22]. Infinite [DDP21]. Infinite-dimensional [DDP21]. infinity [AD22]. information [CCZ22, DR24, EB22, SSL21]. informative [CSSV22]. inhomogeneous [CLM22, GGCvL20]. instrumental [ZLSK22]. instruments [KBG20]. integral [DM23]. integrated [GKV^{+}22]. Integrative [PTW21]. intensity [SMZ22]. interacting [KJJS20, KJJS22]. interaction [BCCP23]. Interactive [BRS23, CW22a]. interference [JPV20, Leu22, LW22, SAH21]. interpolate [LR20a]. interpolated [LS22b]. interpolation [HMRT22, MZ22]. interpolators [CLvdG22]. interval [BZ23a]. intervals [BPS20, CG21a, DHZ21, WG23]. intervention [RGvdL22]. intervention-specific [RGvdL22]. Intrinsic [SLY22, HS21a]. invalid [ZLSK22]. invariance [Dob22, EG21, POLdHG24]. invariance-based [Dob22]. invariant [AO22b, Fan22, GIKM22]. inverse [Han23, LLHL21, MWW23, MNP21, YSK21]. inverse-variance [YSK21]. Irreducibility [DMS20]. Ising [BK20, GM20, XM23]. Isotonic [DZ20b, PS22, DHZ21, FGS21a, GHZ20, HZ20, MPW20, WG23]. Isotropic [AMR20, LSW21]. Iterative [GZ22b, CPT23].

jackknife [BCRT21]. jittering [CHK^{+}24]. Joint [GM20, LHY20]. Just [LR20a].

Kendall [LWL21]. Kernel [LR20a, AR21, HSL24, LPMP24, SS22]. kernels [WP22]. Kesten [ALS23]. knockoffs [BCS20, HJ20, SJ22, WSB^{+}23]. Krause [FGS21a]. Kronecker [DKH21].

labels [KM21a]. LAGO [NLS21]. Large [CFF20, KW22, BG22, CCZ22, DK22, GHO22, HTYZ20, HYF23, LW20, LHY20, LWL21, LR20b, LHBS23, ZZPZ22]. large-dimensional [LW20, LHBS23]. Large-scale [KW22, HTYZ20]. largest [CHP20]. Lasso [CMW23, GĆB22, CHHW21, KFG24, Lah21, CLC21b, LPMP24, MM21]. LASSO-driven [CHHW21]. latent [BDSW22, BCPW23, BW23, BFPM21, Ger24, GX20, HKTY24, SS23, WN22, WZL22]. latent-class [WZL22]. law [EH23]. laws [CHP20]. layer [JMM20, PC20]. layers [GMMM21]. lazy [MZ22]. learn [NLS21]. learn-as-you-go [NLS21]. learnable [SS23]. Learning [ALS23, BK20, DM23, DN21, MW22, YWR24, ABL23, CW21, CCL24, CKP24, CBS20, DW23a, FS23, HS21a, HK22, LL20, LSC^{+}24, LQW^{+}22, NIS^{+}21, OQC22, RCS21, SWJ21, WQW23, ZMW22, ZC23]. least [HMRT22, KP22, Mou22, OvdG21, ZS20]. lengths [Pol21]. level [BE22, RCSN22, WM21]. libraries [CCK24]. Lie [JPV22]. lifetimes [Efr22]. like [BB20]. Likelihood [BM22, BBSMW22, BEV20, BMNT23, CS20, EJ20, FLS^{+}24, GH21, GKM21, Goo22, HSS22, RY21, SG20, TYZ20, YKN24]. Limit [AO22b, HZ20, KMW20, BATU23, BKRS23, CCKK22, HW21a, LWL21, LPYZ24, Lop22, LX23]. Limiting [CHP20, HY22]. limits [CW20, EKJ20, Kle21, LZ22b, MW24]. Linear [HS22, ATW24, BM22, BZ23a, BFZ22, DL22, DS21, FSW21, FGS21b, GZ22a, GvdVZ20, GW22, Han23, HW21b, HS21b, Ing20, KB22, KLN20, KBB^{+}20, LWL21, LGRG23, Mou22, MBR21, SF22, Sta24, ZZPZ22, ZMW22]. Linearized [GMMM21]. linearly [DLS20]. Lipschitz [HJWW20]. loadings [CL24]. Local [BDP23, CBS20, CFM23, HTYZ20, KNBW22, CM22b, CPH23, Eva20, LBY23, MH20, RS20a, SXF24]. localization [VFLRB23, WYR21]. Locally [LZ22a, ZF24, BRS23, DR24, FY23, ZW21]. locating [CM22b]. location [BGvZ12, BGvZ21, DWYZ23, DP22, MS20, PV20a, SG20]. log [FGKS21]. log-concave [FGKS21]. logistic [BMY20, CS20, HTYZ20]. long [BDP23, Hid21, ZLN22]. long-range [ZLN22]. long-run [BDP23]. longitudinal [FNL20, SLY22, WZL22]. loss [LSWMC24, RGvdL22]. loss-based [RGvdL22]. losses [TY23]. low [AFWZ20, CCF21, CHLZ23, DM23, FWZ21, HS21a, LX23, MV21, MW24, Nic24, SW22, XYZ21, XZZ22]. low-degree [SW22]. low-dimensional [DM23]. low-rank [CCF21, CHLZ23, FWZ21, LX23, MV21, MW24, XZZ22]. lower [DSH23, Mou22]. LSS [LHBS23]. lunch [HK22].

machine [LL20]. machines [HS21a]. Majority [LTMC24]. make [BK20]. manifolds [JSLH23]. many [Pol21]. maps [HR21, MBNWW24]. margin [SXF24]. Marginal [KM21a, ALPW22]. marginalized [MH20]. Markov [SPY20, ALPW22, BS23a, DDB20, FZP20, HW23, LQW^{+}22, MH20, RERS23, YZZ22]. Markovian [AL21]. MARS [KFG24]. massive [CP21b]. Matching [DD23]. Matérn [LSW21]. matrices [AFWZ20, AR21, BDP23, BDWW22, CHP20, CLC^{+}21a, CCF21, DZ20c, DY21, DZ24, Fan22, HW20, HY22, LW20, LWL21, LPYZ24, LR20b, LHBS23, LGRG23, MPW20, MWY23, MV21, Mou22, RW23, XYY20, ZZPZ22]. matrix [BM22, BDW21, BSS20, CW20, FWZ21, HKTY24, LHY20, MW24, PC20, QLY23, TV24]. matrix-valued [HKTY24]. Max [FJLX22, LLM20, PR22]. max-stable [PR22]. Max-sum [FJLX22]. maxima [BDH24, BZ23b, DZ20a, DHS20]. maximal [HLM23, RW23]. Maximum [FLS^{+}24, BGvZ12, BGvZ21, CS20, EJ20, Goo22, SG20]. MCMC [ALPW22, PHL20]. Mean [Hop20, BDH21, CW22a, CKP24, Ché24, DM22, DL22, GSS22, LBvdL21, LM21b, RGvdL22, ZZ20]. means [KKZ21, LL20, MH22, XY20]. measure [AC21, DR24, HdBCAM21, SP20, WN22]. measurement [BG21, JPV22, JZLM23]. measurements [ATW24, Nic24]. measures [ALMM22, BB21, Vel22]. Measuring [CLP21b]. median [LL20]. median-of-means [LL20]. Memorization [MZ22]. memory [Hid21]. Mendelian [MWW23, YSK21, ZWH^{+}20]. merging [VWW22]. Message [Fan22, MV21]. method [BZ23b, CHK^{+}24, Che21, CGZ22b, GLP20, MW22, RCSN22, WY20, YKN24]. methods [Dob20, HL24, MPK20, PTW21, PC20, SvZ20]. Metric [DCM24, ABL23, CM22b, HKSW21, KB22, LM21a]. metric-space [LM21a]. metrics [ZZYS20]. Metropolis [STG21]. microclustering [DCT21]. microscopy [KMW21]. mimic [RV22]. Minimax [BFR23, BGL^{+}20, CW21, CP22a, CHL^{+}24, FZP20, HR21, KBRW24, KBW22, LN22, LR20b, LGS21, MGS20, NBW21, NWB22, TY23, BNR23, CS22, Efr21, FNZ22, HW21b, KMW21, Mou22, MJ22]. Minimax-optimal [BGL^{+}20]. minimisation [DDP21]. minimization [MN20]. minimizer [CCZ24]. minimizers [CLvdG22, Han21]. minimizing [SJ22]. minimum [CCZ24, CDRV21, CLvdG22, Cox20, LS22b, RGvdL22]. minimum- [LS22b]. Missing [BS23c, CFMY21, Che22, Efr21, LPYZ24, SBC24, YCF24]. misspecification [DHK^{+}20]. mixed [EJ20, EB22, FSW21, LZ22a, MP23, PWLL21, RERS23]. mixing [WN22]. mixture [ADXR20, AY23, AD22, CLP21a, DP21, JLLX21, KM21b, LZZ21, MK21a, Nda22, SG20]. mixtures [DGMMR21, DWYZ23, LX23, MW22, WY20, YWR24]. MLE [CGZ22b, DKH21]. modal [GH21]. Model [BGL^{+}20, Eva20, GIKM22, Ing20, Tan20, BPS20, BDW21, BEPV22, BK20, CHK^{+}24, CLTZ20, CL24, CCNT21, DLS20, EKJ20, GM20, HSS22, KBB^{+}20, LHY20, Li22, LSC^{+}24, LZZ21, MBR21, Nda22, NCDK24, Pan23, Tia20, WSB^{+}23, XJL20, ZD22, ZMW22]. model- [NCDK24]. Model-assisted [Tan20]. model-based [LSC^{+}24, ZD22]. model-free [KBB^{+}20, Pan23]. model-X [WSB^{+}23]. modeling [Mur21]. models [ADXR20, AY23, AD22, BM22, BDSW22, BR23, BZ23a, BBNW20, BBSMW22, BW23, BFPM21, BFZ22, BMY20, CP22a, CLP21b, CJLZ24, CHLZ23, DHZ21, DCT21, DZ24, DN21, EJ20, EB22, FSW21, FMM23, FLS20, FGS21b, FY23, GvdVZ20, GKM21, GSUC21, GMPE21, GIKM22, GMW21, GX20, HS21b, JZLM23, JO20, KZ21, Kol22, KM21b, LZ22a, LGRG23, MK21a, MN20, PV20b, RS24, SS20, SS23, Sta24, Vel22, WT23, WN22, WZL22, XM23, XWZY24, ZA23, ZC23]. moderate [GSS22]. modern [HLS22]. moment [BDM22]. moments [CHK^{+}24, KKZ21, LS22a, MW22, WY20, ZX22]. Mondrian [MGS20]. Monitoring [HKW21]. monotone [BMNT23, CG21a, DHZ21, WC20]. monotonicity [FGS21a]. Monte [AL21, DMS20, FT23, HBDD20, KJJS20, KJJS22, MMS23]. Motif [BDM22].

MSE [DGR23]. MSE-optimal [DGR23]. multi [CCL24, Cha20, DZ20b, DW23a, GH21, NR20, PC20]. multi-armed [CCL24, Cha20]. multi-dimensional [DZ20b, NR20]. multi-layer [PC20]. multi-modal [GH21]. multi-task [DW23a]. multiclass [HTYZ20]. Multidimensional [KMW20, GR22]. multifactor [DLY22]. multilayer [CLM22, JLLX21]. multimodal [PHL20]. multinomial [BMY20, LLM20]. Multiple [ZVB21, ACR24, ATW24, CCZ22, CR20, DHZ21, FLY23, HZ20, HLM23, WGS^{+}22, ZLSK22]. multiplier [CZ20]. multiply [Cha22a, QTD24]. multiply-imputed [Cha22a]. multireference [DFZ24]. Multiscale [CvdP21, CP21a, KMW20]. multitask [HK22]. Multivariate [FGS21a, GS22, AWS22, BGvZ12, BGvZ21, BSS20, BNX22, CG21b, Ché24, DRR20, DP22, EJ20, FGKS21, GLL22, KBW20, KM21b, LM21b, PR22, PSD24, REZ23, WG23, WZL22, ZVB21].

Navier [NT24]. Near [LLM20, Lop22, EB22]. Near-parametric [LLM20]. nearest [CBS20]. nearly [DL22]. Necessary [KB22, Lah21]. need [XZZ22]. neighbour [CBS20]. Nested [RERS23]. Network [Lei21, AF23, JMM20, JKL21, KL21, LW22, SS23, ZX22]. networks [CLM22, FGZ24, GMMM21, HYF23, JPV20, JLLX21, KLR^{+}20, Kut20, MZ22, PC20, SH20a, SHV24, WYR21]. neural [GMMM21, JMM20, KL21, Kut20, MZ22, SH20a, SHV24]. no [HK22, XZZ22]. no-free-lunch [HK22]. Noise [CFMY21, AV23, DGR23, FGZ24, FLS^{+}24, GLL22, dÁLM21]. Noisy [Han23, AMBL23, DM23, SM20, XYZ21]. non [CP21a, JPV21, ZGV22]. non- [ZGV22]. non-Euclidean [CP21a, JPV21]. Nonasymptotic [JSLH23, RW20a, HRMS21]. Nonclassical [Zhi20]. nonconvex [CPT23, CFMY21, JZLM23, SWJ21]. Nonexchangeable [DCT21]. nonignorable [MK21b]. noninteractive [BRS23]. Nonlinear [AF23, SO23, CJQY24, DM23, JZLM23, LW20, MNP21, PSD24, SW07, SWZ22, ZSW22]. nonlinear-type [SWZ22]. nonmonotone [Che22]. Nonparametric [CPH23, CGC20, DRR20, Efr22, GR22, JPV22, Kut20, NR20, RSH20, SH20a, SBC24, AC24, ADXR20, AY23, AD22, BS23c, BCCG21, CW21, CW22b, CLP21b, CG21a, Cha20, DWW^{+}24, GS22, GLP20, HRMS21, JSLH23, KBW20, KH23, KM21b, LN22, LCI21, MPW23, SG20, SGWH20, SVvZ23, WC20, XS21, ZS22, vdVW21, SHV24]. Nonregular [FNZ22]. nonresponse [MK21b]. nonreversible [AWS22]. nonsmooth [DDJ24, SWJ21]. nonsparse [BFZ22, SF22]. nonspiked [CHP20]. nonstandard [DR21a, LY20]. nonstationary [BR23, DZ20c, DZ23, TCPS23, vDD24]. nonunique [LBvdL21]. norm [AWS22, BDH21, CLvdG22]. norma [LS22b]. normal [BSS20, Che21, Ché24, FY23, KZ21]. normality [DDJ24, QLY23]. normalization [WS20]. normalized [GSS22]. norms [TV24]. note [BGvZ21, FZ21, LBvdL21, ZSW22]. novelty [MLMR24]. null [RV22]. number [KM21b, MK21a, TYZ20, WL22, ZWL20]. numbers [RKP23]. numerical [HL20]. objects [DCM24]. oblique [CCK24]. observations [CSSV22, Ing20, LPYZ24, SLY22]. observed [HW23, KL20, WN22]. Off [HW23, BKRS23, CS22, DSH23, SS22]. Off-policy [HW23]. offline [LSC^{+}24, WQW23]. offs [DH24]. One [LC24, DM22, DMR22]. one-bit [DMR22]. One-step [LC24]. Online [GH21, HKTY24, FRB24, KR20, ZS22]. Only [GHS21]. operator [CP22b, DM23]. operators [BD20]. Optimal [ABL23, ATW24, BM22, BGvZ12, BGvZ21, BKS21, BS23c, BW23, CWZ21, CW22b, CFNP21, CCZ22, CNYY21, Cha22b, CGZ22a, DLS20, DWYZ23, HJWW20, JKL21, LX23, MBR21, PCV23, RCS23, SGWH20, SVvZ23, TSY20, TV24, VFLRB23, WYR21, WY20, XJL20, ADXR20, BB21, BNR23, BMY20, BGL^{+}20, CG21b, DGR23, FZP20, GS22, Han21, HWZ22, HR21, KB22, KL20, LSWMC24, Leu22, LO23, LBvdL21, MBNWW24, MPW20, MK21b, MGS20, Nda22, NBW21, SSL21, SM20, WT23, XYZ21, ZLN22]. Optimality [CCZ24, CGZ22b, LZZ21, OL24, CQZ23, DDJ24, DR21b, KBW22, LLHL21, LM21b, LZ22b, MJ22, ZCW22]. Optimally [MPW23]. Optimization [AO22a, AMBL23, CZ21, CPT23, CFMY21, DR21b, DN21, HMR23, JZLM23]. oracle [BG20, BS23b, RV22]. orbit [FLS^{+}24]. Order [SM23, BZ21, BK20, FZ16, FZ21, HL24, HCCI22, OvdG21]. Order-of-addition [SM23]. ordering [AL21, SM23]. Orthogonal [FS23, CT22, HLS22, PWLL21, SM23]. Osband [FZ21, FZ16]. other [BZ21, DHZ21, RKP23]. outcome [HLM23, LBvdL21]. outcomes [RGvdL22]. outlier [CDRV21]. outliers [BCL^{+}23, CFMY21]. overlapping [BBNW20, ZVB21]. overparametrization [HJ24].

p [VWW22]. p-values [VWW22]. Pairwise [BCCP23, CGZ22a, MW22, PMM^{+}20]. panel [BCPW23, FJLX22, ZZZ24]. panel-based [ZZZ24]. parabolic [Cho20]. parallel [Pol21, WM21]. Parameter [PSD24, ATW24, CWZ21, LSW21, MM21]. parameters [CLTZ20, GM20, Li22, LC24]. Parametric [ZGV22, LLM20]. Partial [CGZ22b, GX20, KS20, PMM^{+}20, ZS20]. partially [HW23, KL20, ZMW22]. particle [CSSV22, FLS^{+}24, KJJS20, KJJS22, LSV20]. partition [DCT21]. partitioning [CFF20]. partitioning-based [CFF20]. passing [MV21, Fan22]. paths [CGC20]. pathwise [BR24]. Pattern [Che22]. PCA [AFW22, CFMY21, Dob20, DH24, RW20a, SS22, YCF24, ZCW22]. PDEs [Cho20]. Penalized [TYZ20, Hua20, HS21b, OQC22]. penalties [CM22a]. performance [DN21]. Permutation [Dob20, PCV23, BKS21, KNBW22, KBW22]. permutations [MPW20, MW22, PS22]. persistent [CGK22, RKP23]. perspective [AD22, BKRS23, KMW21]. perspectives [HMR23]. perturbed [CCF21]. Peskun [AL21]. Phase [CKP24, BV20, CS20, GZ22a, Hua20, MZ22]. phenomenon [BS23b, BDM22, DLY22]. piecewise [BZ23b, CG21b, GHZ20]. planted [LZ22b]. plot [ZD22]. Plugin [MBNWW24]. Poincaré [ALPW22]. point [BCCP23, CNYY21, DM20, FLS20, HKTY24, HKW21, LGS21, VFLRB23, WYR21, XWZY24]. points [EB22, WZVS22, ZWL20]. Poisson [JZLM23, RSH20]. policy [HW23, LQW^{+}22]. polygon [JKL21]. polynomial [Hop20, JSLH23]. polynomials [CG21b, SW22]. population [LHY20]. populations [HY22]. position [SS23]. positive [FLS20]. positives [BNR20]. positivity [LUZ21, REZ23]. possible [DK22, GLL22, RV22]. possibly [LBvdL21, ZLSK22]. Post [BNR20, ZJ23, BPS20, KLYL20, KBB^{+}20, PTW21]. post-model-selection [BPS20]. Post-selection [ZJ23, KBB^{+}20, PTW21]. Posterior [RvdP20, SHSS^{+}21, KK20, Mur21, NT24, SM20, ZG20]. posteriors [AR20]. power [EH23, WSB^{+}23]. Powerful [SJ22]. practically [BDH24]. practice [LL20]. pre [CL24]. pre-averaging [CL24]. Precise [HJ24, JS22, LS22b]. precision [BDP23, DZ20c, LR20b]. Prediction [OvdG21, Tia20, BCRT23, DLS20, KB22, ZZZ24]. predictions [BK20]. Predictive [BCRT21, HHPS20, MJ22, SL23]. predictors [HLM23, JPV21]. prefactors [JSLH23]. presence [DK21, SAH21]. prices [OW21]. pricing [BCPW23]. Principal [FSW21, BDWW22, DY21, KLN20, PRV20]. principle [FWZ21, FRB24, FZ16, FZ21]. principles [CCZ24]. prior [RSH20]. priors [BSS20, MNP21, NT24, RS20b]. privacy [CWZ21, LBY23, RS20a]. private [AMBL23, AV23, BRS23, CHK^{+}24, DR24, LTMC24]. probability [BB21, KR20]. problem [Cha20, NIS^{+}21]. problems [AO22a, DH24, Han23, MNP21, PCV23]. procedure [MK21a]. procedures [BGvZ12, BGvZ21, DR21a, GHS21, WGS^{+}22]. process [BATU23, DDG21, HPTZ22, JT21, Li22, MNP21, Mur21, NT24]. processes [BCCP23, CL22, DRR20, FZP20, Hid21, HW23, HCCI22, LQW^{+}22, MPK20, MP23, YZZ22, ZW21, ZLN22, vDD24]. profession [FS21]. profile [ZWH^{+}20]. profiles [DCM24]. Projected [WQW23]. projection [KBW20]. projections [KPST23]. Projective [SS23]. promotion [BEPV22]. properties [BCST20, CFF20, CVFL22, HS21b, RERS23, RS20b]. proportion [KR20]. proportions [GHS21]. Propriety [Mur21]. protocol [CW22a]. proxy [KRSW21]. pseudo [ALPW22, GKM21]. pseudo-likelihood [GKM21]. pseudo-marginal [ALPW22]. pure [LN22].

quadratic [BRS23, CZ21]. quantification [BG20, CR21, MNP21]. quantify [vDD24]. quantifying [CS22]. quantile [BCPW23, HdBCAM21, HPTZ22]. quantiles [GS22, KP23]. quantization [DMR22]. quantum [FY23, LN22].

Random [BS23c, LW22, AFWZ20, BD20, BM22, CPT23, CVFL22, DZ20a, DCT21, DD23, DCM24, GSS22, GHO22, GGCvL20, HJ24, KB22, LPYZ24, LSW21, SS20, SGWH20, STG21, TCPS23]. randomization [Dob22, MWW23, YSK21, ZWH^{+}20]. randomized [Leu22, LH23]. range [SWFD21, ZLN22]. Rank [CL24, LEV21, ZXZL24, AFWZ20, CP22b, CCF21, CHLZ23, DHS20, FWZ21, HYF23, LWL21, LX23, MV21, MW24, RW23, SHDH22, XYZ21, XZZ22].

Rank-based [LEV21, ZXZL24]. ranking [CGZ22a, CGZ22b]. ranks [GS22, HdBCAM21]. Rao [YWR24]. rare [DK22, PS20]. rare/weak [PS20]. Rate [Leu22, WT23, CW21, CW22b, CCZ22, DHK^{+}20, FL22, Han21, KL21, MLMR24, RV22, SXF24, SM20, TY23, ZL24]. Rate-optimal [Leu22, WT23, SM20]. Rates [DFZ24, HS21b, BB21, BKS21, BFR23, CWZ21, CCK24, GS22, HS21a, HJWW20, Hop20, KBRW24, LGS21, LLM20, MGS20, RS20a, TSY20, XJL20, ZG20]. ratio [BM22, BMNT23, HSS22]. ratios [MW24]. reach [ABL23]. recollections [Che21]. Reconciling [NCDK24, RY21, ZD22]. reconstructability [SJ22]. reconstruction [KL20, RW20a]. recovery [BGL^{+}20, CGZ22b, DD23, FWZ21, GZ22b, Nda22, RSH20]. rectangular [TV24]. reduction [CJQY24, DQW20, KLYL20, KZ21, LL22, LS22a]. reference [BNR20, Mur21, ZZZ24]. Refined [GSS22]. reflections [Che21]. regime [LLSS21]. regimes [LSWMC24, OQC22]. regions [EB22, FLS20, LY20]. regression [BATU23, BGvZ12, BG20, BGvZ21, BCPW23, BV20, BM23, BNX22, BJ23, CS20, CCK24, CM22a, CG21a, CM22b, CS22, DSUC23, DZ20b, DHZ21, DS21, FGS21a, GZ22a, GW22, GLP20, GSUC21, HZ20, HTYZ20, HS23, HJ24, HHPS20, HPTZ22, Ing20, JP20, JPV21, JPV22, JT21, JSLH23, KR20, KL21, KBB^{+}20, Kut20, LLSS21, LPMP24, LR20a, LLHL21, LM21a, MPW23, OQC22, PS22, PLD21, RvdP20, SH20a, SHV24, SGWH20, SF22, WG23, XWZY24, ZP22, ZS22, ZGV22, dÁLM21]. regression-based [OQC22]. regressions [GHO22, dRdCMP21]. regret [BNR23]. regrets [CP22a]. regular [XM23]. regularization [BNR23, DGMMR21, ZGMS24, dÁLM21]. regularized [BZ23a, CLvdG22, HSL24, HS23, JLLX21, LM21a, OvdG21, Tan20, WSB^{+}23]. reinforcement [LSC^{+}24, WQW23]. related [DH24]. Relaxing [BNR23, FGRS22, HJ20]. relevant [BDH21]. ReLU [Kut20, FGZ24, SH20a, SHV24]. replicating [WGS^{+}22]. representation [Lei21]. Rerandomization [WL22]. Rerandomized [MWW23]. resampling [CSSV22, ZLN22]. research [BFdGG21]. residual [HYF23]. residuals [Vel22]. resilience [ZJS22]. resolution [KMW21]. response [BNX22, CJLZ24, FGS21b, PLD21]. responses [BJ23, FNZ22, JP20, JPV21]. restricted [GX20]. restrictions [KH23]. results [DDG21, SS20]. reversible [AL21, BS23a, GR22]. revisited [NIS^{+}21, ZP22]. reward [LQW^{+}22]. Rho [BB20]. Rho-Bayes [BB20]. Ridge [ZP22]. Ridgeless [LR20a, HMRT22]. Riemannian [SLY22]. right [DR24]. risk [CFNP21, CLvdG22, CS22, DR21a, GLCS20, Han21, Han23, HHPS20, Hua20, MN20, Mou22, QTD24]. risk-fairness [CS22]. RKHS [MPW23, ZZYS20]. RKHS-based [MPW23, ZZYS20]. Robust [BB20, BCS20, BFL21, CZ20, DL22, FLY23, GLN20, KBW20, KKZ21, LL20, LM21b, MZ20, OW21, SM20, CFMY21, CCNT21, DM22, DWW^{+}24, DW23a, DN21, FWZ21, HJ24, Hua20, NCDK24, QTD24, WT23, YZZ22, ZWH^{+}20, ZJS22]. robustness [CLvdG22, LBY23]. root [BE22, BNR23, Lei21]. root-entropic [BNR23]. rotationally [Fan22]. rough [CHL^{+}24]. rule [LBvdL21, Pol21]. Run [Pol21, BDP23].

saddlepoint [Goo22]. Safe [DP21]. Sample [YZZ22, Bha20, CHP20, CFF20, GW22, GGCvL20, HSL24, HY22, KRSW21, LHY20, LSC^{+}24, LHBS23, MMS23, Mou22, QLY23, XYY20, XY20, YSK21, ZZPZ22, ZGMS24, ZWH^{+}20, Zhi22, ZWL20]. sample-splitting [ZWL20]. sampling [BJ22, BDM22, DP21, HTYZ20, HW21a, LSV20]. Scalable [HPTZ22, BV20]. scalar [BR24]. scale [CCZ22, GHO22, HTYZ20, KW22, MS20]. scales [HT23]. scanning [KMW20]. scenario [AL21]. Schatten [TV24]. schemes [CSSV22, PSD24]. score [ZWH^{+}20]. ScreeNOT [DGR23]. search [Tia20]. Second [BZ21, SMZ22]. Second-order [BZ21]. sectional [FJLX22]. Segmentation [FLS20]. selection [BPS20, BMNT23, CDRV21, Eva20, GKM21, GLP20, GIKM22, HMR23, HCCI22, Ing20, JT21, KBB^{+}20, Lah21, LTMC24, PTW21, RCS23, WWM20, ZL24, ZWL20, ZJ23]. self [GSS22, WS20]. self-normalization [WS20]. self-normalized [GSS22]. selfnormalization [WZVS22]. semi [CBS20]. semi-supervised [CBS20]. Semiparametric [MK21b, RvdV20, WZL22, BEV20, CJLZ24, ZGV22, dRdCMP21]. sensitive [PS20]. sensitivity [DK22]. separability [BD20]. separable [DY21]. sequence [BS23a, HSS22, HKW21]. sequences [ACR24, HRMS21, WN22]. sequential [FZP20, FT23, HBDD20, HW23, KJJS20, KJJS22, MMS23]. series [AvD20, BR23, BZ23b, CFF20, CNYY21, Cha22b, CL24, DZ20c, DZ23, Ger24, HKTY24, LLS24, SW07, Vel22, WS20, YKN24, ZVB21, ZSW22]. Set [Han21, BE22]. sets [RCSN22, RS20b]. settings [KR20]. Settling [LSC^{+}24]. shape [BS23a, KH23, RCSN22, XS21]. shape-constrained [BS23a]. Sharp [ACR24, BR24, CPT23, CQZ23, Efr21, KBG20, Nda22]. shift [KM21a, MPW23, QTD24]. short [SWFD21]. short-range [SWFD21]. shrinkage [DZ24, FWZ21, FGRS22, LW20]. sieve [RS20b, ZS22]. sign [LTMC24]. signal [BM22, CC23]. signals [CQZ23, GHZ20, GLL22, SO23, WGS^{+}22]. signed [JKL21]. signed-polygon [JKL21]. significance [FNL20]. significant [BDH24]. signs [HdBCAM21]. similarity [BR24]. simple [AC21]. simplices [NIS^{+}21]. Simultaneous [KR20, ZF24]. Single [BM23, FLS^{+}24]. single-particle [FLS^{+}24]. Singular [BDW21, DGR23, EB22, MWY23]. Singularity [DHK^{+}20, KM21a]. SIR [TSY20]. size [BGvZ12, BGvZ21, DDB20]. sizes [ZVB21]. sketching [GW22]. slab [CR20]. sliced [LLHL21]. slicing [DQW20]. sliding [BZ23b]. SLOPE [BKRS23]. smooth [HR21, KZ21, Kol22, LSWMC24, MBNWW24, NWB22, TCPS23]. smoothing [DC23, LPMP24]. Sobolev [ZS22]. solution [Cha20]. Some [BCST20, Che21, BM22]. Sort [MK21a]. space [CT22, CM22b, CHHW21, DKA20, EFdH^{+}22, GIKM22, LM21a]. space-filling [CT22]. space-time [EFdH^{+}22]. spaces [DZ20b, HKSW21, KB22]. Sparse [BV20, DGMMR21, GZ22a, TSY20, ALS23, ACR24, BBSMW22, BCCG21, CW20, CZ21, CQZ23, CCNT21, FMM23, GLN20, GMW21, LGS21, MM21, MJ22, SLY22, SS23, WYR21, ZS20]. sparsity [BMNT23]. Spatial [EFdH^{+}22, KP23, LEV21, Leu22]. SPDEs [ATW24]. specific [RGvdL22]. Spectral [CL22, HSL24, LPYZ24, PC20, AFW22, AR21, BM22, CGZ22b, LWL21, LZZ21, SW07, XYY20, ZW21, ZZPZ22, ZSW22]. spectrum [GKV^{+}22]. spherical [CM21, PV20a]. spherically [MS20]. spike [CR20]. Spiked [DY21, BM22, BDWW22, CHP20, DLS20, EKJ20, JO20, LHY20, ZZPZ22]. spikes [LHBS23]. spline [HS21b]. split [ZD22]. split-plot [ZD22]. splitting [PSD24, ZWL20]. square [ZA23]. squared [GZ22a]. squares [HMRT22, KP22, Mou22, OvdG21, ZS20]. stability [ZJ23]. stabilizing [RKP23]. stable [BR24, MP23, PR22, SL23]. stage [BGvZ12, BGvZ21, LFR21]. standard [ERV20]. StarTrek [ZL24]. state [WQW23]. state-action [WQW23]. states [LN22]. stationarity [AvD20]. stationary [BCCP23, BZ23b, CGK22, MPK20, ZW21]. Statist [BGvZ21]. statistic [BKS21]. Statistical [BDWW22, CW20, CLTZ20, CHL^{+}24, Dre23, DH24, GMW21, LZ22b, MNP21, ZWH^{+}20, AO22a, CLC^{+}21a, CP21b, FS23, FY23, HWZ22, HMR23, JS22, KLYL20, KMW20, KMW21, LPYZ24, NIS^{+}21, NR20, SWJ21, SWZ22, SS22, LBvdL21]. Statistical-computational [DH24]. Statistically [XYZ21]. Statistics [PS22, BM22, BDH24, DCM24, HXWP21, HL24, JKL21, LWL21, LLM20, POLdHG24, QLY23, RKP23, WSB^{+}23, ZZPZ22, ZJS22]. status [Efr21]. Stein [BZ21, Che21, EG21, MH22, Str21, vdVW21]. step [BEV20, DDB20, LC24]. Stigum [ALS23]. Stochastic [CP22a, BB21, CLTZ20, Cho20, CGC20, DDJ24, DDB20, DR21b, PSD24, XJL20, ZS22]. Stokes [NT24]. stopping [Sta24]. streams [Pol21]. strength [PWLL21]. Strong [GKM21]. structural [BFPM21, LLS24, vDD24]. structure [Dre23, GZ22b, GLP20, HKTY24, KW22, WN22, YLFS22]. structured [BBNW20, BK20, GvdVZ20, Han21, KR20, LGRG23]. structures [BFZ22, DM23, LZ22b]. studies [NLS21]. study [CT22, SM23, WC20]. sub [DL22, Hop20]. sub-Gaussian [DL22, Hop20]. subgraph [BDM22]. subgroup [RCS23]. submanifolds [TY23]. SubOptimality [CGZ22b]. subordinated [ZLN22]. subsampling [HYF23]. subsets [EJ20]. Subspace [CLC^{+}21a, BDW21]. successive [KPST23]. sufficient [BJ22, CJQY24, DQW20, KB22, Lah21, LL22]. sum [FJLX22]. summary [YSK21, ZWH^{+}20]. summary-data [YSK21, ZWH^{+}20]. sums [DZ20a, GSS22]. Sup [AWS22, BDH21]. Sup-norm [AWS22, BDH21]. super [BS23b]. super-oracle [BS23b]. superhedging [OW21]. SuperMix [DGMMR21]. superresolution [KMW21]. Supervised [WWZ^{+}24, CBS20]. support [HS21a, RSH20]. sure [Cox20, BZ21, FGRS22]. surfaces [BD20]. Surprises [HMRT22]. surrogate [LSWMC24, SWJ21]. Survival [CLP21a, CvdP21, HLM23, PV20b, WZL22]. symmetric [MS20]. symmetry [GIKM22, XS21]. synchrosqueezing [SWZ22]. systems [KJJS20, KJJS22].

Table [Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano24e, Ano24f, Ano24g, Ano24h]. tables [DK22]. tackling [MPW23]. tail [ES21, Mou22]. tailed [AC24, BCCG21, DSUC23, FWZ21, GSUC21, GMPE21, HY22, KP22]. tails [FGZ24]. Tanner [BKRS23]. TAP [CFM23]. targeted [RGvdL22]. targeting [PHL20]. task [DW23a]. teacher [vdGK21]. tempered [AR20]. temporal [XWZY24]. Tensor [LZ22b, CL24, DH24, HWZ22, JLLX21, XYZ21]. tensors [XZZ22]. Test [FNL20, HW20, ALMM22, BD20, DWW^{+}24, SWFD21, XY20, ZA23]. Testability [BFZ22]. Testing [AvD20, BJ22, BDH24, BCL^{+}23, BP24, CP22b, GM22, JO20, KH23, PRV20, YLFS22, ACR24, BKS21, BS23c, BR24, BDH21, CCZ22, CR20, CMW23, CQZ23, CPH23, DHS20, FLY23, GW22, GS22, GGCvL20, GHS21, HXWP21, JZLM23, JKL21, KRSW21, KMW21, NBW21, NCDK24, POLdHG24, SP20, SVvZ23, WS20, ZXZL24, vdVW21]. tests [AV23, BM22, BKS21, Bha20, Cha22a, Dob22, DW23b, FZP20, FJLX22, HSL24, HSS22, KBW20, KNBW22, KBW22, PLD21, PS20, SHDH22]. their [AR20, AMR20, DR21a, ZC23]. theorem [CCKK22, EG21, HK22, LWL21, Lop22, dRdCMP21]. theorems [AO22b, GSS22, HW21a, LPYZ24]. Theoretical [ZZ20, BCST20, YZZ22]. Theory [LL20, AFW22, CHL^{+}24, HZ20, KS20, KMW20, LS22b, SW07, ZSW22]. thesis [vdGK21]. threshold [ALS23, DD23]. thresholding [DGR23, GLN20, SF22, ZP22]. thresholds [Bha20, FNZ22]. thumb [Pol21]. tie [LO23]. tie-breaker [LO23]. Tierney [AL21]. Time [HRMS21, AvD20, BR23, BEPV22, BZ23b, CNYY21, Cha22b, CM22b, CL24, CHHW21, DL22, DZ20c, DZ23, EFdH^{+}22, Ger24, HKTY24, Hop20, LCI21, LLS24, RW23, RGvdL22, SO23, SW07, SWZ22, Vel22, WS20, YKN24, ZSW22, ZVB21]. time-dependent [LCI21]. time-frequency [SWZ22]. Time-uniform [HRMS21]. time-varying [RW23]. top [CGZ22b]. top- [CGZ22b]. topic [BBSMW22]. topics [KPST23]. Total [LUZ21, LM21a, REZ23, FGS21a, GLCS20, KYZ21, OvdG21, dÁLM21]. trace [LHY20]. trade [BKRS23, CS22, DSH23, DH24, SS22]. trade-off [BKRS23, CS22, DSH23, SS22]. trade-offs [DH24]. training [HJ24, JS22, MZ22]. trajectories [CM22b]. Transfer [CW21, CCL24, CKP24, RCS21]. transform [SWZ22]. transform-toward [SWZ22]. transformations [AO22b]. transformed [DLS20]. transition [CKP24, CS20, GZ22a, Hua20, MZ22]. transitions [BV20]. transport [GS22, HR21, MBNWW24]. transportation [BB21, HdBCAM21]. treatment [BJZ24, DWW^{+}24, JPV20, LSWMC24, LW22, LBvdL21, OQC22, SAH21, SM23, SSL21, Tan20, YNB^{+}22]. tree [BK20]. tree-structured [BK20]. trees [CCK24, CG21b, MGS20, RvdP20]. trek [BDSW22]. trend [CC23, EFdH^{+}22, GLCS20]. trimmed [LM21b]. tuning [FLY23, MM21]. Two [BEV20, BRS23, GW22, GGCvL20, WM21, BGvZ12, BGvZ21, Bha20, DK22, GMMM21, HSL24, HKTY24, JMM20, KRSW21, KKZ21, LLSS21, LFR21, Nda22, XY20, YSK21, ZWH^{+}20, ZXZL24]. two-layer [JMM20]. two-layers [GMMM21]. Two-level [WM21]. two-regime [LLSS21]. Two-sample [GW22, GGCvL20, Bha20, HSL24, KRSW21, XY20, YSK21, ZWH^{+}20]. two-stage [BGvZ12, BGvZ21, LFR21]. Two-step [BEV20]. two-way [HKTY24]. Tyler [GLN20]. type [CJN24, GSS22, SWZ22, Zhi22]. types [Pol21].

unbalanced [CLC^{+}21a]. Uncertainty [CR21, BG20, CCZ24, HTYZ20, MNP21]. understandings [YZZ22]. unified [WC20]. Uniform [AY23, CM22b, HW21a, MM21, DN21, HRMS21]. Uniformly [BPS20]. uniqueness [Cox20, DKH21]. univariate [FGS21b, GLCS20]. Universal [BJ23, HYF23, HKSW21]. Universality [HS23]. universally [SHDH22]. unknown [CW22a, Ché24, GLL22, MPW20, MS20, PS22, RV22, SAH21, TY23]. unlabeled [KLR^{+}20]. Unmixing [NIS^{+}21]. unobserved [LH23, YNB^{+}22]. upper [RW20a]. using [BCCP23, BNR20, EJ20, GS22, KS20, Kut20, LSWMC24, Lei21, MP23, RW23, SH20a, SHV24, Tan20, Vel22, WGS^{+}22, YWR24, ZWH^{+}20].

Valid [KBB^{+}20, BPS20, POLdHG24]. validated [CLC21b]. validation [CC23]. validity [Kle21, KK20, MPK20]. value [DGR23, Dre23, EH23, HHPS20, SSL21]. value-at-risk [HHPS20]. valued [CM22b, HKTY24, LM21a, LC24, vDD24]. values [BCL^{+}23, MWY23, VW21, VWW22]. Variable [BMNT23, JT21, BDSW22, BGL^{+}20, GLP20, HMR23, Lah21, WWM20, ZL24]. variables [BNX22, BFPM21, GSS22, LUZ21, ZLSK22]. variance [BDP23, CW22a, Cha22b, DSH23, LLM20, MWW23, SWFD21, SGWH20, YSK21]. variation [CG21b, FGS21a, GLCS20, LM21a, OvdG21, dÁLM21]. Variational [RW20b, AR20, BKRS23, KR24, OL24, ZG20, ZZ20]. VARs [BG22]. varying [BDH21, RW23]. vector [BDW21, CCNT21, DL22, GKM21, HS21a, NR20, SHDH22, WT23]. vectors [DZ20a, MS20, ZXZL24]. versus [BRS23, PMM^{+}20, SS22]. via [ALPW22, AMBL23, BDH24, BKS21, BDM22, BFR23, BNR23, CLP21a, CHK^{+}24, CZ20, DN21, DH24, FGS21a, GW22, HYF23, HL24, JMM20, JLLX21, KFG24, KBW20, MK21a, MV21, SJ22, WS20, WZVS22, WY20, XS21, ZWL20, ZJ23]. Volatility [JLL21, CHL^{+}24, CGK22, DZ24, SMZ22]. vote [LTMC24].

walk [STG21]. warping [CM22b]. Wasserstein [BBSMW22, CLP21b, NWB22, PLD21, RS24, YWR24, ZGMS24]. way [HKTY24]. ways [VWW22]. weak [ALPW22, BE22, DK22, LS22a, LLM20, PRV20, PS20]. weakly [CSSV22, dRdCMP21]. weighted [MWW23, XJL20, YSK21]. weights [WQW23]. Which [WWM20]. white [dÁLM21]. Whittle [BDP23, RY21].

Wigner [EKJ20]. Wilks [dRdCMP21]. Willem [BFdGG21, FS21, vdGK21]. winner [MWW23]. winsorized [GSS22]. without [Cox20, Efr21]. Wordlength [TX21]. working [HS21a]. Worst [PMM^{+}20]. Worst-case [PMM^{+}20]. Wynn [FGS21b].

X [WSB^{+}23].

Zwet [BFdGG21, FS21, vdGK21].

References

Aamari:2023:ORE

[ABL23] Eddie Aamari, Clément Berenfeld, and Clément Levrard. Optimal reach estimation and metric learning. Annals of Statistics, 51(3):1086–1108, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-3/
Optimal-reach-estimation-and-metric-learning/10.1214/
23-AOS2281.full.

Azadkia:2021:SMC

[AC21] Mona Azadkia and Sourav Chatterjee. A simple measure of conditional dependence. Annals of Statistics, 49(6):3070– 3102, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-6/
A-simple-measure-of-conditional-dependence/10.1214/21-
AOS2073.full.

Agapiou:2024:HTB

[AC24] Sergios Agapiou and Ismaël Castillo. Heavy-tailed Bayesian nonparametric adaptation. Annals of Statistics, 52(4):1433– 1459, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-4/
Heavy-tailed-Bayesian-nonparametric-adaptation/10.1214/
24-AOS2397.full.

Abraham:2024:SMT

[ACR24] Kweku Abraham, Ismaël Castillo, and Étienne Roquain. Sharp multiple testing boundary for sparse sequences. Annals of Statistics, 52(4):1564–1591, August 2024. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-4/Sharp-multiple-testing-
boundary-for-sparse-sequences/10.1214/24-AOS2404.full.

Argiento:2022:IFB

[AD22] Raffaele Argiento and Maria De Iorio. Is infinity that far? A Bayesian nonparametric perspective of finite mixture models. Annals of Statistics, 50(5):2641–2663, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-5/Is-infinity-that-far-
A-Bayesian-nonparametric-perspective-of-finite/10.1214/
22-AOS2201.full.

Aragam:2020:INM

[ADXR20] Bryon Aragam, Chen Dan, Eric P. Xing, and Pradeep Ravikumar. Identifiability of nonparametric mixture models and Bayes optimal clustering. Annals of Statistics, 48(4):2277– 2302, August 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-4/
Identifiability-of-nonparametric-mixture-models-and-
Bayes-optimal-clustering/10.1214/19-AOS1887.full.

Armillotta:2023:NNA

[AF23] Mirko Armillotta and Konstantinos Fokianos. Nonlinear network autoregression. Annals of Statistics, 51(6):2526– 2552, December 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-6/
Nonlinear-network-autoregression/10.1214/23-AOS2345.
full.

Abbe:2022:TPS

[AFW22] Emmanuel Abbe, Jianqing Fan, and Kaizheng Wang. An ℓ_{p} theory of PCA and spectral clustering. Annals of Statistics, 50(4): 2359–2385, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-4/
An-%e2%84%93p-theory-of-PCA-and-spectral-clustering/
10.1214/22-AOS2196.full.

Abbe:2020:EEA

[AFWZ20] Emmanuel Abbe, Jianqing Fan, Kaizheng Wang, and Yiqiao Zhong. Entrywise eigenvector analysis of random matrices with low expected rank. Annals of Statistics, 48(3): 1452–1474, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-3/
Entrywise-eigenvector-analysis-of-random-matrices-with-
low-expected-rank/10.1214/19-AOS1854.full.

Andrieu:2021:PTO

[AL21] Christophe Andrieu and Samuel Livingstone. Peskun–Tierney ordering for Markovian Monte Carlo: Beyond the reversible scenario. Annals of Statistics, 49(4):1958–1981, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-49/issue-4/PeskunTierney-
ordering-for-Markovian-Monte-Carlo--Beyond-the-reversible/
10.1214/20-AOS2008.full.

Albert:2022:ATI

[ALMM22] Mélisande Albert, Béatrice Laurent, Amandine Marrel, and Anouar Meynaoui. Adaptive test of independence based on HSIC measures. Annals of Statistics, 50(2):858–879, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-50/issue-2/Adaptive-test-
of-independence-based-on-HSIC-measures/10.1214/21-AOS2129.
full.

Andrieu:2022:CMC

[ALPW22] Christophe Andrieu, Anthony Lee, Sam Power, and Andi Q. Wang. Comparison of Markov chains via weak Poincaré inequalities with application to pseudo-marginal MCMC. Annals of Statistics, 50(6):3592–3618, December 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-6/Comparison-of-Markov-
chains-via-weak-Poincar%c3%a9-inequalities-with-application/
10.1214/22-AOS2241.full.

Abbe:2023:LSG

[ALS23] Emmanuel Abbe, Shuangping Li, and Allan Sly. Learning sparse graphons and the generalized Kesten–Stigum threshold. Annals of Statistics, 51(2):599–623, April 2023. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-2/Learning-sparse-graphons-
and-the-generalized-KestenStigum-threshold/10.1214/23-
AOS2262.full.

Avella-Medina:2023:DPI

[AMBL23] Marco Avella-Medina, Casey Bradshaw, and Po-Ling Loh. Differentially private inference via noisy optimization. Annals of Statistics, 51(5):2067–2092, October 2023. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-5/Differentially-private-
inference-via-noisy-optimization/10.1214/23-AOS2321.
full.

Anderes:2020:ICF

[AMR20] Ethan Anderes, Jesper Møller, and Jakob G. Rasmussen. Isotropic covariance functions on graphs and their edges. Annals of Statistics, 48(4):2478–2503, August 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-4/Isotropic-covariance-
functions-on-graphs-and-their-edges/10.1214/19-AOS1896.
full.

Anonymous:2023:EBa

[Ano23a] Anonymous. Editorial board. Annals of Statistics, 51(1): ??, February 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-1/
Editorial-Board/aos511p1.full.

Anonymous:2023:EBb

[Ano23b] Anonymous. Editorial board. Annals of Statistics, 51 (2):??, April 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-2/
Editorial-Board/aos512p1.full.

Anonymous:2023:EBc

[Ano23c] Anonymous. Editorial board. Annals of Statistics, 51 (3):??, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-3/
Editorial-Board/aos513p1.full.

Anonymous:2023:EBd

[Ano23d] Anonymous. Editorial board. Annals of Statistics, 51 (4):??, August 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-4/
Editorial-Board/aos514p1.full.

Anonymous:2023:EBe

[Ano23e] Anonymous. Editorial board. Annals of Statistics, 51(5): ??, October 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-5/
Editorial-Board/aos515p1.full.

Anonymous:2023:EBf

[Ano23f] Anonymous. Editorial board. Annals of Statistics, 51(6): ??, December 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-6/
Editorial-Board/aos516p1.full.

Anonymous:2023:TCa

[Ano23g] Anonymous. Table of contents. Annals of Statistics, 51 (1):??, February 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-1/
Table-of-Contents/aos511c1.full.

Anonymous:2023:TCb

[Ano23h] Anonymous. Table of contents. Annals of Statistics, 51 (2):??, April 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-2/
Table-of-Contents/aos512c1.full.

Anonymous:2023:TCc

[Ano23i] Anonymous. Table of contents. Annals of Statistics, 51 (3):??, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-3/
Table-of-Contents/aos513c1.full.

Anonymous:2023:TCd

[Ano23j] Anonymous. Table of contents. Annals of Statistics, 51 (4):??, August 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-4/
Table-of-Contents/aos514c1.full.

Anonymous:2023:TCe

[Ano23k] Anonymous. Table of contents. Annals of Statistics, 51 (5):??, October 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-5/
Table-of-Contents/aos515c1.full.

Anonymous:2023:TCf

[Ano23l] Anonymous. Table of contents. Annals of Statistics, 51 (6):??, December 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-6/
Table-of-Contents/aos516c1.full.

Anonymous:2024:EBa

[Ano24a] Anonymous. Editorial board. Annals of Statistics, 52(1): ??, February 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-1/
Editorial-Board/aos521p1.full.

Anonymous:2024:EBb

[Ano24b] Anonymous. Editorial board. Annals of Statistics, 52 (2):??, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-2/
Editorial-Board/aos522p1.full.

Anonymous:2024:EBc

[Ano24c] Anonymous. Editorial board. Annals of Statistics, 52 (3):??, June 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-3/
Editorial-Board/aos523p1.full.

Anonymous:2024:EBd

[Ano24d] Anonymous. Editorial board. Annals of Statistics, 52 (4):??, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-4/
Editorial-Board/aos524p1.full.

Anonymous:2024:TCa

[Ano24e] Anonymous. Table of contents. Annals of Statistics, 52 (1):??, February 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-1/
Table-of-Contents/aos521c1.full.

Anonymous:2024:TCb

[Ano24f] Anonymous. Table of contents. Annals of Statistics, 52 (2):??, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-2/
Table-of-Contents/aos522c1.full.

Anonymous:2024:TCc

[Ano24g] Anonymous. Table of contents. Annals of Statistics, 52 (3):??, June 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-3/
Table-of-Contents/aos523c1.full.

Anonymous:2024:TCd

[Ano24h] Anonymous. Table of contents. Annals of Statistics, 52 (4):??, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-4/
Table-of-Contents/aos524c1.full.

Aswani:2022:OHF

[AO22a] Anil Aswani and Matt Olfat. Optimization hierarchy for fair statistical decision problems. Annals of Statistics, 50(6):3144– 3173, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-6/
Optimization-hierarchy-for-fair-statistical-decision-
problems/10.1214/22-AOS2217.full.

Austern:2022:LTD

[AO22b] Morgane Austern and Peter Orbanz. Limit theorems for distributions invariant under groups of transformations. Annals of Statistics, 50(4):1960–1991, August 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-4/Limit-theorems-for-distributions-
invariant-under-groups-of-transformations/10.1214/21-
AOS2165.full.

Alquier:2020:CTP

[AR20] Pierre Alquier and James Ridgway. Concentration of tempered posteriors and of their variational approximations. Annals of Statistics, 48(3):1475–1497, June 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-3/Concentration-of-tempered-
posteriors-and-of-their-variational-approximations/10.
1214/19-AOS1855.full.

Amini:2021:CKM

[AR21] Arash A. Amini and Zahra S. Razaee. Concentration of kernel matrices with application to kernel spectral clustering. Annals of Statistics, 49(1):531–556, February 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-1/Concentration-of-kernel-
matrices-with-application-to-kernel-spectral-clustering/
10.1214/20-AOS1967.full.

Altmeyer:2024:OPE

[ATW24] Randolf Altmeyer, Anton Tiepner, and Martin Wahl. Optimal parameter estimation for linear SPDEs from multiple measurements. Annals of Statistics, 52(4):1307–1333, August 2024.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-4/Optimal-parameter-estimation-
for-linear-SPDEs-from-multiple-measurements/10.1214/
24-AOS2364.full.

Awan:2023:CND

[AV23] Jordan Awan and Salil Vadhan. Canonical noise distributions and private hypothesis tests. Annals of Statistics, 51 (2):547–572, April 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-2/
Canonical-noise-distributions-and-private-hypothesis-
tests/10.1214/23-AOS2259.full.

Aue:2020:TSF

[AvD20] Alexander Aue and Anne van Delft. Testing for stationarity of functional time series in the frequency domain. Annals of Statistics, 48(5):2505–2547, October 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-5/Testing-for-stationarity-
of-functional-time-series-in-the-frequency/10.1214/19-
AOS1895.full.

Aeckerle-Willems:2022:SNA

[AWS22] Cathrine Aeckerle-Willems and Claudia Strauch. Sup-norm adaptive drift estimation for multivariate nonreversible diffusions. Annals of Statistics, 50(6):3484–3509, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-6/Sup-norm-adaptive-drift-
estimation-for-multivariate-nonreversible-diffusions/
10.1214/22-AOS2237.full.

Aragam:2023:UCN

[AY23] Bryon Aragam and Ruiyi Yang. Uniform consistency in nonparametric mixture models. Annals of Statistics, 51(1):362– 390, February 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-1/
Uniform-consistency-in-nonparametric-mixture-models/
10.1214/22-AOS2255.full.

Barthelme:2023:GPR

[BATU23] Simon Barthelmé, Pierre-Olivier Amblard, Nicolas Tremblay, and Konstantin Usevich. Gaussian process regression in the flat limit. Annals of Statistics, 51(6):2471– 2505, December 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-6/
Gaussian-process-regression-in-the-flat-limit/10.1214/
23-AOS2336.full.

Baraud:2020:RBL

[BB20] Yannick Baraud and Lucien Birgé. Robust Bayes-like estimation: Rho-Bayes estimation. Annals of Statistics, 48(6):3699– 3720, December 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-6/
Robust-Bayes-like-estimation-Rho-Bayes-estimation/10.
1214/20-AOS1948.full.

Bercu:2021:ADC

[BB21] Bernard Bercu and Jérémie Bigot. Asymptotic distribution and convergence rates of stochastic algorithms for entropic optimal transportation between probability measures. Annals of Statistics, 49(2):968–987, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
49/issue-2/Asymptotic-distribution-and-convergence-rates-
of-stochastic-algorithms-for-entropic/10.1214/20-AOS1987.
full.

Bing:2020:AES

[BBNW20] Xin Bing, Florentina Bunea, Yang Ning, and Marten Wegkamp. Adaptive estimation in structured factor models with applications to overlapping clustering. Annals of Statistics, 48(4): 2055–2081, August 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-4/
Adaptive-estimation-in-structured-factor-models-with-
applications-to-overlapping/10.1214/19-AOS1877.full.

Bing:2022:LES

[BBSMW22] Xin Bing, Florentina Bunea, Seth Strimas-Mackey, and Marten Wegkamp. Likelihood estimation of sparse topic distributions in topic models and its applications to Wasserstein document distance calculations. Annals of Statistics, 50(6):3307– 3333, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-6/
Likelihood-estimation-of-sparse-topic-distributions-
in-topic-models-and/10.1214/22-AOS2229.full.

Borgs:2021:CNE

[BCCG21] Christian Borgs, Jennifer T. Chayes, Henry Cohn, and Shirshendu Ganguly. Consistent nonparametric estimation for heavy-tailed sparse graphs. Annals of Statistics, 49(4):1904– 1930, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-4/
Consistent-nonparametric-estimation-for-heavy-tailed-
sparse-graphs/10.1214/20-AOS1985.full.

Ba:2023:PIF

[BCCP23] Ismaïla Ba, Jean-François Coeurjolly, and Francisco Cuevas-Pacheco. Pairwise interaction function estimation of stationary Gibbs point processes using basis expansion. Annals of Statistics, 51(3):1134–1158, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-3/
Pairwise-interaction-function-estimation-of-stationary-
Gibbs-point-processes-using/10.1214/23-AOS2284.full.

Bates:2023:TOC

[BCL^{+}23] Stephen Bates, Emmanuel Candès, Lihua Lei, Yaniv Romano, and Matteo Sesia. Testing for outliers with conformal pvalues. Annals of Statistics, 51(1):149–178, February 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-1/Testing-for-outliers-
with-conformal-p-values/10.1214/22-AOS2244.full.

Belloni:2023:HDL

[BCPW23] Alexandre Belloni, Mingli Chen, Oscar Hernan Madrid Padilla, and Zixuan (Kevin) Wang. High-dimensional latent panel quantile regression with an application to asset pricing. Annals of Statistics, 51(1):96–121, February 2023. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-1/High-dimensional-latent-
panel-quantile-regression-with-an-application-to/10.
1214/22-AOS2223.full.

Barber:2021:PIJ

[BCRT21] Rina Foygel Barber, Emmanuel J. Candès, Aaditya Ramdas, and Ryan J. Tibshirani. Predictive inference with the jackknife+. Annals of Statistics, 49(1):486–507, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-1/Predictive-inference-
with-the-jackknife/10.1214/20-AOS1965.full.

Barber:2023:CPB

[BCRT23] Rina Foygel Barber, Emmanuel J. Candès, Aaditya Ramdas, and Ryan J. Tibshirani. Conformal prediction beyond exchangeability. Annals of Statistics, 51(2):816–845, April 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-2/Conformal-prediction-
beyond-exchangeability/10.1214/23-AOS2276.full.

Barber:2020:RIK

[BCS20] Rina Foygel Barber, Emmanuel J. Candès, and Richard J. Samworth. Robust inference with knockoffs. Annals of Statistics, 48 (3):1409–1431, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-3/
Robust-inference-with-knockoffs/10.1214/19-AOS1852.full.

Biau:2020:STP

[BCST20] Gérard Biau, Benoît Cadre, Maxime Sangnier, and Ugo Tanielian. Some theoretical properties of GANS. Annals of Statistics, 48(3):1539–1566, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
48/issue-3/Some-theoretical-properties-of-GANS/10.1214/
19-AOS1858.full.

Bagchi:2020:TSC

[BD20] Pramita Bagchi and Holger Dette. A test for separability in covariance operators of random surfaces. Annals of Statistics, 48 (4):2303–2322, August 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-4/
A-test-for-separability-in-covariance-operators-of-random-
surfaces/10.1214/19-AOS1888.full.

Bucher:2021:DGV

[BDH21] Axel Bücher, Holger Dette, and Florian Heinrichs. Are deviations in a gradually varying mean relevant? A testing approach based on sup-norm estimators. Annals of Statistics, 49(6): 3583–3617, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-6/
Are-deviations-in-a-gradually-varying-mean-relevant-
A-testing/10.1214/21-AOS2098.full.

Bastian:2024:TPS

[BDH24] Patrick Bastian, Holger Dette, and Johannes Heiny. Testing for practically significant dependencies in high dimensions via bootstrapping maxima of U-statistics. Annals of Statistics, 52 (2):628–653, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-2/
Testing-for-practically-significant-dependencies-in-
high-dimensions-via-bootstrapping/10.1214/24-AOS2361.
full.

Bhattacharya:2022:MES

[BDM22] Bhaswar B. Bhattacharya, Sayan Das, and Sumit Mukherjee. Motif estimation via subgraph sampling: The fourth-moment phenomenon. Annals of Statistics, 50(2):987–1011, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-2/Motif-estimation-via-
subgraph-sampling-The-fourth-moment-phenomenon/10.1214/
21-AOS2134.full.

Baek:2023:LWE

[BDP23] Changryong Baek, Marie-Christine Düker, and Vladas Pipiras. Local Whittle estimation of high-dimensional long-run variance and precision matrices. Annals of Statistics, 51(6): 2386–2414, December 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-6/

Local-Whittle-estimation-of-high-dimensional-long-run-
variance-and/10.1214/23-AOS2330.full.

Barber:2022:HTC

[BDSW22] Rina Foygel Barber, Mathias Drton, Nils Sturma, and Luca Weihs. Half-trek criterion for identifiability of latent variable models. Annals of Statistics, 50(6):3174–3196, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-6/Half-trek-criterion-
for-identifiability-of-latent-variable-models/10.1214/
22-AOS2221.full.

Bao:2021:SVS

[BDW21] Zhigang Bao, Xiucai Ding, and Ke Wang. Singular vector and singular subspace distribution for the matrix denoising model. Annals of Statistics, 49(1):370–392, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-1/Singular-vector-and-
singular-subspace-distribution-for-the-matrix-denoising/
10.1214/20-AOS1960.full.

Bao:2022:SIP

[BDWW22] Zhigang Bao, Xiucai Ding, Jingming Wang, and Ke Wang. Statistical inference for principal components of spiked covariance matrices. Annals of Statistics, 50(2):1144–1169, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-2/Statistical-inference-
for-principal-components-of-spiked-covariance-matrices/
10.1214/21-AOS2143.full.

Berthet:2022:CRW

[BE22] Philippe Berthet and John H. J. Einmahl. Cube root weak convergence of empirical estimators of a density level set. Annals of Statistics, 50(3):1423–1446, June 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-3/Cube-root-weak-convergence-
of-empirical-estimators-of-a-density/10.1214/21-AOS2157.
full.

Beyhum:2022:EPT

[BEPV22] Jad Beyhum, Anouar El Ghouch, François Portier, and Ingrid Van Keilegom. On an extension of the promotion time cure model. Annals of Statistics, 50(1):537– 559, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-1/
On-an-extension-of-the-promotion-time-cure-model/10.
1214/21-AOS2119.full.

Bravo:2020:TSS

[BEV20] Francesco Bravo, Juan Carlos Escanciano, and Ingrid Van Keilegom. Two-step semiparametric empirical likelihood inference. Annals of Statistics, 48(1):1–26, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

https://projecteuclid.org/euclid.aos/1581930123.

Bickel:2021:WVZ

[BFdGG21] Peter Bickel, Marta Fiocco, Mathisca de Gunst, and Friedrich Götze. Willem van Zwet’s research. Annals of Statistics, 49(5): 2439–2447, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-5/
Willem-van-Zwets-research/10.1214/21-AOS2060.full.

Brecheteau:2021:RBC

[BFL21] Claire Brécheteau, Aurélie Fischer, and Clément Levrard. Robust Bregman clustering. Annals of Statistics, 49(3): 1679–1701, June 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-3/
Robust-Bregman-clustering/10.1214/20-AOS2018.full.

Bongers:2021:FSC

[BFPM21] Stephan Bongers, Patrick Forré, Jonas Peters, and Joris M. Mooij. Foundations of structural causal models with cycles and latent variables. Annals of Statistics, 49(5):2885– 2915, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-5/
Foundations-of-structural-causal-models-with-cycles-
and-latent-variables/10.1214/21-AOS2064.full.

Bilodeau:2023:MRC

[BFR23] Blair Bilodeau, Dylan J. Foster, and Daniel M. Roy. Minimax rates for conditional density estimation via empirical entropy. Annals of Statistics, 51(2):762–790, April 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-51/issue-2/Minimax-rates-
for-conditional-density-estimation-via-empirical-entropy/
10.1214/23-AOS2270.full.

Bradic:2022:THD

[BFZ22] Jelena Bradic, Jianqing Fan, and Yinchu Zhu. Testability of high-dimensional linear models with nonsparse structures. Annals of Statistics, 50(2):615–639, April 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-2/Testability-of-high-
dimensional-linear-models-with-nonsparse-structures/
10.1214/19-AOS1932.full.

Belitser:2020:EBO

[BG20] Eduard Belitser and Subhashis Ghosal. Empirical Bayes oracle uncertainty quantification for regression. Annals of Statistics, 48(6):3113–3137, December 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
48/issue-6/Empirical-Bayes-oracle-uncertainty-quantification-
for-regression/10.1214/19-AOS1845.full.

Belomestny:2021:DDU

[BG21] Denis Belomestny and Alexander Goldenshluger. Density deconvolution under general assumptions on the distribution of measurement errors. Annals of Statistics, 49(2): 615–649, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-2/
Density-deconvolution-under-general-assumptions-on-the-
distribution-of-measurement/10.1214/20-AOS1969.full.

Bykhovskaya:2022:CLV

[BG22] Anna Bykhovskaya and Vadim Gorin. Cointegration in large VARs. Annals of Statistics, 50(3):1593–1617, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-50/issue-3/Cointegration-
in-large-VARs/10.1214/21-AOS2164.full.

Bunea:2020:MAV

[BGL^{+}20] Florentina Bunea, Christophe Giraud, Xi Luo, Martin Royer, and Nicolas Verzelen. Model assisted variable clustering: Minimaxoptimal recovery and algorithms. Annals of Statistics, 48(1): 111–137, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/euclid.aos/1581930128.

Belitser:2012:OTS

[BGvZ12] Eduard Belitser, Subhashis Ghosal, and Harry van Zanten. Optimal two-stage procedures for estimating location and size of the maximum of a multivariate regression function. Annals of Statistics, 40(6):2850–2876, December 2012. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:/

/projecteuclid.org/euclid.aos/1360332186. See correction

[BGvZ21].

Belitser:2021:CNS

[BGvZ21] Eduard Belitser, Subhashis Ghosal, and Harry van Zanten. Correction note: “Optimal two-stage procedures for estimating location and size of the maximum of a multivariate regression function” Ann. Statist. 40 (2012) 2850–2876. Annals of Statistics, 49 (1):612–613, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-1/
Correction-note--Optimal-two-stage-procedures-for-estimating-
location/10.1214/20-AOS1993.full. See [BGvZ12].

Bhattacharya:2020:ADD

[Bha20] Bhaswar B. Bhattacharya. Asymptotic distribution and detection thresholds for two-sample tests based on geometric graphs. Annals of Statistics, 48(5):2879–2903, October 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-5/Asymptotic-distribution-
and-detection-thresholds-for-two-sample-tests-based/
10.1214/19-AOS1913.full.

Barber:2022:TGF

[BJ22] Rina Foygel Barber and Lucas Janson. Testing goodness-of-fit and conditional independence with approximate co-sufficient sampling. Annals of Statistics, 50(5):2514–2544, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-5/Testing-goodness-of-
fit-and-conditional-independence-with-approximate-co/
10.1214/22-AOS2187.full.

Blanchard:2023:URA

[BJ23] Moïse Blanchard and Patrick Jaillet. Universal regression with adversarial responses. Annals of Statistics, 51(3):1401– 1426, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-3/
Universal-regression-with-adversarial-responses/10.1214/
23-AOS2299.full.

Bradic:2024:HDI

[BJZ24] Jelena Bradic, Weijie Ji, and Yuqian Zhang. High-dimensional inference for dynamic treatment effects. Annals of Statistics, 52(2): 415–440, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-52/issue-2/High-
dimensional-inference-for-dynamic-treatment-effects/
10.1214/24-AOS2352.full.

Bresler:2020:LTS

[BK20] Guy Bresler and Mina Karzand. Learning a tree-structured Ising model in order to make predictions. Annals of Statistics, 48 (2):713–737, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/euclid.aos/1590480031.

Bu:2023:CST

[BKRS23] Zhiqi Bu, Jason M. Klusowski, Cynthia Rush, and Weijie J. Su. Characterizing the SLOPE trade-off: A variational perspective and the Donoho–Tanner limit. Annals of Statistics, 51 (1):33–61, February 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-1/

Characterizing-the-SLOPE-trade-off--A-variational-perspective-
and/10.1214/22-AOS2194.full.

Berrett:2021:ORI

[BKS21] Thomas B. Berrett, Ioannis Kontoyiannis, and Richard J. Samworth. Optimal rates for independence testing via U-statistic permutation tests. Annals of Statistics, 49(5):2457–2490, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-49/issue-5/Optimal-rates-
for-independence-testing-via-U-statistic-permutation-
tests/10.1214/20-AOS2041.full.

Banerjee:2022:OSD

[BM22] Debapratim Banerjee and Zongming Ma. Optimal signal detection in some spiked random matrix models: Likelihood ratio tests and linear spectral statistics. Annals of Statistics, 50(4): 1910–1932, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-4/
Optimal-signal-detection-in-some-spiked-random-matrix-
models/10.1214/21-AOS2150.full.

Bhattacharjee:2023:SIF

[BM23] Satarupa Bhattacharjee and Hans-Georg Müller. Single index Fréchet regression. Annals of Statistics, 51(4):1770–1798, August 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-51/issue-4/Single-index-
Fr%c3%a9chet-regression/10.1214/23-AOS2307.full.

Butucea:2023:VSM

[BMNT23] Cristina Butucea, Enno Mammen, Mohamed Ndaoud, and Alexandre B. Tsybakov. Variable selection, monotone likelihood ratio and group sparsity. Annals of Statistics, 51(1): 312–333, February 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-1/
Variable-selection-monotone-likelihood-ratio-and-group-
sparsity/10.1214/22-AOS2251.full.

Bu:2020:ODM

[BMY20] Xianwei Bu, Dibyen Majumdar, and Jie Yang. D-optimal designs for multinomial logistic models. Annals of Statistics, 48 (2):983–1000, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/euclid.aos/1590480042.

Blanchard:2020:PHC

[BNR20] Gilles Blanchard, Pierre Neuvial, and Etienne Roquain. Post hoc confidence bounds on false positives using reference families. Annals of Statistics, 48(3):1281–1303, June 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-3/Post-hoc-confidence-
bounds-on-false-positives-using-reference-families/10.
1214/19-AOS1847.full.

Bilodeau:2023:RDA

[BNR23] Blair Bilodeau, Jeffrey Negrea, and Daniel M. Roy. Relaxing the i.i.d. assumption: Adaptively minimax optimal regret via root-entropic regularization. Annals of Statistics, 51(4): 1850–1876, August 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-4/
Relaxing-the-iid-assumption--Adaptively-minimax-optimal-
regret-via/10.1214/23-AOS2315.full.

Bing:2022:AEM

[BNX22] Xin Bing, Yang Ning, and Yaosheng Xu. Adaptive estimation in multivariate response regression with hidden variables. Annals of Statistics, 50(2):640–672, April 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-2/Adaptive-estimation-
in-multivariate-response-regression-with-hidden-variables/
10.1214/21-AOS2059.full.

Bucher:2024:TIH

[BP24] Axel Bücher and Cambyse Pakzad. Testing for independence in high dimensions based on empirical copulas. Annals of Statistics, 52(1):311–334, February 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
52/issue-1/Testing-for-independence-in-high-dimensions-
based-on-empirical-copulas/10.1214/23-AOS2348.full.

Bachoc:2020:UVC

[BPS20] François Bachoc, David Preinerstorfer, and Lukas Steinberger. Uniformly valid confidence intervals post-model-selection. Annals of Statistics, 48(1):440–463, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

https://projecteuclid.org/euclid.aos/1581930142.

Basu:2023:GMN

[BR23] Sumanta Basu and Suhasini Subba Rao. Graphical models for nonstationary time series. Annals of Statistics, 51(4): 1453–1483, August 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-4/
Graphical-models-for-nonstationary-time-series/10.1214/
22-AOS2205.full.

Brutsche:2024:SAP

[BR24] Johannes Brutsche and Angelika Rohde. Sharp adaptive and pathwise stable similarity testing for scalar ergodic diffusions. Annals of Statistics, 52(3):1127–1151, June 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-3/Sharp-adaptive-and-pathwise-
stable-similarity-testing-for-scalar-ergodic/10.1214/
24-AOS2386.full.

Butucea:2023:IVN

[BRS23] Cristina Butucea, Angelika Rohde, and Lukas Steinberger. Interactive versus noninteractive locally differentially private estimation: Two elbows for the quadratic functional. Annals of Statistics, 51(2):464–486, April 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
51/issue-2/Interactive-versus-noninteractive-locally-
differentially-private-estimation--Two-elbows/10.1214/
22-AOS2254.full.

Berg:2023:ESC

[BS23a] Stephen Berg and Hyebin Song. Efficient shape-constrained inference for the autocovariance sequence from a reversible Markov chain. Annals of Statistics, 51(6):2440–2470, December 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-

of-statistics/volume-51/issue-6/Efficient-shape-constrained-
inference-for-the-autocovariance-sequence-from-a/10.
1214/23-AOS2335.full.

Berrett:2023:EFE

[BS23b] Thomas B. Berrett and Richard J. Samworth. Efficient functional estimation and the super-oracle phenomenon. Annals of Statistics, 51(2):668–690, April 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
51/issue-2/Efficient-functional-estimation-and-the-super-
oracle-phenomenon/10.1214/23-AOS2265.full.

Berrett:2023:ONT

[BS23c] Thomas B. Berrett and Richard J. Samworth. Optimal nonparametric testing of missing completely at random and its connections to compatibility. Annals of Statistics, 51(5):2170– 2193, October 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-5/
Optimal-nonparametric-testing-of-Missing-Completely-
At-Random-and-its/10.1214/23-AOS2326.full.

Berger:2020:BAC

[BSS20] James O. Berger, Dongchu Sun, and Chengyuan Song. Bayesian analysis of the covariance matrix of a multivariate normal distribution with a new class of priors. Annals of Statistics, 48(4):2381–2403, August 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-4/Bayesian-analysis-of-
the-covariance-matrix-of-a-multivariate-normal/10.1214/
19-AOS1891.full.

Bertsimas:2020:SHD

[BV20] Dimitris Bertsimas and Bart Van Parys. Sparse high-dimensional regression: Exact scalable algorithms and phase transitions. Annals of Statistics, 48(1):300–323, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

https://projecteuclid.org/euclid.aos/1581930136.

Bing:2023:ODA

[BW23] Xin Bing and Marten Wegkamp. Optimal discriminant analysis in high-dimensional latent factor models. Annals of Statistics, 51(3):1232–1257, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-3/
Optimal-discriminant-analysis-in-high-dimensional-latent-
factor-models/10.1214/23-AOS2289.full.

Bellec:2021:SOS

[BZ21] Pierre C. Bellec and Cun-Hui Zhang. Second-order Stein: SURE for SURE and other applications in high-dimensional inference. Annals of Statistics, 49(4):1864–1903, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-49/issue-4/Second-order-
Stein--SURE-for-SURE-and-other-applications/10.1214/
20-AOS2005.full.

Bellec:2023:DCR

[BZ23a] Pierre C. Bellec and Cun-Hui Zhang. Debiasing convex regularized estimators and interval estimation in linear models. Annals of Statistics, 51(2):391–436, April 2023. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-2/Debiasing-convex-regularized-
estimators-and-interval-estimation-in-linear-models/
10.1214/22-AOS2243.full.

Bucher:2023:DSB

[BZ23b] Axel Bücher and Leandra Zanger. On the disjoint and sliding block maxima method for piecewise stationary time series. Annals of Statistics, 51(2):573–598, April 2023. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-2/On-the-disjoint-and-
sliding-block-maxima-method-for-piecewise/10.1214/23-
AOS2260.full.

Cannings:2020:LNN

[CBS20] Timothy I. Cannings, Thomas B. Berrett, and Richard J. Samworth. Local nearest neighbour classification with applications to semi-supervised learning. Annals of Statistics, 48(3):1789–1814, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-48/issue-3/Local-nearest-

neighbour-classification-with-applications-to-semi-supervised-
learning/10.1214/19-AOS1868.full.

Chaudhuri:2023:CVF

[CC23] Anamitra Chaudhuri and Sabyasachi Chatterjee. A crossvalidation framework for signal denoising with applications to trend filtering, dyadic CART and beyond. Annals of Statistics, 51 (4):1534–1560, August 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-4/
A-cross-validation-framework-for-signal-denoising-with-
applications-to/10.1214/23-AOS2283.full.

Chen:2021:AHE

[CCF21] Yuxin Chen, Chen Cheng, and Jianqing Fan. Asymmetry helps: Eigenvalue and eigenvector analyses of asymmetrically perturbed low-rank matrices. Annals of Statistics, 49(1):435– 458, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-1/
Asymmetry-helps--Eigenvalue-and-eigenvector-analyses-
of-asymmetrically-perturbed/10.1214/20-AOS1963.full.

Cattaneo:2024:CRO

[CCK24] Matias D. Cattaneo, Rajita Chandak, and Jason M. Klusowski. Convergence rates of oblique regression trees for flexible function libraries. Annals of Statistics, 52(2):466–490, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-2/Convergence-rates-of-
oblique-regression-trees-for-flexible-function-libraries/
10.1214/24-AOS2354.full.

Chernozhuokov:2022:ICL

[CCKK22] Victor Chernozhuokov, Denis Chetverikov, Kengo Kato, and Yuta Koike. Improved central limit theorem and bootstrap approximations in high dimensions. Annals of Statistics, 50(5): 2562–2586, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-5/
Improved-central-limit-theorem-and-bootstrap-approximations-
in-high-dimensions/10.1214/22-AOS2193.full.

Cai:2024:TLC

[CCL24] Changxiao Cai, T. Tony Cai, and Hongzhe Li. Transfer learning for contextual multi-armed bandits. Annals of Statistics, 52 (1):207–232, February 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-1/
Transfer-learning-for-contextual-multi-armed-bandits/
10.1214/23-AOS2341.full.

Comminges:2021:ARE

[CCNT21] L. Comminges, O. Collier, M. Ndaoud, and A. B. Tsybakov. Adaptive robust estimation in sparse vector model. Annals of Statistics, 49(3):1347–1377, June 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-3/Adaptive-robust-estimation-
in-sparse-vector-model/10.1214/20-AOS2002.full.

Cao:2022:OFD

[CCZ22] Hongyuan Cao, Jun Chen, and Xianyang Zhang. Optimal false discovery rate control for large scale multiple testing with auxiliary information. Annals of Statistics, 50(2): 807–857, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-2/
Optimal-false-discovery-rate-control-for-large-scale-
multiple-testing/10.1214/21-AOS2128.full.

Cai:2024:EIM

[CCZ24] T. Tony Cai, Ran Chen, and Yuancheng Zhu. Estimation and inference for minimizer and minimum of convex functions: Optimality, adaptivity and uncertainty principles. Annals of Statistics, 52(1):392–411, February 2024. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-1/Estimation-and-inference-
for-minimizer-and-minimum-of-convex-functions/10.1214/
24-AOS2355.full.

Carpentier:2021:EME

[CDRV21] Alexandra Carpentier, Sylvain Delattre, Etienne Roquain, and Nicolas Verzelen. Estimating minimum effect with outlier selection. Annals of Statistics, 49(1):272–294, February 2021.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-1/Estimating-minimum-effect-
with-outlier-selection/10.1214/20-AOS1956.full.

Cattaneo:2020:LSP

[CFF20] Matias D. Cattaneo, Max H. Farrell, and Yingjie Feng. Large sample properties of partitioning-based series estimators. Annals of Statistics, 48(3):1718–1741, June 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-3/Large-sample-properties-
of-partitioning-based-series-estimators/10.1214/19-AOS1865.
full.

Celentano:2023:LCT

[CFM23] Michael Celentano, Zhou Fan, and Song Mei. Local convexity of the TAP free energy and AMP convergence for. Annals of Statistics, 51(2):519–546, April 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
51/issue-2/Local-convexity-of-the-TAP-free-energy-and-
AMP-convergence/10.1214/23-AOS2257.full.

Chen:2021:BCN

[CFMY21] Yuxin Chen, Jianqing Fan, Cong Ma, and Yuling Yan. Bridging convex and nonconvex optimization in robust PCA: Noise, outliers and missing data. Annals of Statistics, 49(5):2948– 2971, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-5/
Bridging-convex-and-nonconvex-optimization-in-robust-
PCA--Noise/10.1214/21-AOS2066.full.

Chatelain:2020:IAC

[CFN20] Simon Chatelain, Anne-Laure Fougères, and Johanna G. Neslehová. Inference for Archimax copulas. Annals of Statistics, 48 (2):1025–1051, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/euclid.aos/1590480044.

Camerlenghi:2021:ODR

[CFNP21] Federico Camerlenghi, Stefano Favaro, Zacharie Naulet, and Francesca Panero. Optimal disclosure risk assessment. Annals of Statistics, 49(2):723–744, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
49/issue-2/Optimal-disclosure-risk-assessment/10.1214/
20-AOS1975.full.

Chakraborty:2021:CCI

[CG21a] Moumita Chakraborty and Subhashis Ghosal. Coverage of credible intervals in nonparametric monotone regression. Annals of Statistics, 49(2):1011–1028, April 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-2/Coverage-of-credible-
intervals-in-nonparametric-monotone-regression/10.1214/
20-AOS1989.full.

Chatterjee:2021:AEM

[CG21b] Sabyasachi Chatterjee and Subhajit Goswami. Adaptive estimation of multivariate piecewise polynomials and bounded variation functions by optimal decision trees. Annals of Statistics, 49(5): 2531–2551, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-5/
Adaptive-estimation-of-multivariate-piecewise-polynomials-
and-bounded-variation-functions/10.1214/20-AOS2045.full.

Comte:2020:NDE

[CGC20] Fabienne Comte and Valentine Genon-Catalot. Nonparametric drift estimation for i.i.d. paths of stochastic differential equations. Annals of Statistics, 48(6):3336–3365, December 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-6/Nonparametric-drift-
estimation-for-iid-paths-of-stochastic-differential-
equations/10.1214/19-AOS1933.full.

Chronopoulos:2022:CBP

[CGK22] Ilias Chronopoulos, Liudas Giraitis, and George Kapetanios. Choosing between persistent and stationary volatility. Annals of Statistics, 50(6):3466–3483, December 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-6/Choosing-between-persistent-
and-stationary-volatility/10.1214/22-AOS2236.full.

Chen:2022:OFR

[CGZ22a] Pinhan Chen, Chao Gao, and Anderson Y. Zhang. Optimal full ranking from pairwise comparisons. Annals of Statistics, 50 (3):1775–1805, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-3/
Optimal-full-ranking-from-pairwise-comparisons/10.1214/
22-AOS2175.full.

Chen:2022:PRT

[CGZ22b] Pinhan Chen, Chao Gao, and Anderson Y. Zhang. Partial recovery for top-k ranking: Optimality of MLE and SubOptimality of the spectral method. Annals of Statistics, 50(3): 1618–1652, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-3/
Partial-recovery-for-top-k-ranking--Optimality-of-MLE/
10.1214/21-AOS2166.full.

Chan:2020:MAB

[Cha20] Hock Peng Chan. The multi-armed bandit problem: An efficient nonparametric solution. Annals of Statistics, 48(1):346–373, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL https://projecteuclid.org/euclid.
aos/1581930138.

Chan:2022:GFT

[Cha22a] Kin Wai Chan. General and feasible tests with multiply-imputed datasets. Annals of Statistics, 50(2):930–948, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-2/General-and-feasible-
tests-with-multiply-imputed-datasets/10.1214/21-AOS2132.
full.

Chan:2022:ODB

[Cha22b] Kin Wai Chan. Optimal difference-based variance estimators in time series: a general framework. Annals of Statistics, 50 (3):1376–1400, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-3/
Optimal-difference-based-variance-estimators-in-time-
series--A/10.1214/21-AOS2154.full.

Chen:2021:SMN

[Che21] Louis H. Y. Chen. Stein’s method of normal approximation: Some recollections and reflections. Annals of Statistics, 49(4): 1850–1863, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-4/
Steins-method-of-normal-approximation-Some-recollections-
and-reflections/10.1214/21-AOS2083.full.

Chen:2022:PGG

[Che22] Yen-Chi Chen. Pattern graphs: a graphical approach to nonmonotone missing data. Annals of Statistics, 50(1):129– 146, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-1/
Pattern-graphs-A-graphical-approach-to-nonmonotone-missing-
data/10.1214/21-AOS2094.full.

Chetelat:2024:EIM

[Ché24] Didier Chételat. Erratum: Improved multivariate normal mean estimation with unknown covariance when p is greater than n. Annals of Statistics, 52(1):412, February 2024. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-1/Erratum--Improved-multivariate-
normal-mean-estimation-with-unknown-covariance/10.1214/
23-AOS2344.full.

Chernozhukov:2021:LDI

[CHHW21] Victor Chernozhukov, Wolfgang Karl Härdle, Chen Huang, and Weining Wang. LASSO-driven inference in time and space. Annals of Statistics, 49(3):1702–1735, June 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-3/LASSO-driven-inference-
in-time-and-space/10.1214/20-AOS2019.full.

Chang:2024:EDP

[CHK^{+}24] Jinyuan Chang, Qiao Hu, Eric D. Kolaczyk, Qiwei Yao, and Fengting Yi. Edge differentially private estimation in the βmodel via jittering and method of moments. Annals of Statistics, 52(2):708–728, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-2/
Edge-differentially-private-estimation-in-the-%ce%b2-
model-via-jittering/10.1214/24-AOS2365.full.

Chong:2024:SIR

[CHL^{+}24] Carsten H. Chong, Marc Hoffmann, Yanghui Liu, Mathieu Rosenbaum, and Grégoire Szymansky. Statistical inference for rough volatility: Minimax theory. Annals of Statistics, 52(4): 1277–1306, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-4/
Statistical-inference-for-rough-volatility-Minimax-theory/
10.1214/23-AOS2343.full.

Chernozhukov:2023:ILR

[CHLZ23] Victor Chernozhukov, Christian Hansen, Yuan Liao, and Yinchu Zhu. Inference for low-rank models. Annals of Statistics, 51 (3):1309–1330, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-3/
Inference-for-low-rank-models/10.1214/23-AOS2293.full.

Chong:2020:HFA

[Cho20] Carsten Chong. High-frequency analysis of parabolic stochastic PDEs. Annals of Statistics, 48(2):1143–1167, April 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL https://projecteuclid.org/euclid.aos/1590480049.

Cai:2020:LLD

[CHP20] T. Tony Cai, Xiao Han, and Guangming Pan. Limiting laws for divergent spiked eigenvalues and largest nonspiked eigenvalue of sample covariance matrices. Annals of Statistics, 48 (3):1255–1280, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-3/
Limiting-laws-for-divergent-spiked-eigenvalues-and-largest-
nonspiked-eigenvalue/10.1214/18-AOS1798.full.

Chen:2024:DEI

[CJLZ24] Xi Chen, Wenbo Jing, Weidong Liu, and Yichen Zhang. Distributed estimation and inference for semiparametric binary response models. Annals of Statistics, 52(3):922–

947, June 2024.  CODEN ASTSC7.  ISSN 0090-5364

(print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-3/
Distributed-estimation-and-inference-for-semiparametric-
binary-response-models/10.1214/24-AOS2376.full.

Cattaneo:2024:BAI

[CJN24] Matias D. Cattaneo, Michael Jansson, and Kenichi Nagasawa. Bootstrap-assisted inference for generalized grenander-type estimators. Annals of Statistics, 52(4):1509–1533, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-4/Bootstrap-assisted-inference-
for-generalized-Grenander-type-estimators/10.1214/24-
AOS2402.full.

Chen:2024:DNS

[CJQY24] YinFeng Chen, YuLing Jiao, Rui Qiu, and Zhou Yu. Deep nonlinear sufficient dimension reduction. Annals of Statistics, 52 (3):1201–1226, June 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-3/
Deep-nonlinear-sufficient-dimension-reduction/10.1214/
24-AOS2390.full.

Cai:2024:TLF

[CKP24] T. Tony Cai, Dongwoo Kim, and Hongming Pu. Transfer learning for functional mean estimation: Phase transition and adaptive algorithms. Annals of Statistics, 52(2): 654–678, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-2/
Transfer-learning-for-functional-mean-estimation--Phase-
transition-and/10.1214/24-AOS2362.full.

Cheysson:2022:SEH

[CL22] Felix Cheysson and Gabriel Lang. Spectral estimation of Hawkes processes from count data. Annals of Statistics, 50 (3):1722–1746, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-3/
Spectral-estimation-of-Hawkes-processes-from-count-data/
10.1214/22-AOS2173.full.

Chen:2024:RFL

[CL24] Weilin Chen and Clifford Lam. Rank and factor loadings estimation in time series tensor factor model by pre-averaging. Annals of Statistics, 52(1):364–391, February 2024. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-1/Rank-and-factor-loadings-
estimation-in-time-series-tensor-factor/10.1214/23-AOS2350.
full.

Cai:2021:SEU

[CLC^{+}21a] Changxiao Cai, Gen Li, Yuejie Chi, H. Vincent Poor, and Yuxin Chen. Subspace estimation from unbalanced and incomplete data matrices: ℓ_{2,∞} statistical guarantees. Annals of Statistics, 49(2):944–967, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-2/
Subspace-estimation-from-unbalanced-and-incomplete-data-
matrices--%e2%84%932/10.1214/20-AOS1986.full.

Chetverikov:2021:CVL

[CLC21b] Denis Chetverikov, Zhipeng Liao, and Victor Chernozhukov. On cross-validated Lasso in high dimensions. Annals of Statistics, 49(3):1300–1317, June 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-3/
On-cross-validated-Lasso-in-high-dimensions/10.1214/
20-AOS2000.full.

Chen:2022:GIC

[CLM22] Shuxiao Chen, Sifan Liu, and Zongming Ma. Global and individualized community detection in inhomogeneous multilayer networks. Annals of Statistics, 50(5):2664–2693, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-5/Global-and-individualized-
community-detection-in-inhomogeneous-multilayer-networks/
10.1214/22-AOS2202.full.

Camerlenghi:2021:SAH

[CLP21a] Federico Camerlenghi, Antonio Lijoi, and Igor Prünster. Survival analysis via hierarchically dependent mixture hazards. Annals of Statistics, 49(2):863–884, April 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-2/Survival-analysis-via-
hierarchically-dependent-mixture-hazards/10.1214/20-
AOS1982.full.

Catalano:2021:MDW

[CLP21b] Marta Catalano, Antonio Lijoi, and Igor Prünster. Measuring dependence in the Wasserstein distance for Bayesian nonparametric models. Annals of Statistics, 49(5):2916– 2947, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-5/
Measuring-dependence-in-the-Wasserstein-distance-for-
Bayesian-nonparametric-models/10.1214/21-AOS2065.full.

Chen:2020:SIM

[CLTZ20] Xi Chen, Jason D. Lee, Xin T. Tong, and Yichen Zhang. Statistical inference for model parameters in stochastic gradient descent. Annals of Statistics, 48(1):251–273, February 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL https://projecteuclid.org/euclid.aos/1581930134.

Chinot:2022:RMN

[CLvdG22] Geoffrey Chinot, Matthias Löffler, and Sara van de Geer. On the robustness of minimum norm interpolators and regularized empirical risk minimizers. Annals of Statistics, 50(4): 2306–2333, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-4/
On-the-robustness-of-minimum-norm-interpolators-and-
regularized-empirical/10.1214/22-AOS2190.full.

Caponera:2021:ASF

[CM21] Alessia Caponera and Domenico Marinucci. Asymptotics for spherical functional autoregressions. Annals of Statistics, 49(1): 346–369, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-1/
Asymptotics-for-spherical-functional-autoregressions/
10.1214/20-AOS1959.full.

Celentano:2022:FBH

[CM22a] Michael Celentano and Andrea Montanari. Fundamental barriers to high-dimensional regression with convex penalties. Annals of Statistics, 50(1):170–196, February 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-1/Fundamental-barriers-
to-high-dimensional-regression-with-convex-penalties/
10.1214/21-AOS2100.full.

Chen:2022:UCL

[CM22b] Yaqing Chen and Hans-Georg Müller. Uniform convergence of local Fréchet regression with applications to locating extrema and time warping for metric space valued trajectories. Annals of Statistics, 50(3):1573–1592, June 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-3/Uniform-convergence-
of-local-Fr%c3%a9chet-regression-with-applications-to-
locating/10.1214/21-AOS2163.full.

Celentano:2023:LGG

[CMW23] Michael Celentano, Andrea Montanari, and Yuting Wei. The lasso with general Gaussian designs with applications to hypothesis testing. Annals of Statistics, 51(5):2194–2220, October 2023. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-51/issue-5/The-
Lasso-with-general-Gaussian-designs-with-applications-
to-hypothesis/10.1214/23-AOS2327.full.

Chan:2021:OCP

[CNYY21] Ngai Hang Chan, Wai Leong Ng, Chun Yip Yau, and Haihan Yu. Optimal change-point estimation in time series. Annals of Statistics, 49(4):2336–2355, August 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-4/Optimal-change-point-
estimation-in-time-series/10.1214/20-AOS2039.full.

Cox:2020:ASU

[Cox20] Gregory Cox. Almost sure uniqueness of a global minimum without convexity. Annals of Statistics, 48(1):584–606, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/euclid.aos/
1581930148.

Chandler:2021:MGF

[CP21a] Gabriel Chandler and Wolfgang Polonik. Multiscale geometric feature extraction for high-dimensional and non-Euclidean data with applications. Annals of Statistics, 49(2):988– 1010, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-2/
Multiscale-geometric-feature-extraction-for-high-dimensional-
and-non-Euclidean/10.1214/20-AOS1988.full.

Chen:2021:DSI

[CP21b] Song Xi Chen and Liuhua Peng. Distributed statistical inference for massive data. Annals of Statistics, 49(5):2851– 2869, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-5/
Distributed-statistical-inference-for-massive-data/10.
1214/21-AOS2062.full.

Cai:2022:SCA

[CP22a] T. Tony Cai and Hongming Pu. Stochastic continuum-armed bandits with additive models: Minimax regrets and adaptive algorithm. Annals of Statistics, 50(4):2179–2204, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-4/Stochastic-continuum-
armed-bandits-with-additive-models--Minimax-regrets/
10.1214/22-AOS2182.full.

Charkaborty:2022:TRC

[CP22b] Anirvan Charkaborty and Victor M. Panaretos. Testing for the rank of a covariance operator. Annals of Statistics, 50(6): 3510–3537, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-6/
Testing-for-the-rank-of-a-covariance-operator/10.1214/
22-AOS2238.full.

Christgau:2023:NCL

[CPH23] Alexander Mangulad Christgau, Lasse Petersen, and Niels Richard Hansen. Nonparametric conditional local independence testing. Annals of Statistics, 51(5):2116–2144, October 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-5/Nonparametric-conditional-
local-independence-testing/10.1214/23-AOS2323.full.

Chandrasekher:2023:SGC

[CPT23] Kabir Aladin Chandrasekher, Ashwin Pananjady, and Christos Thrampoulidis. Sharp global convergence guarantees for iterative nonconvex optimization with random data. Annals of Statistics, 51(1):179–210, February 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
51/issue-1/Sharp-global-convergence-guarantees-for-iterative-
nonconvex-optimization-with-random/10.1214/22-AOS2246.
full.

Chen:2023:SOH

[CQZ23] Song Xi Chen, Yumou Qiu, and Shuyi Zhang. Sharp optimality for high-dimensional covariance testing under sparse signals. Annals of Statistics, 51(5):1921–1945, October 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-5/Sharp-optimality-for-
high-dimensional-covariance-testing-under-sparse-signals/
10.1214/23-AOS2310.full.

Castillo:2020:SSE

[CR20] Ismaël Castillo and Étienne Roquain. On spike and slab empirical Bayes multiple testing. Annals of Statistics, 48(5): 2548–2574, October 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-5/
On-spike-and-slab-empirical-Bayes-multiple-testing/10.
1214/19-AOS1897.full.

Castillo:2021:UQB

[CR21] Ismaël Castillo and Veronika Rocková. Uncertainty quantification for Bayesian CART. Annals of Statistics, 49(6):3482– 3509, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-6/
Uncertainty-quantification-for-Bayesian-CART/10.1214/
21-AOS2093.full.

Candes:2020:PTE

[CS20] Emmanuel J. Candès and Pragya Sur. The phase transition for the existence of the maximum likelihood estimate in highdimensional logistic regression. Annals of Statistics, 48(1):27–42, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL https://projecteuclid.org/euclid.
aos/1581930124.

Chzhen:2022:MFQ

[CS22] Evgenii Chzhen and Nicolas Schreuder. A minimax framework for quantifying risk-fairness trade-off in regression. Annals of Statistics, 50(4):2416–2442, August 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-4/A-minimax-framework-
for-quantifying-risk-fairness-trade-off-in/10.1214/22-
AOS2198.full.

Chopin:2022:RSP

[CSSV22] Nicolas Chopin, Sumeetpal S. Singh, Tomás Soto, and Matti Vihola. On resampling schemes for particle filters with weakly informative observations. Annals of Statistics, 50(6):3197– 3222, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-6/
On-resampling-schemes-for-particle-filters-with-weakly-
informative-observations/10.1214/22-AOS2222.full.

Chen:2022:SOA

[CT22] Guanzhou Chen and Boxin Tang. A study of orthogonal array-based designs under a broad class of space-filling criteria. Annals of Statistics, 50(5):2925–2949, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-5/A-study-of-orthogonal-
array-based-designs-under-a-broad/10.1214/22-AOS2215.
full.

Castillo:2021:MBS

[CvdP21] Ismaël Castillo and Stéphanie van der Pas. Multiscale Bayesian survival analysis. Annals of Statistics, 49(6):3559– 3582, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-6/
Multiscale-Bayesian-survival-analysis/10.1214/21-AOS2097.
full.

Chi:2022:APH

[CVFL22] Chien-Ming Chi, Patrick Vossler, Yingying Fan, and Jinchi Lv. Asymptotic properties of high-dimensional random forests. Annals of Statistics, 50(6):3415–3438, December 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-6/Asymptotic-properties-
of-high-dimensional-random-forests/10.1214/22-AOS2234.
full.

Cai:2020:SCL

[CW20] T. Tony Cai and Yihong Wu. Statistical and computational limits for sparse matrix detection. Annals of Statistics, 48(3): 1593–1614, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-3/
Statistical-and-computational-limits-for-sparse-matrix-
detection/10.1214/19-AOS1860.full.

Cai:2021:TLN

[CW21] T. Tony Cai and Hongji Wei. Transfer learning for nonparametric classification: Minimax rate and adaptive classifier. Annals of Statistics, 49(1):100–128, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-1/Transfer-learning-for-
nonparametric-classification--Minimax-rate-and-adaptive/
10.1214/20-AOS1949.full.

Cai:2022:DAG

[CW22a] T. Tony Cai and Hongji Wei. Distributed adaptive Gaussian mean estimation with unknown variance: Interactive protocol helps adaptation. Annals of Statistics, 50(4):1992– 2020, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-4/
Distributed-adaptive-Gaussian-mean-estimation-with-unknown-
variance--Interactive/10.1214/21-AOS2167.full.

Cai:2022:DNF

[CW22b] T. Tony Cai and Hongji Wei. Distributed nonparametric function estimation: Optimal rate of convergence and cost of adaptation. Annals of Statistics, 50(2):698–725, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-2/Distributed-nonparametric-
function-estimation--Optimal-rate-of-convergence-and/
10.1214/21-AOS2124.full.

Cai:2021:CPO

[CWZ21] T. Tony Cai, Yichen Wang, and Linjun Zhang. The cost of privacy: Optimal rates of convergence for parameter estimation with differential privacy. Annals of Statistics, 49(5):2825– 2850, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-49/issue-5/The-
cost-of-privacy--Optimal-rates-of-convergence-for/10.
1214/21-AOS2058.full.

Chen:2020:RIM

[CZ20] Xi Chen and Wen-Xin Zhou. Robust inference via multiplier bootstrap. Annals of Statistics, 48(3):1665–1691, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-3/Robust-inference-via-
multiplier-bootstrap/10.1214/19-AOS1863.full.

Cai:2021:COA

[CZ21] T. Tony Cai and Linjun Zhang. A convex optimization approach to high-dimensional sparse quadratic discriminant analysis. Annals of Statistics, 49(3):1537–1568, June 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-3/A-convex-optimization-
approach-to-high-dimensional-sparse-quadratic-discriminant/
10.1214/20-AOS2012.full.

delAlamo:2021:FCT

[dÁLM21] Miguel del Álamo, Housen Li, and Axel Munk. Frameconstrained total variation regularization for white noise regression. Annals of Statistics, 49(3):1318–1346, June 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-3/Frame-constrained-total-
variation-regularization-for-white-noise-regression/
10.1214/20-AOS2001.full.

Dau:2023:BSA

[DC23] Hai-Dang Dau and Nicolas Chopin. On backward smoothing algorithms. Annals of Statistics, 51(5):2145–2169, October 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-5/On-backward-smoothing-
algorithms/10.1214/23-AOS2324.full.

Dubey:2024:MSE

[DCM24] Paromita Dubey, Yaqing Chen, and Hans-Georg Müller. Metric statistics: Exploration and inference for random objects with distance profiles. Annals of Statistics, 52(2): 757–792, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-2/
Metric-statistics--Exploration-and-inference-for-random-
objects-with/10.1214/24-AOS2368.full.

DiBenedetto:2021:NRP

[DCT21] Giuseppe Di Benedetto, François Caron, and Yee Whye Teh. Nonexchangeable random partition models for microclustering. Annals of Statistics, 49(4):1931–1957, August 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-4/Nonexchangeable-random-
partition-models-for-microclustering/10.1214/20-AOS2003.
full.

Ding:2023:MRT

[DD23] Jian Ding and Hang Du. Matching recovery threshold for correlated random graphs. Annals of Statistics, 51(4):1718– 1743, August 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-4/
Matching-recovery-threshold-for-correlated-random-graphs/
10.1214/23-AOS2305.full.

Dieuleveut:2020:BGB

[DDB20] Aymeric Dieuleveut, Alain Durmus, and Francis Bach. Bridging the gap between constant step size stochastic gradient descent and Markov chains. Annals of Statistics, 48(3):1348–1382, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-48/issue-3/Bridging-the-
gap-between-constant-step-size-stochastic-gradient-descent/
10.1214/19-AOS1850.full.

Davezies:2021:EPR

[DDG21] Laurent Davezies, Xavier D’Haultfoeuille, and Yannick Guyonvarch. Empirical process results for exchangeable arrays. Annals of Statistics, 49(2):845–862, April 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-2/Empirical-process-results-
for-exchangeable-arrays/10.1214/20-AOS1981.full.

Davis:2024:ANO

[DDJ24] Damek Davis, Dmitriy Drusvyatskiy, and Liwei Jiang. Asymptotic normality and optimality in nonsmooth stochastic approximation. Annals of Statistics, 52(4):1485–1508, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-4/Asymptotic-normality-
and-optimality-in-nonsmooth-stochastic-approximation/
10.1214/24-AOS2401.full.

Daudel:2021:IDG

[DDP21] Kamélia Daudel, Randal Douc, and François Portier. Infinitedimensional gradient-based descent for alpha-divergence minimisation. Annals of Statistics, 49(4):2250–2270, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-4/Infinite-dimensional-
gradient-based-descent-for-alpha-divergence-minimisation/
10.1214/20-AOS2035.full.

Dou:2024:REH

[DFZ24] Zehao Dou, Zhou Fan, and Harrison H. Zhou. Rates of estimation for high-dimensional multireference alignment. Annals of Statistics, 52(1):261–284, February 2024. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-1/Rates-of-estimation-
for-high-dimensional-multireference-alignment/10.1214/
23-AOS2346.full.

DeCastro:2021:SSR

[DGMMR21] Y. De Castro, S. Gadat, C. Marteau, and C. Maugis-Rabusseau. SuperMix: Sparse regularization for mixtures. Annals of Statistics, 49(3):1779–1809, June 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-3/SuperMix-Sparse-regularization-
for-mixtures/10.1214/20-AOS2022.full.

Donoho:2023:SEM

[DGR23] David Donoho, Matan Gavish, and Elad Romanov. ScreeNOT: Exact MSE-optimal singular value thresholding in correlated noise. Annals of Statistics, 51(1):122–148, February 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-1/ScreeNOT--Exact-MSE-
optimal-singular-value-thresholding-in-correlated/10.
1214/22-AOS2232.full.

Dudeja:2024:SCT

[DH24] Rishabh Dudeja and Daniel Hsu. Statistical-computational trade-offs in tensor PCA and related problems via communication complexity. Annals of Statistics, 52(1):131– 156, February 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-1/
Statistical-computational-trade-offs-in-tensor-PCA-and-
related-problems/10.1214/23-AOS2331.full.

Dwivedi:2020:SMC

[DHK^{+}20] Raaz Dwivedi, Nhat Ho, Koulik Khamaru, Martin J. Wainwright, Michael I. Jordan, and Bin Yu. Singularity, misspecification and the convergence rate of EM. Annals of Statistics, 48(6):

3161–3182, December 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-6/
Singularity-misspecification-and-the-convergence-rate-
of-EM/10.1214/19-AOS1924.full.

Drton:2020:HDC

[DHS20] Mathias Drton, Fang Han, and Hongjian Shi. High-dimensional consistent independence testing with maxima of rank correlations. Annals of Statistics, 48(6):3206–3227, December 2020. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-48/issue-6/High-
dimensional-consistent-independence-testing-with-maxima-
of-rank-correlations/10.1214/19-AOS1926.full.

Deng:2021:CIM

[DHZ21] Hang Deng, Qiyang Han, and Cun-Hui Zhang. Confidence intervals for multiple isotonic regression and other monotone models. Annals of Statistics, 49(4):2021–2052, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-4/Confidence-intervals-
for-multiple-isotonic-regression-and-other-monotone-
models/10.1214/20-AOS2025.full.

Duffy:2021:EIP

[DK21] James A. Duffy and Ioannis Kasparis. Estimation and inference in the presence of fractional. Annals of Statistics, 49(2): 1195–1217, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-2/
Estimation-and-inference-in-the-presence-of-fractional-
d1-2/10.1214/20-AOS1998.full.

Donoho:2022:HCC

[DK22] David L. Donoho and Alon Kipnis. Higher criticism to compare two large frequency tables, with sensitivity to possible rare and weak differences. Annals of Statistics, 50(3): 1447–1472, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-3/
Higher-criticism-to-compare-two-large-frequency-tables-
with-sensitivity/10.1214/21-AOS2158.full.

Dette:2020:FDA

[DKA20] Holger Dette, Kevin Kokot, and Alexander Aue. Functional data analysis in the Banach space of continuous functions. Annals of Statistics, 48(2):1168–1192, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1590480050.

Drton:2021:EUK

[DKH21] Mathias Drton, Satoshi Kuriki, and Peter Hoff. Existence and uniqueness of the Kronecker covariance MLE. Annals of Statistics, 49(5):2721–2754, October 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-5/Existence-and-uniqueness-
of-the-Kronecker-covariance-MLE/10.1214/21-AOS2052.full.

Depersin:2022:RSG

[DL22] Jules Depersin and Guillaume Lecué. Robust sub-Gaussian estimation of a mean vector in nearly linear time. Annals of Statistics, 50(1):511–536, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
50/issue-1/Robust-sub-Gaussian-estimation-of-a-mean-
vector-in-nearly/10.1214/21-AOS2118.full.

Dobriban:2020:OPL

[DLS20] Edgar Dobriban, William Leeb, and Amit Singer. Optimal prediction in the linearly transformed spiked model. Annals of Statistics, 48(1):491–513, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https:/

/projecteuclid.org/euclid.aos/1581930144.

Dette:2022:DAG

[DLY22] Holger Dette, Xin Liu, and Rong-Xian Yue. Design admissibility and de la Garza phenomenon in multifactor experiments. Annals of Statistics, 50(3):1247–1265, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-3/Design-admissibility-
and-de-la-Garza-phenomenon-in-multifactor-experiments/
10.1214/21-AOS2147.full.

Dubey:2020:FCP

[DM20] Paromita Dubey and Hans-Georg Müller. Fréchet change-point detection. Annals of Statistics, 48(6):3312–3335, December 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-6/Fr%c3%a9chet-change-
point-detection/10.1214/19-AOS1930.full.

Dalalyan:2022:AOR

[DM22] Arnak S. Dalalyan and Arshak Minasyan. All-in-one robust estimator of the Gaussian mean. Annals of Statistics, 50(2):1193– 1219, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-50/issue-2/All-
in-one-robust-estimator-of-the-Gaussian-mean/10.1214/
21-AOS2145.full.

Ding:2023:LLD

[DM23] Xiucai Ding and Rong Ma. Learning low-dimensional nonlinear structures from high-dimensional noisy data: an integral operator approach. Annals of Statistics, 51(4):1744–1769, August 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-51/issue-4/Learning-low-
dimensional-nonlinear-structures-from-high-dimensional-
noisy-data/10.1214/23-AOS2306.full.

Dirksen:2022:CEU

[DMR22] Sjoerd Dirksen, Johannes Maly, and Holger Rauhut. Covariance estimation under one-bit quantization. Annals of Statistics, 50(6):3538–3562, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
50/issue-6/Covariance-estimation-under-one-bit-quantization/
10.1214/22-AOS2239.full.

Durmus:2020:IGE

[DMS20] Alain Durmus, Éric Moulines, and Eero Saksman. Irreducibility and geometric ergodicity of Hamiltonian Monte Carlo. Annals of Statistics, 48(6):3545–3564, December 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-

of-statistics/volume-48/issue-6/Irreducibility-and-geometric-
ergodicity-of-Hamiltonian-Monte-Carlo/10.1214/19-AOS1941.
full.

Duchi:2021:LMU

[DN21] John C. Duchi and Hongseok Namkoong. Learning models with uniform performance via distributionally robust optimization. Annals of Statistics, 49(3):1378–1406, June 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-3/Learning-models-with-
uniform-performance-via-distributionally-robust-optimization/
10.1214/20-AOS2004.full.

Dobriban:2020:PMF

[Dob20] Edgar Dobriban. Permutation methods for factor analysis and PCA. Annals of Statistics, 48(5):2824–2847, October 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-5/Permutation-methods-
for-factor-analysis-and-PCA/10.1214/19-AOS1907.full.

Dobriban:2022:CIB

[Dob22] Edgar Dobriban. Consistency of invariance-based randomization tests. Annals of Statistics, 50(4):2443–2466, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-4/Consistency-of-invariance-
based-randomization-tests/10.1214/22-AOS2200.full.

Delyon:2021:SAI

[DP21] Bernard Delyon and François Portier. Safe adaptive importance sampling: a mixture approach. Annals of Statistics, 49 (2):885–917, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-2/
Safe-adaptive-importance-sampling-A-mixture-approach/
10.1214/20-AOS1983.full.

Durre:2022:AEI

[DP22] Alexander Dürre and Davy Paindaveine. Affine-equivariant inference for multivariate location under. Annals of Statistics, 50(5): 2616–2640, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-5/
Affine-equivariant-inference-for-multivariate-location-
under-Lp-loss-functions/10.1214/22-AOS2199.full.

Ding:2020:DSA

[DQW20] Shanshan Ding, Wei Qian, and Lan Wang. Double-slicing assisted sufficient dimension reduction for high-dimensional censored data. Annals of Statistics, 48(4):2132–2154, August 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-4/Double-slicing-assisted-
sufficient-dimension-reduction-for-high-dimensional-
censored/10.1214/19-AOS1880.full.

Duanmu:2021:EAP

[DR21a] Haosui Duanmu and Daniel M. Roy. On extended admissible procedures and their nonstandard Bayes risk. Annals of Statistics, 49 (4):2053–2078, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-4/
On-extended-admissible-procedures-and-their-nonstandard-
Bayes-risk/10.1214/20-AOS2026.full.

Duchi:2021:AOS

[DR21b] John C. Duchi and Feng Ruan. Asymptotic optimality in stochastic optimization. Annals of Statistics, 49(1):21– 48, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-1/
Asymptotic-optimality-in-stochastic-optimization/10.
1214/19-AOS1831.full.

Duchi:2024:RCM

[DR24] John C. Duchi and Feng Ruan. The right complexity measure in locally private estimation: It is not the Fisher information. Annals of Statistics, 52(1):1–51, February 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-1/The-right-complexity-
measure-in-locally-private-estimation--It/10.1214/22-
AOS2227.full.

deChaumaray:2021:WTS

[dRdCMP21] Marie du Roy de Chaumaray, Matthieu Marbac, and Valentin Patilea. Wilks’ theorem for semiparametric regressions with weakly dependent data. Annals of Statistics, 49(6):3228– 3254, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-6/
Wilks-theorem-for-semiparametric-regressions-with-weakly-
dependent-data/10.1214/21-AOS2081.full.

Drees:2023:SIC

[Dre23] Holger Drees. Statistical inference on a changing extreme value dependence structure. Annals of Statistics, 51(4):1824– 1849, August 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-4/
Statistical-inference-on-a-changing-extreme-value-dependence-
structure/10.1214/23-AOS2314.full.

Donnet:2020:NBE

[DRR20] Sophie Donnet, Vincent Rivoirard, and Judith Rousseau. Nonparametric Bayesian estimation for multivariate Hawkes processes. Annals of Statistics, 48(5):2698–2727, October 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-48/issue-5/Nonparametric-
Bayesian-estimation-for-multivariate-Hawkes-processes/
10.1214/19-AOS1903.full.

Dobriban:2021:DLR

[DS21] Edgar Dobriban and Yue Sheng. Distributed linear regression by averaging. Annals of Statistics, 49(2):918– 943, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-2/
Distributed-linear-regression-by-averaging/10.1214/20-
AOS1984.full.

Derumigny:2023:LBB

[DSH23] Alexis Derumigny and Johannes Schmidt-Hieber. On lower bounds for the bias-variance trade-off. Annals of Statistics, 51 (4):1510–1533, August 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-4/
On-lower-bounds-for-the-bias-variance-trade-off/10.1214/
23-AOS2279.full.

Daouia:2023:IER

[DSUC23] Abdelaati Daouia, Gilles Stupfler, and Antoine Usseglio-Carleve. Inference for extremal regression with dependent heavytailed data. Annals of Statistics, 51(5):2040–2066, October 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-51/issue-5/Inference-for-
extremal-regression-with-dependent-heavy-tailed-data/
10.1214/23-AOS2320.full.

Duan:2023:ARM

[DW23a] Yaqi Duan and Kaizheng Wang. Adaptive and robust multi-task learning. Annals of Statistics, 51(5):2015–2039, October 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-5/Adaptive-and-robust-
multi-task-learning/10.1214/23-AOS2319.full.

Dumbgen:2023:NAT

[DW23b] Lutz Dümbgen and Jon A. Wellner. A new approach to tests and confidence bands for distribution functions. Annals of Statistics, 51(1):260–289, February 2023. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-1/A-new-approach-to-tests-
and-confidence-bands-for-distribution/10.1214/22-AOS2249.
full.

Doss:2024:NDR

[DWW^{+}24] Charles R. Doss, Guangwei Weng, Lan Wang, Ira Moscovice, and Tongtan Chantarat. A nonparametric doubly robust test for a continuous treatment effect. Annals of Statistics, 52(4): 1592–1615, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-4/
A-nonparametric-doubly-robust-test-for-a-continuous-
treatment-effect/10.1214/24-AOS2405.full.

Doss:2023:OEH

[DWYZ23] Natalie Doss, Yihong Wu, Pengkun Yang, and Harrison H. Zhou. Optimal estimation of high-dimensional Gaussian location mixtures. Annals of Statistics, 51(1):62–95, February 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-1/Optimal-estimation-of-
high-dimensional-Gaussian-location-mixtures/10.1214/
22-AOS2207.full.

Ding:2021:SSC

[DY21] Xiucai Ding and Fan Yang. Spiked separable covariance matrices and principal components. Annals of Statistics, 49(2): 1113–1138, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-2/
Spiked-separable-covariance-matrices-and-principal-components/
10.1214/20-AOS1995.full.

Deng:2020:BGA

[DZ20a] Hang Deng and Cun-Hui Zhang. Beyond Gaussian approximation: Bootstrap for maxima of sums of independent random vectors. Annals of Statistics, 48(6):3643– 3671, December 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-6/
Beyond-Gaussian-approximation--Bootstrap-for-maxima-
of-sums-of/10.1214/20-AOS1946.full.

Deng:2020:IRM

[DZ20b] Hang Deng and Cun-Hui Zhang. Isotonic regression in multidimensional spaces and graphs. Annals of Statistics, 48(6): 3672–3698, December 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-6/
Isotonic-regression-in-multi-dimensional-spaces-and-
graphs/10.1214/20-AOS1947.full.

Ding:2020:EIP

[DZ20c] Xiucai Ding and Zhou Zhou. Estimation and inference for precision matrices of nonstationary time series. Annals of Statistics, 48(4):2455–2477, August 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-4/Estimation-and-inference-
for-precision-matrices-of-nonstationary-time-series/
10.1214/19-AOS1894.full.

Ding:2023:AAN

[DZ23] Xiucai Ding and Zhou Zhou. AutoRegressive approximations to nonstationary time series with inference and applications. Annals of Statistics, 51(3):1207–1231, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-3/AutoRegressive-approximations-
to-nonstationary-time-series-with-inference-and-applications/
10.1214/23-AOS2288.full.

Ding:2024:HDC

[DZ24] Yi Ding and Xinghua Zheng. High-dimensional covariance matrices under dynamic volatility models: Asymptotics and shrinkage estimation. Annals of Statistics, 52(3):1027–1049, June 2024. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-52/issue-3/High-
dimensional-covariance-matrices-under-dynamic-volatility-
models--Asymptotics/10.1214/24-AOS2381.full.

Ekvall:2022:CRN

[EB22] Karl Oskar Ekvall and Matteo Bottai. Confidence regions near singular information and boundary points with applications to mixed models. Annals of Statistics, 50(3):1806– 1832, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-3/
Confidence-regions-near-singular-information-and-boundary-
points-with-applications/10.1214/22-AOS2177.full.

Einmahl:2022:SDS

[EFdH^{+}22] John H. J. Einmahl, Ana Ferreira, Laurens de Haan, Cláudia Neves, and Chen Zhou. Spatial dependence and spacetime trend in extreme events. Annals of Statistics, 50(1): 30–52, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-1/
Spatial-dependence-and-spacetime-trend-in-extreme-events/
10.1214/21-AOS2067.full.

Efromovich:2021:SMD

[Efr21] Sam Efromovich. Sharp minimax distribution estimation for current status censoring with or without missing. Annals of Statistics, 49(1):568–589, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
49/issue-1/Sharp-minimax-distribution-estimation-for-
current-status-censoring-with-or/10.1214/20-AOS1970.
full.

Efromovich:2022:NBD

[Efr22] Sam Efromovich. Nonparametric bivariate density estimation for censored lifetimes. Annals of Statistics, 50(5):2767–2792, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-50/issue-5/Nonparametric-
bivariate-density-estimation-for-censored-lifetimes/
10.1214/22-AOS2209.full.

Eaton:2021:CSI

[EG21] Morris L. Eaton and Edward I. George. Charles Stein and invariance: Beginning with the Hunt–Stein theorem. Annals of Statistics, 49(4):1815–1822, August 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-4/Charles-Stein-and-invariance-
Beginning-with-the-HuntStein-theorem/10.1214/21-AOS2075.
full.

Einmahl:2023:EVI

[EH23] John H. J. Einmahl and Yi He. Extreme value inference for heterogeneous power law data. Annals of Statistics, 51(3): 1331–1356, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-3/
Extreme-value-inference-for-heterogeneous-power-law-
data/10.1214/23-AOS2294.full.

Ekvall:2020:CML

[EJ20] Karl Oskar Ekvall and Galin L. Jones. Consistent maximum likelihood estimation using subsets with applications to multivariate mixed models. Annals of Statistics, 48(2):932–952, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/euclid.aos/
1590480040.

ElAlaoui:2020:FLD

[EKJ20] Ahmed El Alaoui, Florent Krzakala, and Michael Jordan. Fundamental limits of detection in the spiked Wigner model. Annals of Statistics, 48(2):863–885, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1590480037.

Edelmann:2020:DSD

[ERV20] Dominic Edelmann, Donald Richards, and Daniel Vogel. The distance standard deviation. Annals of Statistics, 48(6):3395–3416, December 2020. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-48/issue-6/The-
distance-standard-deviation/10.1214/19-AOS1935.full.

Einmahl:2021:ETC

[ES21] John H. J. Einmahl and Johan Segers. Empirical tail copulas for functional data. Annals of Statistics, 49(5):2672– 2696, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-5/
Empirical-tail-copulas-for-functional-data/10.1214/21-
AOS2050.full.

Evans:2020:MSL

[Eva20] Robin J. Evans. Model selection and local geometry. Annals of Statistics, 48(6):3513–3544, December 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-6/Model-selection-and-
local-geometry/10.1214/19-AOS1940.full.

Fan:2022:AMP

[Fan22] Zhou Fan. Approximate Message Passing algorithms for rotationally invariant matrices. Annals of Statistics, 50(1):197– 224, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-1/
Approximate-Message-Passing-algorithms-for-rotationally-
invariant-matrices/10.1214/21-AOS2101.full.

Feng:2021:AML

[FGKS21] Oliver Y. Feng, Adityanand Guntuboyina, Arlene K. H. Kim, and Richard J. Samworth. Adaptation in multivariate log-concave density estimation. Annals of Statistics, 49(1):129–153, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-49/issue-1/Adaptation-in-
multivariate-log-concave-density-estimation/10.1214/
20-AOS1950.full.

Fathi:2022:RGA

[FGRS22] Max Fathi, Larry Goldstein, Gesine Reinert, and Adrien Saumard. Relaxing the Gaussian assumption in shrinkage and SURE in high dimension. Annals of Statistics, 50(5): 2737–2766, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-5/
Relaxing-the-Gaussian-assumption-in-shrinkage-and-SURE-
in-high/10.1214/22-AOS2208.full.

Fang:2021:MEI

[FGS21a] Billy Fang, Adityanand Guntuboyina, and Bodhisattva Sen. Multivariate extensions of isotonic regression and total variation denoising via entire monotonicity and Hardy–Krause variation. Annals of Statistics, 49(2):769–792, April 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-2/Multivariate-extensions-
of-isotonic-regression-and-total-variation-denoising-
via/10.1214/20-AOS1977.full.

Freise:2021:AWA

[FGS21b] Fritjof Freise, Norbert Gaffke, and Rainer Schwabe. The adaptive Wynn algorithm in generalized linear models with univariate response. Annals of Statistics, 49(2):702–722, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-49/issue-2/The-
adaptive-Wynn-algorithm-in-generalized-linear-models-
with-univariate/10.1214/20-AOS1974.full.

Fan:2024:HDN

[FGZ24] Jianqing Fan, Yihong Gu, and Wen-Xin Zhou. How do noise tails impact on deep ReLU networks? Annals of Statistics, 52 (4):1845–1871, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-4/
How-do-noise-tails-impact-on-deep-ReLU-networks/10.1214/
24-AOS2428.full.

Feng:2022:MST

[FJLX22] Long Feng, Tiefeng Jiang, Binghui Liu, and Wei Xiong. Maxsum tests for cross-sectional independence of high-dimensional panel data. Annals of Statistics, 50(2):1124–1143, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-50/issue-2/Max-sum-tests-
for-cross-sectional-independence-of-high-dimensional/
10.1214/21-AOS2142.full.

Fithian:2022:CCF

[FL22] William Fithian and Lihua Lei. Conditional calibration for false discovery rate control under dependence. Annals of Statistics, 50(6):3091–3118, December 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-6/Conditional-calibration-
for-false-discovery-rate-control-under-dependence/10.
1214/21-AOS2137.full.

Fang:2020:SEC

[FLS20] Xiao Fang, Jian Li, and David Siegmund. Segmentation and estimation of change-point models: False positive control and confidence regions. Annals of Statistics, 48(3): 1615–1647, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-3/
Segmentation-and-estimation-of-change-point-models--
False-positive/10.1214/19-AOS1861.full.

Fan:2024:MLH

[FLS^{+}24] Zhou Fan, Roy R. Lederman, Yi Sun, Tianhao Wang, and Sheng Xu. Maximum likelihood for high-noise group orbit estimation and single-particle cryo-EM. Annals of Statistics, 52 (1):52–77, February 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-1/
Maximum-likelihood-for-high-noise-group-orbit-estimation-
and-single/10.1214/23-AOS2292.full.

Fan:2023:RHD

[FLY23] Jianqing Fan, Zhipeng Lou, and Mengxin Yu. Robust highdimensional tuning free multiple testing. Annals of Statistics, 51 (5):2093–2115, October 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-5/
Robust-high-dimensional-tuning-free-multiple-testing/
10.1214/23-AOS2322.full.

Fan:2023:BFS

[FMM23] Jianqing Fan, Ricardo P. Masini, and Marcelo C. Medeiros. Bridging factor and sparse models. Annals of Statistics, 51(4): 1692–1717, August 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-4/
Bridging-factor-and-sparse-models/10.1214/23-AOS2304.
full.

Fang:2020:TSH

[FNL20]  Ethan X. Fang, Yang Ning, and Runze Li.  Test of

significance for high-dimensional longitudinal data. Annals of Statistics, 48(5):2622–2645, October 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-5/Test-of-significance-
for-high-dimensional-longitudinal-data/10.1214/19-AOS1900.
full.

Feng:2022:NME

[FNZ22] Huijie Feng, Yang Ning, and Jiwei Zhao. Nonregular and minimax estimation of individualized thresholds in high dimension with binary responses. Annals of Statistics, 50(4): 2284–2305, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-4/
Nonregular-and-minimax-estimation-of-individualized-
thresholds-in-high-dimension/10.1214/22-AOS2188.full.

Fischer:2024:OCP

[FRB24] Lasse Fischer, Marta Bofill Roig, and Werner Brannath. The online closure principle. Annals of Statistics, 52(2):817–841, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-52/issue-2/The-
online-closure-principle/10.1214/24-AOS2370.full.

Fisher:2021:WVZ

[FS21] Nicholas I. Fisher and Adrian F. M. Smith. Willem van Zwet’s contributions to the profession. Annals of Statistics, 49(5): 2432–2438, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-5/
Willem-van-Zwets-contributions-to-the-profession/10.
1214/21-AOS2053.full.

Foster:2023:OSL

[FS23] Dylan J. Foster and Vasilis Syrgkanis. Orthogonal statistical learning. Annals of Statistics, 51(3):879–908, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-3/Orthogonal-statistical-
learning/10.1214/23-AOS2258.full.

Fan:2021:PCL

[FSW21] Zhou Fan, Yi Sun, and Zhichao Wang. Principal components in linear mixed models with general bulk. Annals of Statistics, 49 (3):1489–1513, June 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-3/
Principal-components-in-linear-mixed-models-with-general-
bulk/10.1214/20-AOS2010.full.

Finke:2023:CSM

[FT23] Axel Finke and Alexandre H. Thiery. Conditional sequential Monte Carlo in high dimensions. Annals of Statistics, 51 (2):437–463, April 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-2/
Conditional-sequential-Monte-Carlo-in-high-dimensions/
10.1214/22-AOS2252.full.

Fan:2021:SPH

[FWZ21] Jianqing Fan, Weichen Wang, and Ziwei Zhu. A shrinkage principle for heavy-tailed data: High-dimensional robust low-rank matrix recovery. Annals of Statistics, 49(3): 1239–1266, June 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-3/
A-shrinkage-principle-for-heavy-tailed-data--High-dimensional/
10.1214/20-AOS1980.full.

Fujiwara:2023:EEL

[FY23] Akio Fujiwara and Koichi Yamagata. Efficiency of estimators for locally asymptotically normal quantum statistical models. Annals of Statistics, 51(3):1159–1182, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-51/issue-3/Efficiency-of-
estimators-for-locally-asymptotically-normal-quantum-
statistical-models/10.1214/23-AOS2285.full.

Fissler:2016:HOE

[FZ16] Tobias Fissler and Johanna F. Ziegel. Higher order elicitability and Osband’s principle. Annals of Statistics, 44(4):1680–1707, August 2016. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1467894712. See correction [FZ21].

Fissler:2021:CNH

[FZ21] Tobias Fissler and Johanna F. Ziegel. Correction note: Higher order elicitability and Osband’s principle. Annals of Statistics, 49(1):614, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-1/
Correction-note-Higher-order-elicitability-and-Osbands-
principle/10.1214/20-AOS2014.full. See [FZ16].

Fauss:2020:MOS

[FZP20] Michael Fauß, Abdelhak M. Zoubir, and H. Vincent Poor. Minimax optimal sequential hypothesis tests for Markov processes. Annals of Statistics, 48(5):2599–2621, October 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-

of-statistics/volume-48/issue-5/Minimax-optimal-sequential-
hypothesis-tests-for-Markov-processes/10.1214/19-AOS1899.
full.

Guo:2022:DDL

[GĆB22] Zijian Guo, Domagoj Ćevid, and Peter Bühlmann. Doubly debiased lasso: High-dimensional inference under hidden confounding. Annals of Statistics, 50(3):1320–1347, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-3/Doubly-debiased-lasso-
High-dimensional-inference-under-hidden-confounding/
10.1214/21-AOS2152.full.

Gerhardus:2024:CCA

[Ger24] Andreas Gerhardus. Characterization of causal ancestral graphs for time series with latent confounders. Annals of Statistics, 52 (1):103–130, February 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-1/
Characterization-of-causal-ancestral-graphs-for-time-
series-with-latent/10.1214/23-AOS2325.full.

Gao:2021:ADH

[GFLS21] Lan Gao, Yingying Fan, Jinchi Lv, and Qi-Man Shao. Asymptotic distributions of high-dimensional distance correlation inference. Annals of Statistics, 49(4):1999–2020, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-4/Asymptotic-distributions-
of-high-dimensional-distance-correlation-inference/10.
1214/20-AOS2024.full.

Ghoshdastidar:2020:TSH

[GGCvL20] Debarghya Ghoshdastidar, Maurilio Gutzeit, Alexandra Carpentier, and Ulrike von Luxburg. Two-sample hypothesis testing for inhomogeneous random graphs. Annals of Statistics, 48(4):2208– 2229, August 2020. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-48/issue-4/Two-
sample-hypothesis-testing-for-inhomogeneous-random-graphs/
10.1214/19-AOS1884.full.

Gerber:2021:OIM

[GH21] Mathieu Gerber and Kari Heine. Online inference with multimodal likelihood functions. Annals of Statistics, 49(6):3103– 3126, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-6/
Online-inference-with-multi-modal-likelihood-functions/
10.1214/21-AOS2076.full.

Ghosh:2022:BLS

[GHO22] Swarnadip Ghosh, Trevor Hastie, and Art B. Owen. Backfitting for large scale crossed random effects regressions. Annals of Statistics, 50(1):560–583, February 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-1/Backfitting-for-large-
scale-crossed-random-effects-regressions/10.1214/21-
AOS2121.full.

Goeman:2021:OCT

[GHS21] Jelle J. Goeman, Jesse Hemerik, and Aldo Solari. Only closed testing procedures are admissible for controlling false discovery proportions. Annals of Statistics, 49(2):1218–1238, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-2/Only-closed-testing-
procedures-are-admissible-for-controlling-false-discovery/
10.1214/20-AOS1999.full.

Gao:2020:EIP

[GHZ20] Chao Gao, Fang Han, and Cun-Hui Zhang. On estimation of isotonic piecewise constant signals. Annals of Statistics, 48 (2):629–654, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/euclid.aos/1590480028.

Graczyk:2022:MSS

[GIKM22] Piotr Graczyk, Hideyuki Ishi, Bartosz Kołodziejek, and Hélène Massam. Model selection in the space of Gaussian models invariant by symmetry. Annals of Statistics, 50(3): 1747–1774, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-3/

Model-selection-in-the-space-of-Gaussian-models-invariant-
by/10.1214/22-AOS2174.full.

Ghosh:2021:SSC

[GKM21] Satyajit Ghosh, Kshitij Khare, and George Michailidis. Strong selection consistency of Bayesian vector autoregressive models based on a pseudo-likelihood approach. Annals of Statistics, 49 (3):1267–1299, June 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-3/
Strong-selection-consistency-of-Bayesian-vector-autoregressive-
models-based-on/10.1214/20-AOS1992.full.

Goto:2022:ICS

[GKV^{+}22] Yuichi Goto, Tobias Kley, Ria Van Hecke, Stanislav Volgushev, Holger Dette, and Marc Hallin. The integrated copula spectrum. Annals of Statistics, 50(6):3563–3591, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-50/issue-6/The-
integrated-copula-spectrum/10.1214/22-AOS2240.full.

Guntuboyina:2020:ARB

[GLCS20] Adityanand Guntuboyina, Donovan Lieu, Sabyasachi Chatterjee, and Bodhisattva Sen. Adaptive risk bounds in univariate total variation denoising and trend filtering. Annals of Statistics, 48 (1):205–229, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/euclid.aos/1581930132.

Gassiat:2022:DUN

[GLL22] Élisabeth Gassiat, Sylvain Le Corff, and Luc Lehéricy. Deconvolution with unknown noise distribution is possible for multivariate signals. Annals of Statistics, 50(1):303–323, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-50/issue-1/Deconvolution-
with-unknown-noise-distribution-is-possible-for-multivariate-
signals/10.1214/21-AOS2106.full.

Goes:2020:RSC

[GLN20] John Goes, Gilad Lerman, and Boaz Nadler. Robust sparse covariance estimation by thresholding Tyler’s M-estimator. Annals of Statistics, 48(1):86–110, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1581930127.

Giordano:2020:GVS

[GLP20] Francesco Giordano, Soumendra Nath Lahiri, and Maria Lucia Parrella. GRID: A variable selection and structure discovery method for high dimensional nonparametric regression. Annals of Statistics, 48(3):1848–1874, June 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-3/GRID--A-variable-selection-
and-structure-discovery-method-for/10.1214/19-AOS1846.
full.

Ghosal:2020:JEP

[GM20] Promit Ghosal and Sumit Mukherjee. Joint estimation of parameters in Ising model. Annals of Statistics, 48(2):785–810, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/euclid.aos/
1590480034.

Gao:2022:TEC

[GM22] Chao Gao and Zongming Ma. Testing equivalence of clustering. Annals of Statistics, 50(1):407–429, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-1/Testing-equivalence-
of-clustering/10.1214/21-AOS2113.full.

Ghorbani:2021:LTL

[GMMM21] Behrooz Ghorbani, Song Mei, Theodor Misiakiewicz, and Andrea Montanari. Linearized two-layers neural networks in high dimension. Annals of Statistics, 49(2):1029–1054, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-2/Linearized-two-layers-
neural-networks-in-high-dimension/10.1214/20-AOS1990.
full.

Gnecco:2021:CDH

[GMPE21] Nicola Gnecco, Nicolai Meinshausen, Jonas Peters, and Sebastian Engelke. Causal discovery in heavy-tailed models. Annals of Statistics, 49(3):1755–1778, June 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-3/Causal-discovery-in-
heavy-tailed-models/10.1214/20-AOS2021.full.

Gregory:2021:SIS

[GMW21] Karl Gregory, Enno Mammen, and Martin Wahl. Statistical inference in sparse high-dimensional additive models. Annals of Statistics, 49(3):1514–1536, June 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-3/Statistical-inference-
in-sparse-high-dimensional-additive-models/10.1214/20-
AOS2011.full.

Goodman:2022:AAS

[Goo22] Jesse Goodman. Asymptotic accuracy of the saddlepoint approximation for maximum likelihood estimation. Annals of Statistics, 50(4):2021–2046, August 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-4/Asymptotic-accuracy-
of-the-saddlepoint-approximation-for-maximum-likelihood-
estimation/10.1214/22-AOS2169.full.

Giordano:2022:NBI

[GR22] Matteo Giordano and Kolyan Ray. Nonparametric Bayesian inference for reversible multidimensional diffusions. Annals of Statistics, 50(5):2872–2898, October 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-5/Nonparametric-Bayesian-
inference-for-reversible-multidimensional-diffusions/
10.1214/22-AOS2213.full.

Ghosal:2022:MRQ

[GS22] Promit Ghosal and Bodhisattva Sen. Multivariate ranks and quantiles using optimal transport: Consistency, rates and nonparametric testing. Annals of Statistics, 50(2):1012–1037, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-50/issue-2/Multivariate-
ranks-and-quantiles-using-optimal-transport--Consistency-
rates/10.1214/21-AOS2136.full.

Gao:2022:RCT

[GSS22] Lan Gao, Qi-Man Shao, and Jiasheng Shi. Refined Cramértype moderate deviation theorems for general self-normalized sums with applications to dependent random variables and winsorized mean. Annals of Statistics, 50(2):673–697, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-50/issue-2/Refined-Cram%
c3%a9r-type-moderate-deviation-theorems-for-general-
self-normalized/10.1214/21-AOS2122.full.

Girard:2021:ECE

[GSUC21] Stéphane Girard, Gilles Stupfler, and Antoine Usseglio-Carleve. Extreme conditional expectile estimation in heavy-tailed heteroscedastic regression models. Annals of Statistics, 49(6):3358– 3382, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-6/
Extreme-conditional-expectile-estimation-in-heavy-tailed-
heteroscedastic-regression-models/10.1214/21-AOS2087.
full.

Gao:2020:GFB

[GvdVZ20] Chao Gao, Aad W. van der Vaart, and Harrison H. Zhou. A general framework for Bayes structured linear models. Annals of Statistics, 48(5):2848–2878, October 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-5/A-general-framework-
for-Bayes-structured-linear-models/10.1214/19-AOS1909.
full.

Gao:2022:TST

[GW22] Fengnan Gao and Tengyao Wang. Two-sample testing of high-dimensional linear regression coefficients via complementary sketching. Annals of Statistics, 50(5):2950–2972, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-50/issue-5/Two-
sample-testing-of-high-dimensional-linear-regression-
coefficients-via/10.1214/22-AOS2216.full.

Gu:2020:PIR

[GX20] Yuqi Gu and Gongjun Xu. Partial identifiability of restricted latent class models. Annals of Statistics, 48(4):2082– 2107, August 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-4/
Partial-identifiability-of-restricted-latent-class-models/
10.1214/19-AOS1878.full.

Gamarnik:2022:SHD

[GZ22a] David Gamarnik and Ilias Zadik. Sparse high-dimensional linear regression. estimating squared error and a phase transition. Annals of Statistics, 50(2):880–903, April 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-2/Sparse-high-dimensional-
linear-regression-Estimating-squared-error-and-a/10.
1214/21-AOS2130.full.

Gao:2022:IAD

[GZ22b] Chao Gao and Anderson Y. Zhang. Iterative algorithm for discrete structure recovery. Annals of Statistics, 50(2): 1066–1094, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-2/
Iterative-algorithm-for-discrete-structure-recovery/
10.1214/21-AOS2140.full.

Han:2021:SSG

[Han21] Qiyang Han. Set structured global empirical risk minimizers are rate optimal in general dimensions. Annals of Statistics, 49(5): 2642–2671, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-5/
Set-structured-global-empirical-risk-minimizers-are-
rate-optimal-in/10.1214/21-AOS2049.full.

Han:2023:NLI

[Han23] Qiyang Han. Noisy linear inverse problems under convex constraints: Exact risk asymptotics in high dimensions. Annals of Statistics, 51(4):1611–1638, August 2023. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-

of-statistics/volume-51/issue-4/Noisy-linear-inverse-
problems-under-convex-constraints--Exact-risk/10.1214/
23-AOS2301.full.

Heng:2020:CSM

[HBDD20] Jeremy Heng, Adrian N. Bishop, George Deligiannidis, and Arnaud Doucet. Controlled sequential Monte Carlo. Annals of Statistics, 48(5):2904–2929, October 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-5/Controlled-sequential-
Monte-Carlo/10.1214/19-AOS1914.full.

Huang:2022:COS

[HCCI22] Hsueh-Han Huang, Ngai Hang Chan, Kun Chen, and Ching-Kang Ing. Consistent order selection for ARFIMA processes. Annals of Statistics, 50(3):1297–1319, June 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-3/Consistent-order-selection-
for-ARFIMA-processes/10.1214/21-AOS2149.full.

Hallin:2021:DQF

[HdBCAM21] Marc Hallin, Eustasio del Barrio, Juan Cuesta-Albertos, and Carlos Matrán. Distribution and quantile functions, ranks and signs in dimension d: a measure transportation approach. Annals of Statistics, 49(2):1139–1165, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-49/issue-2/Distribution-
and-quantile-functions-ranks-and-signs-in-dimension-
d/10.1214/20-AOS1996.full.

He:2020:ICV

[HHPS20] Yi He, Yanxi Hou, Liang Peng, and Haipeng Shen. Inference for conditional value-at-risk of a predictive regression. Annals of Statistics, 48(6):3442–3464, December 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-6/Inference-for-conditional-
value-at-risk-of-a-predictive-regression/10.1214/19-
AOS1937.full.

Hidalgo:2021:BLM

[Hid21] Javier Hidalgo. Bootstrap long memory processes in the frequency domain. Annals of Statistics, 49(3):1407–1435, June 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-49/issue-3/Bootstrap-long-
memory-processes-in-the-frequency-domain/10.1214/20-
AOS2006.full.

Huang:2020:RAK

[HJ20] Dongming Huang and Lucas Janson. Relaxing the assumptions of knockoffs by conditioning. Annals of Statistics, 48(5): 3021–3042, October 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-5/
Relaxing-the-assumptions-of-knockoffs-by-conditioning/
10.1214/19-AOS1920.full.

Hassani:2024:COA

[HJ24] Hamed Hassani and Adel Javanmard. The curse of overparametrization in adversarial training: Precise analysis of robust generalization for random features regression. Annals of Statistics, 52(2):441–465, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
52/issue-2/The-curse-of-overparametrization-in-adversarial-
training--Precise-analysis/10.1214/24-AOS2353.full.

Han:2020:ORE

[HJWW20] Yanjun Han, Jiantao Jiao, Tsachy Weissman, and Yihong Wu. Optimal rates of entropy estimation over Lipschitz balls. Annals of Statistics, 48(6):3228–3250, December 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-6/Optimal-rates-of-entropy-
estimation-over-Lipschitz-balls/10.1214/19-AOS1927.full.

Hanneke:2022:NFL

[HK22] Steve Hanneke and Samory Kpotufe. A no-free-lunch theorem for multitask learning. Annals of Statistics, 50(6):3119–3143, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-50/issue-6/A-no-

free-lunch-theorem-for-multitask-learning/10.1214/22-
AOS2189.full.

Hanneke:2021:UBC

[HKSW21] Steve Hanneke, Aryeh Kontorovich, Sivan Sabato, and Roi Weiss. Universal Bayes consistency in metric spaces. Annals of Statistics, 49(4):2129–2150, August 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-4/Universal-Bayes-consistency-
in-metric-spaces/10.1214/20-AOS2029.full.

He:2024:OCP

[HKTY24] Yong He, Xinbing Kong, Lorenzo Trapani, and Long Yu. Online change-point detection for matrix-valued time series with latent two-way factor structure. Annals of Statistics, 52(4): 1646–1670, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-4/
Online-change-point-detection-for-matrix-valued-time-
series-with/10.1214/24-AOS2410.full.

Horvath:2021:MCP

[HKW21] Lajos Horváth, Piotr Kokoszka, and Shixuan Wang. Monitoring for a change point in a sequence of distributions. Annals of Statistics, 49(4):2271–2291, August 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-4/Monitoring-for-a-change-
point-in-a-sequence-of-distributions/10.1214/20-AOS2036.
full.

Hong:2020:NB

[HL20] Han Hong and Jessie Li. The numerical bootstrap. Annals of Statistics, 48(1):397–412, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1581930140.

He:2024:HOC

[HL24] Shengyi He and Henry Lam. Higher-order coverage errors of batching methods via Edgeworth expansions on tstatistics. Annals of Statistics, 52(4):1360–1383, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-52/issue-4/Higher-order-
coverage-errors-of-batching-methods-via-Edgeworth-expansions/
10.1214/24-AOS2377.full.

Huang:2023:EEM

[HLM23] Tzu-Jung Huang, Alex Luedtke, and Ian W. McKeague. Efficient estimation of the maximal association between multiple predictors and a survival outcome. Annals of Statistics, 51(5): 1965–1988, October 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-5/
Efficient-estimation-of-the-maximal-association-between-
multiple-predictors-and/10.1214/23-AOS2313.full.

He:2022:NFD

[HLS22] Yuanzhen He, C. Devon Lin, and Fasheng Sun. A new and flexible design construction for orthogonal arrays for modern applications. Annals of Statistics, 50(3):1473–1489, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-3/A-new-and-flexible-design-
construction-for-orthogonal-arrays-for/10.1214/21-AOS2159.
full.

Hazimeh:2023:GVS

[HMR23] Hussein Hazimeh, Rahul Mazumder, and Peter Radchenko. Grouped variable selection with discrete optimization: Computational and statistical perspectives. Annals of Statistics, 51 (1):1–32, February 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-1/
Grouped-variable-selection-with-discrete-optimization-
-Computational-and-statistical/10.1214/21-AOS2155.full.

Hastie:2022:SHD

[HMRT22] Trevor Hastie, Andrea Montanari, Saharon Rosset, and Ryan J. Tibshirani. Surprises in high-dimensional ridgeless least squares interpolation. Annals of Statistics, 50(2):949–986, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-50/issue-2/Surprises-in-
high-dimensional-ridgeless-least-squares-interpolation/
10.1214/21-AOS2133.full.

Hopkins:2020:MES

[Hop20] Samuel B. Hopkins. Mean estimation with sub-Gaussian rates in polynomial time. Annals of Statistics, 48(2):1193–1213, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/euclid.aos/
1590480051.

He:2022:SEI

[HPTZ22] Xuming He, Xiaoou Pan, Kean Ming Tan, and Wen-Xin Zhou. Scalable estimation and inference for censored quantile regression process. Annals of Statistics, 50(5):2899–2924, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-5/Scalable-estimation-
and-inference-for-censored-quantile-regression-process/
10.1214/22-AOS2214.full.

Hutter:2021:MES

[HR21] Jan-Christian Hütter and Philippe Rigollet. Minimax estimation of smooth optimal transport maps. Annals of Statistics, 49 (2):1166–1194, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-2/
Minimax-estimation-of-smooth-optimal-transport-maps/
10.1214/20-AOS1997.full.

Howard:2021:TUN

[HRMS21] Steven R. Howard, Aaditya Ramdas, Jon McAuliffe, and Jasjeet Sekhon. Time-uniform, nonparametric, nonasymptotic confidence sequences. Annals of Statistics, 49(2):1055–1080, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-49/issue-2/Time-uniform-
nonparametric-nonasymptotic-confidence-sequences/10.
1214/20-AOS1991.full.

Hamm:2021:ALR

[HS21a] Thomas Hamm and Ingo Steinwart. Adaptive learning rates for support vector machines working on data with low intrinsic dimension. Annals of Statistics, 49(6):3153– 3180, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-6/

Adaptive-learning-rates-for-support-vector-machines-
working-on-data/10.1214/21-AOS2078.full.

Huang:2021:APP

[HS21b] Jianhua Z. Huang and Ya Su. Asymptotic properties of penalized spline estimators in concave extended linear models: Rates of convergence. Annals of Statistics, 49(6):3383– 3407, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-6/
Asymptotic-properties-of-penalized-spline-estimators-
in-concave-extended-linear/10.1214/21-AOS2088.full.

Huang:2022:LBC

[HS22] Yijian Huang and Martin G. Sanda. Linear biomarker combination for constrained classification. Annals of Statistics, 50(5): 2793–2815, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-5/
Linear-biomarker-combination-for-constrained-classification/
10.1214/22-AOS2210.full.

Han:2023:URR

[HS23] Qiyang Han and Yandi Shen. Universality of regularized regression estimators in high dimensions. Annals of Statistics, 51(4): 1799–1823, August 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-4/
Universality-of-regularized-regression-estimators-in-
high-dimensions/10.1214/23-AOS2309.full.

Hagrass:2024:SRK

[HSL24] Omar Hagrass, Bharath Sriperumbudur, and Bing Li. Spectral regularized kernel two-sample tests. Annals of Statistics, 52 (3):1076–1101, June 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-3/
Spectral-regularized-kernel-two-sample-tests/10.1214/
24-AOS2383.full.

Han:2022:HDA

[HSS22] Qiyang Han, Bodhisattva Sen, and Yandi Shen. Highdimensional asymptotics of likelihood ratio tests in the Gaussian sequence model under convex constraints. Annals of Statistics, 50 (1):376–406, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-1/
High-dimensional-asymptotics-of-likelihood-ratio-tests-
in-the-Gaussian/10.1214/21-AOS2111.full.

Hoffmann:2023:DDE

[HT23] Marc Hoffmann and Mathias Trabs. Dispersal density estimation across scales. Annals of Statistics, 51(3):1258– 1281, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-3/
Dispersal-density-estimation-across-scales/10.1214/23-
AOS2290.full.

Han:2020:LUS

[HTYZ20] Lei Han, Kean Ming Tan, Ting Yang, and Tong Zhang. Local uncertainty sampling for large-scale multiclass logistic regression. Annals of Statistics, 48(3):1770–1788, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-3/Local-uncertainty-sampling-
for-large-scale-multiclass-logistic-regression/10.1214/
19-AOS1867.full.

Huang:2020:ARP

[Hua20] Hanwen Huang. Asymptotic risk and phase transition of ℓ_{1}- penalized robust estimator. Annals of Statistics, 48(5):3090– 3111, October 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-5/
Asymptotic-risk-and-phase-transition-of-l_1-penalized-
robust-estimator/10.1214/19-AOS1923.full.

Han:2020:THD

[HW20] Yuefeng Han and Wei Biao Wu. Test for high dimensional covariance matrices. Annals of Statistics, 48(6):3565–3588, December 2020. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-48/issue-6/Test-
for-high-dimensional-covariance-matrices/10.1214/20-
AOS1943.full.

Han:2021:CSD

[HW21a] Qiyang Han and Jon A. Wellner. Complex sampling designs: Uniform limit theorems and applications. Annals of Statistics, 49 (1):459–485, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-1/
Complex-sampling-designs-Uniform-limit-theorems-and-
applications/10.1214/20-AOS1964.full.

Hirshberg:2021:AML

[HW21b] David A. Hirshberg and Stefan Wager. Augmented minimax linear estimation. Annals of Statistics, 49(6):3206– 3227, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-6/
Augmented-minimax-linear-estimation/10.1214/21-AOS2080.
full.

Hu:2023:PEP

[HW23] Yuchen Hu and Stefan Wager. Off-policy evaluation in partially observed Markov decision processes under sequential ignorability. Annals of Statistics, 51(4):1561–1585, August 2023. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-51/issue-4/Off-
policy-evaluation-in-partially-observed-Markov-decision-
processes-under/10.1214/23-AOS2287.full.

Han:2022:OSC

[HWZ22] Rungang Han, Rebecca Willett, and Anru R. Zhang. An optimal statistical and computational framework for generalized tensor estimation. Annals of Statistics, 50(1):1– 29, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-1/
An-optimal-statistical-and-computational-framework-for-
generalized-tensor-estimation/10.1214/21-AOS2061.full.

He:2021:AIS

[HXWP21] Yinqiu He, Gongjun Xu, Chong Wu, and Wei Pan. Asymptotically independent U-statistics in high-dimensional testing. Annals of Statistics, 49(1):154–181, February 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-1/Asymptotically-independent-
U-statistics-in-high-dimensional-testing/10.1214/20-
AOS1951.full.

Heiny:2022:LDE

[HY22] Johannes Heiny and Jianfeng Yao. Limiting distributions for eigenvalues of sample correlation matrices from heavy-tailed populations. Annals of Statistics, 50(6):3249–3280, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-6/Limiting-distributions-
for-eigenvalues-of-sample-correlation-matrices-from-
heavy/10.1214/22-AOS2226.full.

Han:2023:URI

[HYF23] Xiao Han, Qing Yang, and Yingying Fan. Universal rank inference via residual subsampling with application to large networks. Annals of Statistics, 51(3):1109–1133, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-3/Universal-rank-inference-
via-residual-subsampling-with-application-to-large/10.
1214/23-AOS2282.full.

Han:2020:LDT

[HZ20] Qiyang Han and Cun-Hui Zhang. Limit distribution theory for block estimators in multiple isotonic regression. Annals of Statistics, 48(6):3251–3282, December 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-6/Limit-distribution-theory-
for-block-estimators-in-multiple-isotonic-regression/
10.1214/19-AOS1928.full.

Imbens:2021:CB

[IM21] Guido Imbens and Konrad Menzel. A causal bootstrap. Annals of Statistics, 49(3):1460–1488, June 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-3/A-causal-bootstrap/10.
1214/20-AOS2009.full.

Ing:2020:MSH

[Ing20] Ching-Kang Ing. Model selection for high-dimensional linear regression with dependent observations. Annals of Statistics, 48 (4):1959–1980, August 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-4/
Model-selection-for-high-dimensional-linear-regression-
with-dependent-observations/10.1214/19-AOS1872.full.

Jin:2021:OAS

[JKL21] Jiashun Jin, Zheng Tracy Ke, and Shengming Luo. Optimal adaptivity of signed-polygon statistics for network testing. Annals of Statistics, 49(6):3408–3433, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-6/Optimal-adaptivity-of-
signed-polygon-statistics-for-network-testing/10.1214/
21-AOS2089.full.

Jacod:2021:VC

[JLL21] Jean Jacod, Jia Li, and Zhipeng Liao. Volatility coupling. Annals of Statistics, 49(4):1982–1998, August 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-4/Volatility-coupling/
10.1214/20-AOS2023.full.

Jing:2021:CDM

[JLLX21] Bing-Yi Jing, Ting Li, Zhongyuan Lyu, and Dong Xia. Community detection on mixture multilayer networks via regularized tensor decomposition. Annals of Statistics, 49(6):3181– 3205, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-6/
Community-detection-on-mixture-multilayer-networks-via-
regularized-tensor-decomposition/10.1214/21-AOS2079.
full.

Javanmard:2020:ATL

[JMM20] Adel Javanmard, Marco Mondelli, and Andrea Montanari. Analysis of a two-layer neural network via displacement convexity. Annals of Statistics, 48(6):3619–3642, December 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-48/issue-6/Analysis-of-a-
two-layer-neural-network-via-displacement-convexity/
10.1214/20-AOS1945.full.

Johnstone:2020:THD

[JO20] Iain M. Johnstone and Alexei Onatski. Testing in highdimensional spiked models. Annals of Statistics, 48(3):1231– 1254, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-3/
Testing-in-high-dimensional-spiked-models/10.1214/18-
AOS1697.full.

Jeon:2020:ARH

[JP20] Jeong Min Jeon and Byeong U. Park. Additive regression with Hilbertian responses. Annals of Statistics, 48(5):2671– 2697, October 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-5/
Additive-regression-with-Hilbertian-responses/10.1214/
19-AOS1902.full.

Jagadeesan:2020:DET

[JPV20] Ravi Jagadeesan, Natesh S. Pillai, and Alexander Volfovsky. Designs for estimating the treatment effect in networks with interference. Annals of Statistics, 48(2):679–712, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

https://projecteuclid.org/euclid.aos/1590480030.

Jeon:2021:ARN

[JPV21] Jeong Min Jeon, Byeong U. Park, and Ingrid Van Keilegom. Additive regression for non-Euclidean responses and predictors. Annals of Statistics, 49(5):2611–2641, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-5/Additive-regression-
for-non-Euclidean-responses-and-predictors/10.1214/21-
AOS2048.full.

Jeon:2022:NRL

[JPV22] Jeong Min Jeon, Byeong U. Park, and Ingrid Van Keilegom. Nonparametric regression on Lie groups with measurement errors. Annals of Statistics, 50(5):2973–3008, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-50/issue-5/Nonparametric-
regression-on-Lie-groups-with-measurement-errors/10.
1214/22-AOS2218.full.

Javanmard:2022:PSA

[JS22] Adel Javanmard and Mahdi Soltanolkotabi. Precise statistical analysis of classification accuracies for adversarial training. Annals of Statistics, 50(4):2127–2156, August 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-4/Precise-statistical-
analysis-of-classification-accuracies-for-adversarial-
training/10.1214/22-AOS2180.full.

Jiao:2023:DNR

[JSLH23] Yuling Jiao, Guohao Shen, Yuanyuan Lin, and Jian Huang. Deep nonparametric regression on approximate manifolds: Nonasymptotic error bounds with polynomial prefactors. Annals of Statistics, 51(2):691–716, April 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
51/issue-2/Deep-nonparametric-regression-on-approximate-
manifolds--Nonasymptotic-error-bounds/10.1214/23-AOS2266.
full.

Jiang:2021:VSC

[JT21] Sheng Jiang and Surya T. Tokdar. Variable selection consistency of Gaussian process regression. Annals of Statistics, 49(5): 2491–2505, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-5/
Variable-selection-consistency-of-Gaussian-process-regression/
10.1214/20-AOS2043.full.

Jiang:2023:HDP

[JZLM23] Fei Jiang, Yeqing Zhou, Jianxuan Liu, and Yanyuan Ma. On high-dimensional Poisson models with measurement error: Hypothesis testing for nonlinear nonconvex optimization. Annals of Statistics, 51(1):233–259, February 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-

of-statistics/volume-51/issue-1/On-high-dimensional-
Poisson-models-with-measurement-error--Hypothesis/10.
1214/22-AOS2248.full.

Kirchner:2022:NSC

[KB22] Kristin Kirchner and David Bolin. Necessary and sufficient conditions for asymptotically optimal linear prediction of random fields on compact metric spaces. Annals of Statistics, 50(2): 1038–1065, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-2/
Necessary-and-sufficient-conditions-for-asymptotically-
optimal-linear-prediction-of/10.1214/21-AOS2138.full.

Kuchibhotla:2020:VPS

[KBB^{+}20] Arun K. Kuchibhotla, Lawrence D. Brown, Andreas Buja, Junhui Cai, Edward I. George, and Linda H. Zhao. Valid post-selection inference in model-free linear regression. Annals of Statistics, 48(5):2953–2981, October 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-5/Valid-post-selection-
inference-in-model-free-linear-regression/10.1214/19-
AOS1917.full.

Kennedy:2020:SIC

[KBG20] Edward H. Kennedy, Sivaraman Balakrishnan, and Max G’Sell. Sharp instruments for classifying compliers and generalizing causal effects. Annals of Statistics, 48(4):2008– 2030, August 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-4/
Sharp-instruments-for-classifying-compliers-and-generalizing-
causal-effects/10.1214/19-AOS1874.full.

Kennedy:2024:MRH

[KBRW24] Edward H. Kennedy, Sivaraman Balakrishnan, James M. Robins, and Larry Wasserman. Minimax rates for heterogeneous causal effect estimation. Annals of Statistics, 52(2):793–816, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-52/issue-2/Minimax-rates-
for-heterogeneous-causal-effect-estimation/10.1214/24-
AOS2369.full.

Kim:2020:RMN

[KBW20] Ilmun Kim, Sivaraman Balakrishnan, and Larry Wasserman. Robust multivariate nonparametric tests via projection averaging. Annals of Statistics, 48(6):3417–3441, December 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-6/Robust-multivariate-
nonparametric-tests-via-projection-averaging/10.1214/
19-AOS1936.full.

Kim:2022:MOP

[KBW22] Ilmun Kim, Sivaraman Balakrishnan, and Larry Wasserman. Minimax optimality of permutation tests. Annals of Statistics, 50 (1):225–251, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-1/
Minimax-optimality-of-permutation-tests/10.1214/21-AOS2103.
full.

Ki:2024:ML

[KFG24] Dohyeong Ki, Billy Fang, and Adityanand Guntuboyina. MARS via LASSO. Annals of Statistics, 52(3):1102–1126, June 2024. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-52/issue-3/MARS-
via-LASSO/10.1214/24-AOS2384.full.

Komarova:2023:TNS

[KH23] Tatiana Komarova and Javier Hidalgo. Testing nonparametric shape restrictions. Annals of Statistics, 51(6):2299– 2317, December 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-6/
Testing-nonparametric-shape-restrictions/10.1214/23-
AOS2311.full.

Koskela:2020:AGI

[KJJS20] Jere Koskela, Paul A. Jenkins, Adam M. Johansen, and Dario Spanò. Asymptotic genealogies of interacting particle systems with an application to sequential Monte Carlo. Annals of Statistics, 48(1):560–583, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https:

//projecteuclid.org/euclid.aos/1581930147. See erratum

[KJJS22].

Koskela:2022:EAG

[KJJS22] Jere Koskela, Paul A. Jenkins, Adam M. Johansen, and Dario Spanò. Erratum: Asymptotic genealogies of interacting particle systems with an application to sequential Monte Carlo. Annals of Statistics, 50(4):2467–2468, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-4/Erratum--Asymptotic-
genealogies-of-interacting-particle-systems-with-an/
10.1214/21-AOS2135.full. See [KJJS20].

Kolassa:2020:VFE

[KK20] John E. Kolassa and Todd A. Kuffner. On the validity of the formal Edgeworth expansion for posterior densities. Annals of Statistics, 48(4):1940–1958, August 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-4/On-the-validity-of-the-
formal-Edgeworth-expansion-for-posterior/10.1214/19-
AOS1871.full.

Klochkov:2021:RMC

[KKZ21] Yegor Klochkov, Alexey Kroshnin, and Nikita Zhivotovskiy. Robust k-means clustering for distributions with two moments. Annals of Statistics, 49(4):2206–2230, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-4/Robust-k-means-clustering-
for-distributions-with-two-moments/10.1214/20-AOS2033.
full.

Kneip:2020:ORP

[KL20] Alois Kneip and Dominik Liebl. On the optimal reconstruction of partially observed functional data. Annals of Statistics, 48 (3):1692–1717, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-3/
On-the-optimal-reconstruction-of-partially-observed-
functional-data/10.1214/19-AOS1864.full.

Kohler:2021:RCF

[KL21] Michael Kohler and Sophie Langer. On the rate of convergence of fully connected deep neural network regression estimates. Annals of Statistics, 49(4):2231–2249, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-4/On-the-rate-of-convergence-
of-fully-connected-deep-neural/10.1214/20-AOS2034.full.

Kleijn:2021:FVB

[Kle21] B. J. K. Kleijn. Frequentist validity of Bayesian limits. Annals of Statistics, 49(1):182–202, February 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-1/Frequentist-validity-
of-Bayesian-limits/10.1214/20-AOS1952.full.

Koltchinskii:2020:EEL

[KLN20] Vladimir Koltchinskii, Matthias Löffler, and Richard Nickl. Efficient estimation of linear functionals of principal components. Annals of Statistics, 48(1):464–490, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

https://projecteuclid.org/euclid.aos/1581930143.

Kolaczyk:2020:AUN

[KLR^{+}20] Eric D. Kolaczyk, Lizhen Lin, Steven Rosenberg, Jackson Walters, and Jie Xu. Averages of unlabeled networks: Geometric characterization and asymptotic behavior. Annals of Statistics, 48(1):514–538, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https:/

/projecteuclid.org/euclid.aos/1581930145.

Kim:2020:PDR

[KLYL20] Kyongwon Kim, Bing Li, Zhou Yu, and Lexin Li. On post dimension reduction statistical inference. Annals of Statistics, 48 (3):1567–1592, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-3/
On-post-dimension-reduction-statistical-inference/10.
1214/19-AOS1859.full.

Kpotufe:2021:MSB

[KM21a] Samory Kpotufe and Guillaume Martinet. Marginal singularity and the benefits of labels in covariate-shift. Annals of Statistics, 49(6):3299–3323, December 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-6/Marginal-singularity-
and-the-benefits-of-labels-in-covariate-shift/10.1214/
21-AOS2084.full.

Kwon:2021:ENC

[KM21b] Caleb Kwon and Eric Mbakop. Estimation of the number of components of nonparametric multivariate finite mixture models. Annals of Statistics, 49(4):2178–2205, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-49/issue-4/Estimation-of-
the-number-of-components-of-nonparametric-multivariate-
finite/10.1214/20-AOS2032.full.

Konig:2020:MMS

[KMW20] Claudia König, Axel Munk, and Frank Werner. Multidimensional multiscale scanning in exponential families: Limit theory and statistical consequences. Annals of Statistics, 48(2):655–678, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/euclid.aos/
1590480029.

Kulaitis:2021:WRS

[KMW21] Gytis Kulaitis, Axel Munk, and Frank Werner. What is resolution? A statistical minimax testing perspective on superresolution microscopy. Annals of Statistics, 49(4):2292–2312, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-49/issue-4/What-
is-resolution-A-statistical-minimax-testing-perspective-
on-superresolution/10.1214/20-AOS2037.full.

Kim:2022:LPT

[KNBW22] Ilmun Kim, Matey Neykov, Sivaraman Balakrishnan, and Larry Wasserman. Local permutation tests for conditional independence. Annals of Statistics, 50(6):3388–3414, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-6/Local-permutation-tests-
for-conditional-independence/10.1214/22-AOS2233.full.

Koltchinskii:2022:ESF

[Kol22] Vladimir Koltchinskii. Estimation of smooth functionals in highdimensional models: Bootstrap chains and Gaussian approximation. Annals of Statistics, 50(4):2386–2415, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-4/Estimation-of-smooth-
functionals-in-high-dimensional-models--Bootstrap/10.
1214/22-AOS2197.full.

Kuchibhotla:2022:LSE

[KP22] Arun K. Kuchibhotla and Rohit K. Patra. On least squares estimation under heteroscedastic and heavy-tailed errors. Annals of Statistics, 50(1):277–302, February 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-1/On-least-squares-estimation-
under-heteroscedastic-and-heavy-tailed-errors/10.1214/
21-AOS2105.full.

Konen:2023:SQH

[KP23] Dimitri Konen and Davy Paindaveine. Spatial quantiles on the hypersphere. Annals of Statistics, 51(5):2221–2245, October 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-51/issue-5/Spatial-quantiles-
on-the-hypersphere/10.1214/23-AOS2332.full.

Klopp:2023:ATD

[KPST23] Olga Klopp, Maxim Panov, Suzanne Sigalla, and Alexandre B. Tsybakov. Assigning topics to documents by successive projections. Annals of Statistics, 51(5):1989–2014, October 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-5/Assigning-topics-to-
documents-by-successive-projections/10.1214/23-AOS2316.
full.

Katsevich:2020:SHP

[KR20] Eugene Katsevich and Aaditya Ramdas. Simultaneous highprobability bounds on the false discovery proportion in structured, regression and online settings. Annals of Statistics, 48(6): 3465–3487, December 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-6/
Simultaneous-high-probability-bounds-on-the-false-discovery-
proportion-in/10.1214/19-AOS1938.full.

Katsevich:2024:AAG

[KR24] Anya Katsevich and Philippe Rigollet. On the approximation accuracy of Gaussian variational inference. Annals of Statistics, 52(4):1384–1409, August 2024. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-4/On-the-approximation-
accuracy-of-Gaussian-variational-inference/10.1214/24-
AOS2393.full.

Kim:2021:CAP

[KRSW21] Ilmun Kim, Aaditya Ramdas, Aarti Singh, and Larry Wasserman. Classification accuracy as a proxy for two-sample testing. Annals of Statistics, 49(1):411–434, February 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-1/Classification-accuracy-
as-a-proxy-for-two-sample-testing/10.1214/20-AOS1962.
full.

Karmakar:2020:AEC

[KS20] Bikram Karmakar and Dylan S. Small. Assessment of the extent of corroboration of an elaborate theory of a causal hypothesis using partial conjunctions of evidence factors. Annals of Statistics, 48(6):3283–3311, December 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-6/Assessment-of-the-extent-
of-corroboration-of-an-elaborate-theory/10.1214/19-AOS1929.
full.

Kutyniok:2020:DSN

[Kut20] Gitta Kutyniok. Discussion of: “Nonparametric regression using deep neural networks with ReLU activation function”. Annals of Statistics, 48(4):1902–1905, August 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-4/Discussion-of--Nonparametric-

regression-using-deep-neural-networks-with/10.1214/19-
AOS1911.full. See [SH20a].

Kou:2022:LSI

[KW22] Jiyao Kou and Guenther Walther. Large-scale inference with block structure. Annals of Statistics, 50(3):1541– 1572, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-3/
Large-scale-inference-with-block-structure/10.1214/21-
AOS2162.full.

Kong:2021:AED

[KYZ21] Xiangshun Kong, Mingao Yuan, and Wei Zheng. Approximate and exact designs for total effects. Annals of Statistics, 49(3): 1594–1625, June 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-3/
Approximate-and-exact-designs-for-total-effects/10.1214/
20-AOS2015.full.

Koltchinskii:2021:ESF

[KZ21] Vladimir Koltchinskii and Mayya Zhilova. Estimation of smooth functionals in normal models: Bias reduction and asymptotic efficiency. Annals of Statistics, 49(5):2577–2610, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-5/Estimation-of-smooth-
functionals-in-normal-models--Bias-reduction/10.1214/
20-AOS2047.full.

Lahiri:2021:NSC

[Lah21] Soumendra N. Lahiri. Necessary and sufficient conditions for variable selection consistency of the LASSO in high dimensions. Annals of Statistics, 49(2):820–844, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-2/Necessary-and-sufficient-
conditions-for-variable-selection-consistency-of-the/
10.1214/20-AOS1979.full.

Li:2020:AGH

[LAP^{+}20] Haoran Li, Alexander Aue, Debashis Paul, Jie Peng, and Pei Wang. An adaptable generalization of Hotelling’s.

Annals of Statistics, 48(3):1815–1847, June 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-3/An-adaptable-generalization-
of-Hotellings-T2-test-in-high-dimension/10.1214/19-AOS1869.
full.

Luedtke:2021:CNS

[LBvdL21] Alex Luedtke, Aurélien Bibaut, and Mark van der Laan. Correction note: “Statistical inference for the mean outcome under a possibly nonunique optimal treatment rule”. Annals of Statistics, 49(4):2429–2430, August 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-4/Correction-note--Statistical-
inference-for-the-mean-outcome-under/10.1214/20-AOS2031.
full.

Li:2023:RLD

[LBY23] Mengchu Li, Thomas B. Berrett, and Yi Yu. On robustness and local differential privacy. Annals of Statistics, 51 (2):717–737, April 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-2/
On-robustness-and-local-differential-privacy/10.1214/
23-AOS2267.full.

Luedtke:2024:OSE

[LC24] Alex Luedtke and Incheoul Chung. One-step estimation of differentiable Hilbert-valued parameters. Annals of Statistics, 52 (4):1534–1563, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-4/
One-step-estimation-of-differentiable-Hilbert-valued-
parameters/10.1214/24-AOS2403.full.

Lee:2021:BNH

[LCI21] Donald K. K. Lee, Ningyuan Chen, and Hemant Ishwaran. Boosted nonparametric hazards with time-dependent covariates. Annals of Statistics, 49(4):2101–2128, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-4/Boosted-nonparametric-

hazards-with-time-dependent-covariates/10.1214/20-AOS2028.
full.

Lei:2021:NRU

[Lei21] Jing Lei. Network representation using graph root distributions. Annals of Statistics, 49(2):745–768, April 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-2/Network-representation-
using-graph-root-distributions/10.1214/20-AOS1976.full.

Leung:2022:ROC

[Leu22] Michael P. Leung. Rate-optimal cluster-randomized designs for spatial interference. Annals of Statistics, 50(5):3064–3087, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-50/issue-5/Rate-optimal-
cluster-randomized-designs-for-spatial-interference/
10.1214/22-AOS2224.full.

Lalancette:2021:RBE

[LEV21] Michaël Lalancette, Sebastian Engelke, and Stanislav Volgushev. Rank-based estimation under asymptotic dependence and independence, with applications to spatial extremes. Annals of Statistics, 49(5):2552–2576, October 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-5/Rank-based-estimation-
under-asymptotic-dependence-and-independence-with-applications/
10.1214/20-AOS2046.full.

Lin:2021:ITS

[LFR21] Zhantao Lin, Nancy Flournoy, and William F. Rosenberger. Inference for a two-stage enrichment design. Annals of Statistics, 49 (5):2697–2720, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-5/
Inference-for-a-two-stage-enrichment-design/10.1214/
21-AOS2051.full.

Lopuhaa:2023:ELM

[LGRG23] Hendrik Paul Lopuhaä, Valerie Gares, and Anne Ruiz-Gazen. S-estimation in linear models with structured covariance matrices. Annals of Statistics, 51(6):2415–2439, December 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-51/issue-6/S-estimation-
in-linear-models-with-structured-covariance-matrices/
10.1214/23-AOS2334.full.

Liu:2021:MRS

[LGS21] Haoyang Liu, Chao Gao, and Richard J. Samworth. Minimax rates in sparse, high-dimensional change point detection. Annals of Statistics, 49(2):1081–1112, April 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-2/Minimax-rates-in-sparse-
high-dimensional-change-point-detection/10.1214/20-AOS1994.
full.

Liu:2023:IUC

[LH23] Yang Liu and Feifang Hu. The impacts of unobserved covariates on covariate-adaptive randomized experiments. Annals of Statistics, 51(5):1895–1920, October 2023. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-5/The-impacts-of-unobserved-
covariates-on-covariate-adaptive-randomized-experiments/
10.1214/23-AOS2308.full.

Liu:2023:CLL

[LHBS23] Zhijun Liu, Jiang Hu, Zhidong Bai, and Haiyan Song. A CLT for the LSS of large-dimensional sample covariance matrices with diverging spikes. Annals of Statistics, 51(5):2246–2271, October 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-51/issue-5/A-CLT-for-the-
LSS-of-large-dimensional-sample-covariance/10.1214/23-
AOS2333.full.

Li:2020:AJD

[LHY20] Zeng Li, Fang Han, and Jianfeng Yao. Asymptotic joint distribution of extreme eigenvalues and trace of large sample covariance matrix in a generalized spiked population model. Annals of Statistics, 48(6):3138–3160, December 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-6/Asymptotic-joint-distribution-

of-extreme-eigenvalues-and-trace-of-large/10.1214/19-
AOS1882.full.

Li:2022:BFD

[Li22] Cheng Li. Bayesian fixed-domain asymptotics for covariance parameters in a Gaussian process model. Annals of Statistics, 50(6): 3334–3363, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-6/
Bayesian-fixed-domain-asymptotics-for-covariance-parameters-
in-a-Gaussian/10.1214/22-AOS2230.full.

Lecue:2020:RML

[LL20] Guillaume Lecué and Matthieu Lerasle. Robust machine learning by median-of-means: Theory and practice. Annals of Statistics, 48(2):906–931, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/euclid.aos/1590480039.

Lee:2022:FSD

[LL22] Kuang-Yao Lee and Lexin Li. Functional sufficient dimension reduction through average Fréchet derivatives. Annals of Statistics, 50(2):904–929, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-2/
Functional-sufficient-dimension-reduction-through-average-
Fr%c3%a9chet-derivatives/10.1214/21-AOS2131.full.

Lin:2021:OSI

[LLHL21] Qian Lin, Xinran Li, Dongming Huang, and Jun S. Liu. On the optimality of sliced inverse regression in high dimensions. Annals of Statistics, 49(1):1–20, February 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-1/On-the-optimality-of-
sliced-inverse-regression-in-high-dimensions/10.1214/
19-AOS1813.full.

Lopes:2020:BMS

[LLM20] Miles E. Lopes, Zhenhua Lin, and Hans-Georg Müller. Bootstrapping max statistics in high dimensions: Near-parametric rates under weak variance decay and application to functional and multinomial data. Annals of Statistics, 48(2):1214–1229, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL https://projecteuclid.org/euclid.
aos/1590480052.

Li:2024:DES

[LLS24] Degui Li, Runze Li, and Han Lin Shang. Detection and estimation of structural breaks in high-dimensional functional time series. Annals of Statistics, 52(4):1716–1740, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-52/issue-4/Detection-and-
estimation-of-structural-breaks-in-high-dimensional-
functional/10.1214/24-AOS2414.full.

Lee:2021:FDT

[LLSS21] Sokbae Lee, Yuan Liao, Myung Hwan Seo, and Youngki Shin. Factor-driven two-regime regression. Annals of Statistics, 49 (3):1656–1678, June 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-3/
Factor-driven-two-regime-regression/10.1214/20-AOS2017.
full.

Lin:2021:TVR

[LM21a] Zhenhua Lin and Hans-Georg Müller. Total variation regularized Fréchet regression for metric-space valued data. Annals of Statistics, 49(6):3510–3533, December 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-6/Total-variation-regularized-
Fr%c3%a9chet-regression-for-metric-space-valued-data/
10.1214/21-AOS2095.full.

Lugosi:2021:RMM

[LM21b] Gábor Lugosi and Shahar Mendelson. Robust multivariate mean estimation: The optimality of trimmed mean. Annals of Statistics, 49(1):393–410, February 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-1/Robust-multivariate-
mean-estimation-The-optimality-of-trimmed-mean/10.1214/
20-AOS1961.full.

Lahiry:2022:MNE

[LN22] Samriddha Lahiry and Michael Nussbaum. Minimax nonparametric estimation of pure quantum states. Annals of Statistics, 50 (1):430–459, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-1/
Minimax-nonparametric-estimation-of-pure-quantum-states/
10.1214/21-AOS2115.full.

Li:2023:GCO

[LO23] Harrison H. Li and Art B. Owen. A general characterization of optimal tie-breaker designs. Annals of Statistics, 51 (3):1030–1057, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-3/
A-general-characterization-of-optimal-tie-breaker-designs/
10.1214/23-AOS2275.full.

Lopes:2022:CLT

[Lop22] Miles E. Lopes. Central limit theorem and bootstrap approximation in high dimensions: Near. Annals of Statistics, 50(5): 2492–2513, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-5/
Central-limit-theorem-and-bootstrap-approximation-in-
high-dimensions/10.1214/22-AOS2184.full.

Lee:2024:EFL

[LPMP24] Eun Ryung Lee, Seyoung Park, Enno Mammen, and Byeong U. Park. Efficient functional Lasso kernel smoothing for highdimensional additive regression. Annals of Statistics, 52(4): 1741–1773, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-4/
Efficient-functional-Lasso-kernel-smoothing-for-high-
dimensional-additive-regression/10.1214/24-AOS2415.full.

Li:2024:SAG

[LPYZ24] Huiqin Li, Guangming Pan, Yanqing Yin, and Wang Zhou. Spectral analysis of gram matrices with missing at random observations: Convergence, central limit theorems, and applications in statistical inference. Annals of Statistics, 52(3):

1254–1275, June 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-3/
Spectral-analysis-of-gram-matrices-with-missing-at-random-
observations/10.1214/24-AOS2392.full.

Liao:2022:BPL

[LQW^{+}22] Peng Liao, Zhengling Qi, Runzhe Wan, Predrag Klasnja, and Susan A. Murphy. Batch policy learning in average reward Markov decision processes. Annals of Statistics, 50(6):3364–3387, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-50/issue-6/Batch-policy-
learning-in-average-reward-Markov-decision-processes/
10.1214/22-AOS2231.full.

Liang:2020:JIK

[LR20a] Tengyuan Liang and Alexander Rakhlin. Just interpolate: Kernel “Ridgeless” regression can generalize. Annals of Statistics, 48(3):1329–1347, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-3/
Just-interpolate-Kernel-Ridgeless-regression-can-generalize/
10.1214/19-AOS1849.full.

Liu:2020:MEL

[LR20b] Yu Liu and Zhao Ren. Minimax estimation of large precision matrices with bandable Cholesky factor. Annals of Statistics, 48 (4):2428–2454, August 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-4/
Minimax-estimation-of-large-precision-matrices-with-
bandable-Cholesky-factor/10.1214/19-AOS1893.full.

Li:2022:DRF

[LS22a] Bing Li and Jun Song. Dimension reduction for functional data based on weak conditional moments. Annals of Statistics, 50 (1):107–128, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-1/
Dimension-reduction-for-functional-data-based-on-weak-
conditional-moments/10.1214/21-AOS2091.full.

Liang:2022:PHD

[LS22b] Tengyuan Liang and Pragya Sur. A precise high-dimensional asymptotic theory for boosting and minimum-ℓ_{1}-norma interpolated classifiers. Annals of Statistics, 50(3):1669–1695, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-3/A-precise-high-dimensional-
asymptotic-theory-for-boosting-and-minimum/10.1214/22-
AOS2170.full.

Li:2024:SSC

[LSC^{+}24] Gen Li, Laixi Shi, Yuxin Chen, Yuejie Chi, and Yuting Wei. Settling the sample complexity of model-based offline reinforcement learning. Annals of Statistics, 52(1):233–260, February 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-1/Settling-the-sample-
complexity-of-model-based-offline-reinforcement-learning/
10.1214/23-AOS2342.full.

Lee:2020:CCB

[LSV20] Anthony Lee, Sumeetpal S. Singh, and Matti Vihola. Coupled conditional backward sampling particle filter. Annals of Statistics, 48(5):3066–3089, October 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-5/Coupled-conditional-
backward-sampling-particle-filter/10.1214/19-AOS1922.
full.

Loh:2021:FDA

[LSW21] Wei-Liem Loh, Saifei Sun, and Jun Wen. On fixed-domain asymptotics, parameter estimation and isotropic Gaussian random fields with Matérn covariance functions. Annals of Statistics, 49(6):3127–3152, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
49/issue-6/On-fixed-domain-asymptotics-parameter-estimation-
and-isotropic-Gaussian-random/10.1214/21-AOS2077.full.

Laha:2024:FOD

[LSWMC24] Nilanjana Laha, Aaron Sonabend-W, Rajarshi Mukherjee, and Tianxi Cai. Finding the optimal dynamic treatment regimes using smooth Fisher consistent surrogate loss. Annals of Statistics, 52(2):679–707, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-2/
Finding-the-optimal-dynamic-treatment-regimes-using-
smooth-Fisher-consistent/10.1214/24-AOS2363.full.

Liu:2024:MVD

[LTMC24] Weidong Liu, Jiyuan Tu, Xiaojun Mao, and Xi Chen. Majority vote for distributed differentially private sign selection. Annals of Statistics, 52(4):1671–1690, August 2024. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-4/Majority-vote-for-distributed-
differentially-private-sign-selection/10.1214/24-AOS2411.
full.

Lauritzen:2021:TPE

[LUZ21] Steffen Lauritzen, Caroline Uhler, and Piotr Zwiernik. Total positivity in exponential families with application to binary variables. Annals of Statistics, 49(3):1436–1459, June 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-3/Total-positivity-in-
exponential-families-with-application-to-binary-variables/
10.1214/20-AOS2007.full.

Ledoit:2020:ANS

[LW20] Olivier Ledoit and Michael Wolf. Analytical nonlinear shrinkage of large-dimensional covariance matrices. Annals of Statistics, 48 (5):3043–3065, October 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-5/
Analytical-nonlinear-shrinkage-of-large-dimensional-
covariance-matrices/10.1214/19-AOS1921.full.

Li:2022:RGA

[LW22] Shuangning Li and Stefan Wager. Random graph asymptotics for treatment effect estimation under network interference. Annals of Statistics, 50(4):2334–2358, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-50/issue-4/Random-graph-

asymptotics-for-treatment-effect-estimation-under-network-
interference/10.1214/22-AOS2191.full.

Li:2021:CLT

[LWL21] Zeng Li, Qinwen Wang, and Runze Li. Central limit theorem for linear spectral statistics of large dimensional Kendall’s rank correlation matrices and its applications. Annals of Statistics, 49 (3):1569–1593, June 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-3/
Central-limit-theorem-for-linear-spectral-statistics-
of-large-dimensional/10.1214/20-AOS2013.full.

Lyu:2023:OEC

[LX23] Zhongyuan Lyu and Dong Xia. Optimal estimation and computational limit of low-rank Gaussian mixtures. Annals of Statistics, 51(2):646–667, April 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
51/issue-2/Optimal-estimation-and-computational-limit-
of-low-rank-Gaussian-mixtures/10.1214/23-AOS2264.full.

Lee:2020:BCR

[LY20] Stephen M. S. Lee and Puyudi Yang. Bootstrap confidence regions based on M-estimators under nonstandard conditions. Annals of Statistics, 48(1):274–299, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

https://projecteuclid.org/euclid.aos/1581930135.

Lauritzen:2022:LAG

[LZ22a] Steffen Lauritzen and Piotr Zwiernik. Locally associated graphical models and mixed convex exponential families. Annals of Statistics, 50(5):3009–3038, October 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-5/Locally-associated-graphical-
models-and-mixed-convex-exponential-families/10.1214/
22-AOS2219.full.

Luo:2022:TCP

[LZ22b] Yuetian Luo and Anru R. Zhang. Tensor clustering with planted structures: Statistical optimality and computational limits. Annals of Statistics, 50(1):584–613, February 2022.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-1/Tensor-clustering-with-
planted-structures--Statistical-optimality-and-computational/
10.1214/21-AOS2123.full.

Loffler:2021:OSC

[LZZ21] Matthias Löffler, Anderson Y. Zhang, and Harrison H. Zhou. Optimality of spectral clustering in the Gaussian mixture model. Annals of Statistics, 49(5):2506–2530, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-49/issue-5/Optimality-of-
spectral-clustering-in-the-Gaussian-mixture-model/10.
1214/20-AOS2044.full.

Manole:2024:PES

[MBNWW24] Tudor Manole, Sivaraman Balakrishnan, Jonathan Niles-Weed, and Larry Wasserman. Plugin estimation of smooth optimal transport maps. Annals of Statistics, 52(3):966–998, June 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-3/Plugin-estimation-of-
smooth-optimal-transport-maps/10.1214/24-AOS2379.full.

Mukherjee:2021:OLD

[MBR21] Debarghya Mukherjee, Moulinath Banerjee, and Ya’acov Ritov. Optimal linear discriminators for the discrete choice model in growing dimensions. Annals of Statistics, 49(6):3324– 3357, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-6/
Optimal-linear-discriminators-for-the-discrete-choice-
model-in-growing/10.1214/21-AOS2085.full.

Mourtada:2020:MOR

[MGS20] Jaouad Mourtada, Stéphane Gaïffas, and Erwan Scornet. Minimax optimal rates for Mondrian trees and forests. Annals of Statistics, 48(4):2253–2276, August 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-4/Minimax-optimal-rates-
for-Mondrian-trees-and-forests/10.1214/19-AOS1886.full.

Mogensen:2020:MEM

[MH20] Søren Wengel Mogensen and Niels Richard Hansen. Markov equivalence of marginalized local independence graphs. Annals of Statistics, 48(1):539–559, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1581930146.

McCormack:2022:SEF

[MH22] Andrew McCormack and Peter Hoff. The Stein effect for Fréchet means. Annals of Statistics, 50(6):3647–3676, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-6/The-Stein-effect-for-
Fr%c3%a9chet-means/10.1214/22-AOS2245.full.

Mukherjee:2022:MOS

[MJ22] Gourab Mukherjee and Iain M. Johnstone. On minimax optimality of sparse Bayes predictive density estimates. Annals of Statistics, 50(1):81–106, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
50/issue-1/On-minimax-optimality-of-sparse-Bayes-predictive-
density-estimates/10.1214/21-AOS2086.full.

Manole:2021:ENC

[MK21a] Tudor Manole and Abbas Khalili. Estimating the number of components in finite mixture models via the Group–Sort–Fuse procedure. Annals of Statistics, 49(6):3043–3069, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-6/Estimating-the-number-
of-components-in-finite-mixture-models-via/10.1214/21-
AOS2072.full.

Morikawa:2021:SOE

[MK21b] Kosuke Morikawa and Jae Kwang Kim. Semiparametric optimal estimation with nonignorable nonresponse data. Annals of Statistics, 49(5):2991–3014, October 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-5/Semiparametric-optimal-
estimation-with-nonignorable-nonresponse-data/10.1214/
21-AOS2070.full.

Marandon:2024:AND

[MLMR24] Ariane Marandon, Lihua Lei, David Mary, and Etienne Roquain. Adaptive novelty detection with false discovery rate guarantee. Annals of Statistics, 52(1):157–183, February 2024. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-1/Adaptive-novelty-detection-
with-false-discovery-rate-guarantee/10.1214/23-AOS2338.
full.

Miolane:2021:DLU

[MM21] Léo Miolane and Andrea Montanari. The distribution of the Lasso: Uniform control over sparse balls and adaptive parameter tuning. Annals of Statistics, 49(4):2313– 2335, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-4/
The-distribution-of-the-Lasso--Uniform-control-over-
sparse/10.1214/20-AOS2038.full.

Marion:2023:FSC

[MMS23] Joe Marion, Joseph Mathews, and Scott C. Schmidler. Finite-sample complexity of sequential Monte Carlo estimators. Annals of Statistics, 51(3):1357–1375, June 2023. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-3/Finite-sample-complexity-
of-sequential-Monte-Carlo-estimators/10.1214/23-AOS2295.
full.

McGoff:2020:ERM

[MN20] Kevin McGoff and Andrew B. Nobel. Empirical risk minimization and complexity of dynamical models. Annals of Statistics, 48 (4):2031–2054, August 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-4/
Empirical-risk-minimization-and-complexity-of-dynamical-
models/10.1214/19-AOS1876.full.

Monard:2021:SGB

[MNP21] François Monard, Richard Nickl, and Gabriel P. Paternain. Statistical guarantees for Bayesian uncertainty quantification in nonlinear inverse problems with Gaussian process priors.

Annals of Statistics, 49(6):3255–3298, December 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-6/Statistical-guarantees-
for-Bayesian-uncertainty-quantification-in-nonlinear-
inverse-problems/10.1214/21-AOS2082.full.

Mourtada:2022:EMR

[Mou22] Jaouad Mourtada. Exact minimax risk for linear least squares, and the lower tail of sample covariance matrices. Annals of Statistics, 50(4):2157–2178, August 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-4/Exact-minimax-risk-for-
linear-least-squares-and-the-lower/10.1214/22-AOS2181.
full.

Mies:2023:EMF

[MP23] Fabian Mies and Mark Podolskij. Estimation of mixed fractional stable processes using high-frequency data. Annals of Statistics, 51(5):1946–1964, October 2023. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-5/Estimation-of-mixed-
fractional-stable-processes-using-high-frequency-data/
10.1214/23-AOS2312.full.

Meyer:2020:EVF

[MPK20] Marco Meyer, Efstathios Paparoditis, and Jens-Peter Kreiss. Extending the validity of frequency domain bootstrap methods to general stationary processes. Annals of Statistics, 48 (4):2404–2427, August 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-4/
Extending-the-validity-of-frequency-domain-bootstrap-
methods-to-general/10.1214/19-AOS1892.full.

Mao:2020:TOE

[MPW20] Cheng Mao, Ashwin Pananjady, and Martin J. Wainwright. Towards optimal estimation of bivariate isotonic matrices with unknown permutations. Annals of Statistics, 48(6):3183– 3205, December 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-6/
Towards-optimal-estimation-of-bivariate-isotonic-matrices-
with-unknown-permutations/10.1214/19-AOS1925.full.

Ma:2023:OTC

[MPW23] Cong Ma, Reese Pathak, and Martin J. Wainwright. Optimally tackling covariate shift in RKHS-based nonparametric regression. Annals of Statistics, 51(2):738–761, April 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-2/Optimally-tackling-covariate-
shift-in-RKHS-based-nonparametric-regression/10.1214/
23-AOS2268.full.

Maruyama:2020:ABE

[MS20] Yuzo Maruyama and William E. Strawderman. Admissible Bayes equivariant estimation of location vectors for spherically symmetric distributions with unknown scale. Annals of Statistics, 48 (2):1052–1071, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/euclid.aos/1590480045.

Mure:2021:PRP

[Mur21] Joseph Muré. Propriety of the reference posterior distribution in Gaussian process modeling. Annals of Statistics, 49(4): 2356–2377, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-4/
Propriety-of-the-reference-posterior-distribution-in-
Gaussian-process-modeling/10.1214/20-AOS2040.full.

Montanari:2021:ELR

[MV21] Andrea Montanari and Ramji Venkataramanan. Estimation of low-rank matrices via approximate message passing. Annals of Statistics, 49(1):321–345, February 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-1/Estimation-of-low-rank-
matrices-via-approximate-message-passing/10.1214/20-
AOS1958.full.

Mao:2022:LMP

[MW22] Cheng Mao and Yihong Wu. Learning mixtures of permutations: Groups of pairwise comparisons and combinatorial method of moments. Annals of Statistics, 50(4):2231–2255, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-4/Learning-mixtures-of-
permutations--Groups-of-pairwise-comparisons-and/10.
1214/22-AOS2185.full.

Montanari:2024:FLL

[MW24] Andrea Montanari and Yuchen Wu. Fundamental limits of low-rank matrix estimation with diverging aspect ratios. Annals of Statistics, 52(4):1460–1484, August 2024. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-4/Fundamental-limits-of-
low-rank-matrix-estimation-with-diverging-aspect/10.
1214/24-AOS2400.full.

Ma:2023:BWC

[MWW23] Xinwei Ma, Jingshen Wang, and Chong Wu. Breaking the winner’s curse in Mendelian randomization: Rerandomized inverse variance weighted estimator. Annals of Statistics, 51(1): 211–232, February 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-1/
Breaking-the-winners-curse-in-Mendelian-randomization-
-Rerandomized-inverse/10.1214/22-AOS2247.full.

Mei:2023:SVD

[MWY23] Tianxing Mei, Chen Wang, and Jianfeng Yao. On singular values of data matrices with general independent columns. Annals of Statistics, 51(2):624–645, April 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
51/issue-2/On-singular-values-of-data-matrices-with-
general-independent-columns/10.1214/23-AOS2263.full.

Mendelson:2020:RCE

[MZ20] Shahar Mendelson and Nikita Zhivotovskiy. Robust covariance estimation under. Annals of Statistics, 48(3):1648– 1664, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-3/
Robust-covariance-estimation-under-L_4-L_2-norm-equivalence/
10.1214/19-AOS1862.full.

Montanari:2022:IPT

[MZ22] Andrea Montanari and Yiqiao Zhong. The interpolation phase transition in neural networks: Memorization and generalization under lazy training. Annals of Statistics, 50(5):2816–2847, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-50/issue-5/The-
interpolation-phase-transition-in-neural-networks--Memorization-
and/10.1214/22-AOS2211.full.

Neykov:2021:MOC

[NBW21] Matey Neykov, Sivaraman Balakrishnan, and Larry Wasserman. Minimax optimal conditional independence testing. Annals of Statistics, 49(4):2151–2177, August 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-4/Minimax-optimal-conditional-
independence-testing/10.1214/20-AOS2030.full.

Niu:2024:RMX

[NCDK24] Ziang Niu, Abhinav Chakraborty, Oliver Dukes, and Eugene Katsevich. Reconciling model-X and doubly robust approaches to conditional independence testing. Annals of Statistics, 52 (3):895–921, June 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-3/
Reconciling-model-X-and-doubly-robust-approaches-to-
conditional-independence/10.1214/24-AOS2372.full.

Ndaoud:2022:SOR

[Nda22] Mohamed Ndaoud. Sharp optimal recovery in the two component Gaussian mixture model. Annals of Statistics, 50(4): 2096–2126, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-4/
Sharp-optimal-recovery-in-the-two-component-Gaussian-
mixture-model/10.1214/22-AOS2178.full.

Nickl:2024:CID

[Nic24] Richard Nickl. Consistent inference for diffusions from low frequency measurements. Annals of Statistics, 52(2): 519–549, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-2/
Consistent-inference-for-diffusions-from-low-frequency-
measurements/10.1214/24-AOS2357.full.

Najafi:2021:SLS

[NIS^{+}21] Amir Najafi, Saeed Ilchi, Amir Hossein Saberi, Seyed Abolfazl Motahari, Babak H. Khalaj, and Hamid R. Rabiee. On statistical learning of simplices: Unmixing problem revisited. Annals of Statistics, 49(3):1626–1655, June 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-3/On-statistical-learning-
of-simplices-Unmixing-problem-revisited/10.1214/20-AOS2016.
full.

Nevo:2021:ASL

[NLS21] Daniel Nevo, Judith J. Lok, and Donna Spiegelman. Analysis of “learn-as-you-go” (LAGO) studies. Annals of Statistics, 49 (2):793–819, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-2/
Analysis-of-learn-as-you-go-LAGO-studies/10.1214/20-
AOS1978.full.

Nickl:2020:NSI

[NR20] Richard Nickl and Kolyan Ray. Nonparametric statistical inference for drift vector fields of multi-dimensional diffusions. Annals of Statistics, 48(3):1383–1408, June 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-3/Nonparametric-statistical-
inference-for-drift-vector-fields-of-multi-dimensional/
10.1214/19-AOS1851.full.

Nickl:2024:PCD

[NT24] Richard Nickl and Edriss S. Titi. On posterior consistency of data assimilation with Gaussian process priors: The 2D-Navier– Stokes equations. Annals of Statistics, 52(4):1825–1844, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-52/issue-4/On-posterior-
consistency-of-data-assimilation-with-Gaussian-process-
priors/10.1214/24-AOS2427.full.

Niles-Weed:2022:MES

[NWB22] Jonathan Niles-Weed and Quentin Berthet. Minimax estimation of smooth densities in Wasserstein distance. Annals of Statistics, 50(3):1519–1540, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-3/
Minimax-estimation-of-smooth-densities-in-Wasserstein-
distance/10.1214/21-AOS2161.full.

Ohn:2024:AVB

[OL24] Ilsang Ohn and Lizhen Lin. Adaptive variational Bayes: Optimality, computation and applications. Annals of Statistics, 52 (1):335–363, February 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-1/
Adaptive-variational-Bayes-Optimality-computation-and-
applications/10.1214/23-AOS2349.full.

Oh:2022:GEB

[OQC22] Eun Jeong Oh, Min Qian, and Ying Kuen Cheung. Generalization error bounds of dynamic treatment regimes in penalized regression-based learning. Annals of Statistics, 50(4): 2047–2071, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-4/
Generalization-error-bounds-of-dynamic-treatment-regimes-
in-penalized-regression/10.1214/22-AOS2171.full.

Ortelli:2021:PBH

[OvdG21] Francesco Ortelli and Sara van de Geer. Prediction bounds for higher order total variation regularized least squares. Annals of Statistics, 49(5):2755–2773, October 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-5/Prediction-bounds-for-
higher-order-total-variation-regularized-least-squares/
10.1214/21-AOS2054.full.

Obloj:2021:RES

[OW21] Jan Obłój and Johannes Wiesel. Robust estimation of superhedging prices. Annals of Statistics, 49(1):508–530, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-1/Robust-estimation-of-
superhedging-prices/10.1214/20-AOS1966.full.

Panigrahi:2023:CMF

[Pan23] Snigdha Panigrahi. Carving model-free inference. Annals of Statistics, 51(6):2318–2341, December 2023. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-6/Carving-model-free-inference/
10.1214/23-AOS2318.full.

Paul:2020:SMF

[PC20] Subhadeep Paul and Yuguo Chen. Spectral and matrix factorization methods for consistent community detection in multilayer networks. Annals of Statistics, 48(1):230–250, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/euclid.aos/
1581930133.

Pilliat:2023:OPE

[PCV23] Emmanuel Pilliat, Alexandra Carpentier, and Nicolas Verzelen. Optimal permutation estimation in CrowdSourcing problems. Annals of Statistics, 51(3):935–961, June 2023. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-3/Optimal-Permutation-
Estimation-in-CrowdSourcing-problems/10.1214/23-AOS2271.
full.

Pompe:2020:FAM

[PHL20] Emilia Pompe, Chris Holmes, and Krzysztof Latuszyński. A framework for adaptive MCMC targeting multimodal distributions. Annals of Statistics, 48(5):2930–2952, October 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-5/A-framework-for-adaptive-
MCMC-targeting-multimodal-distributions/10.1214/19-AOS1916.
full.

Petersen:2021:WTC

[PLD21] Alexander Petersen, Xi Liu, and Afshin A. Divani. Wasserstein F -tests and confidence bands for the Fréchet regression of density response curves. Annals of Statistics, 49(1):

590–611, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-1/
Wasserstein-F-tests-and-confidence-bands-for-the-Fr%
c3%a9chet-regression/10.1214/20-AOS1971.full.

Pananjady:2020:WCV

[PMM^{+}20] Ashwin Pananjady, Cheng Mao, Vidya Muthukumar, Martin J. Wainwright, and Thomas A. Courtade. Worst-case versus average-case design for estimation from partial pairwise comparisons. Annals of Statistics, 48(2):1072–1097, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

https://projecteuclid.org/euclid.aos/1590480046.

Pollak:2021:RTR

[Pol21] Moshe Pollak. A rule of thumb: Run lengths to false alarm of many types of control charts run in parallel on dependent streams are asymptotically independent. Annals of Statistics, 49 (1):557–567, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-1/
A-rule-of-thumb--Run-lengths-to-false-alarm/10.1214/
20-AOS1968.full.

Perez-Ortiz:2024:SGI

[POLdHG24] Muriel Felipe Pérez-Ortiz, Tyron Lardy, Rianne de Heide, and Peter D. Grünwald. E-statistics, group invariance and anytimevalid testing. Annals of Statistics, 52(4):1410–1432, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-52/issue-4/E-statistics-
group-invariance-and-anytime-valid-testing/10.1214/24-
AOS2394.full.

Padoan:2022:CBI

[PR22] Simone A. Padoan and Stefano Rizzelli. Consistency of Bayesian inference for multivariate max-stable distributions. Annals of Statistics, 50(3):1490–1518, June 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-3/Consistency-of-Bayesian-
inference-for-multivariate-max-stable-distributions/
10.1214/21-AOS2160.full.

Paindaveine:2020:TPC

[PRV20] Davy Paindaveine, Julien Remy, and Thomas Verdebout. Testing for principal component directions under weak identifiability. Annals of Statistics, 48(1):324–345, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

https://projecteuclid.org/euclid.aos/1581930137.

Porter:2020:BHM

[PS20] Thomas Porter and Michael Stewart. Beyond HC: More sensitive tests for rare/weak alternatives. Annals of Statistics, 48(4): 2230–2252, August 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-4/
Beyond-HC-More-sensitive-tests-for-rareweak-alternatives/
10.1214/19-AOS1885.full.

Pananjady:2022:IRU

[PS22] Ashwin Pananjady and Richard J. Samworth. Isotonic regression with unknown permutations: Statistics, computation and adaptation. Annals of Statistics, 50(1):324–350, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-1/Isotonic-regression-
with-unknown-permutations-Statistics-computation-and-
adaptation/10.1214/21-AOS2107.full.

Pilipovic:2024:PEN

[PSD24] Predrag Pilipovic, Adeline Samson, and Susanne Ditlevsen. Parameter estimation in nonlinear multivariate stochastic differential equations based on splitting schemes. Annals of Statistics, 52(2):842–867, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-2/
Parameter-estimation-in-nonlinear-multivariate-stochastic-
differential-equations-based-on/10.1214/24-AOS2371.full.

Panigrahi:2021:IMP

[PTW21] Snigdha Panigrahi, Jonathan Taylor, and Asaf Weinstein. Integrative methods for post-selection inference under convex constraints. Annals of Statistics, 49(5):2803–2824, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-

of-statistics/volume-49/issue-5/Integrative-methods-
for-post-selection-inference-under-convex-constraints/
10.1214/21-AOS2057.full.

Paindaveine:2020:ISL

[PV20a] Davy Paindaveine and Thomas Verdebout. Inference for spherical location under high concentration. Annals of Statistics, 48(5): 2982–2998, October 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-5/
Inference-for-spherical-location-under-high-concentration/
10.1214/19-AOS1918.full.

Patilea:2020:GAC

[PV20b] Valentin Patilea and Ingrid Van Keilegom. A general approach for cure models in survival analysis. Annals of Statistics, 48(4): 2323–2346, August 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-4/
A-general-approach-for-cure-models-in-survival-analysis/
10.1214/19-AOS1889.full.

Pang:2021:CMO

[PWLL21] Shanqi Pang, Jing Wang, Dennis K. J. Lin, and Min-Qian Liu. Construction of mixed orthogonal arrays with high strength. Annals of Statistics, 49(5):2870–2884, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-49/issue-5/Construction-
of-mixed-orthogonal-arrays-with-high-strength/10.1214/
21-AOS2063.full.

Qiu:2023:ANE

[QLY23] Jiaxin Qiu, Zeng Li, and Jianfeng Yao. Asymptotic normality for eigenvalue statistics of a general sample covariance matrix when. Annals of Statistics, 51(3):1427– 1451, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-3/
Asymptotic-normality-for-eigenvalue-statistics-of-a-
general-sample-covariance/10.1214/23-AOS2300.full.

Qiu:2024:EMR

[QTD24] Hongxiang Qiu, Eric Tchetgen Tchetgen, and Edgar Dobriban. Efficient and multiply robust risk estimation under general forms of dataset shift. Annals of Statistics, 52(4):1796–1824, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-52/issue-4/Efficient-and-
multiply-robust-risk-estimation-under-general-forms-
of/10.1214/24-AOS2422.full.

Reeve:2021:ATL

[RCS21] Henry W. J. Reeve, Timothy I. Cannings, and Richard J. Samworth. Adaptive transfer learning. Annals of Statistics, 49(6): 3618–3649, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-6/
Adaptive-transfer-learning/10.1214/21-AOS2102.full.

Reeve:2023:OSS

[RCS23] Henry W. J. Reeve, Timothy I. Cannings, and Richard J. Samworth. Optimal subgroup selection. Annals of Statistics, 51(6): 2342–2365, December 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-6/
Optimal-subgroup-selection/10.1214/23-AOS2328.full.

Rodriguez-Casal:2022:DAM

[RCSN22] Alberto Rodríguez-Casal and Paula Saavedra-Nieves. A data-adaptive method for estimating density level sets under shape conditions. Annals of Statistics, 50(3):1653– 1668, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-3/
A-data-adaptive-method-for-estimating-density-level-
sets-under/10.1214/21-AOS2168.full.

Richardson:2023:NMP

[RERS23] Thomas S. Richardson, Robin J. Evans, James M. Robins, and Ilya Shpitser. Nested Markov properties for acyclic directed mixed graphs. Annals of Statistics, 51(1):334–361, February 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/

annals-of-statistics/volume-51/issue-1/Nested-Markov-
properties-for-acyclic-directed-mixed-graphs/10.1214/
22-AOS2253.full.

Rottger:2023:TPM

[REZ23] Frank Röttger, Sebastian Engelke, and Piotr Zwiernik. Total positivity in multivariate extremes. Annals of Statistics, 51 (3):962–1004, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-3/
Total-positivity-in-multivariate-extremes/10.1214/23-
AOS2272.full.

Rytgaard:2022:CTT

[RGvdL22] Helene C. Rytgaard, Thomas A. Gerds, and Mark J. van der Laan. Continuous-time targeted minimum loss-based estimation of intervention-specific mean outcomes. Annals of Statistics, 50 (5):2469–2491, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-5/
Continuous-time-targeted-minimum-loss-based-estimation-
of-intervention-specific/10.1214/21-AOS2114.full.

Roycraft:2023:BPB

[RKP23] Benjamin Roycraft, Johannes Krebs, and Wolfgang Polonik. Bootstrapping persistent Betti numbers and other stabilizing statistics. Annals of Statistics, 51(4):1484–1509, August 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-51/issue-4/Bootstrapping-
persistent-Betti-numbers-and-other-stabilizing-statistics/
10.1214/23-AOS2277.full.

Rohde:2020:GRC

[RS20a] Angelika Rohde and Lukas Steinberger. Geometrizing rates of convergence under local differential privacy constraints. Annals of Statistics, 48(5):2646–2670, October 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-5/Geometrizing-rates-of-
convergence-under-local-differential-privacy-constraints/
10.1214/19-AOS1901.full.

Rousseau:2020:AFC

[RS20b] Judith Rousseau and Botond Szabo. Asymptotic frequentist coverage properties of Bayesian credible sets for sieve priors. Annals of Statistics, 48(4):2155–2179, August 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-4/Asymptotic-frequentist-
coverage-properties-of-Bayesian-credible-sets-for-sieve/
10.1214/19-AOS1881.full.

Rousseau:2024:WCB

[RS24] Judith Rousseau and Catia Scricciolo. Wasserstein convergence in Bayesian and frequentist deconvolution models. Annals of Statistics, 52(4):1691–1715, August 2024. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-4/Wasserstein-convergence-
in-Bayesian-and-frequentist-deconvolution-models/10.
1214/24-AOS2413.full.

Reiss:2020:NBA

[RSH20] Markus Reiß and Johannes Schmidt-Hieber. Nonparametric Bayesian analysis of the compound Poisson prior for support boundary recovery. Annals of Statistics, 48(3):1432–1451, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-48/issue-3/Nonparametric-
Bayesian-analysis-of-the-compound-Poisson-prior-for-
support/10.1214/19-AOS1853.full.

Roquain:2022:FDR

[RV22] Etienne Roquain and Nicolas Verzelen. False discovery rate control with unknown null distribution: Is it possible to mimic the oracle? Annals of Statistics, 50(2):1095–1123, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-2/False-discovery-rate-
control-with-unknown-null-distribution--Is/10.1214/21-
AOS2141.full.

Rockova:2020:PCB

[RvdP20]  Veronika Rocková and Stéphanie van der Pas.  Poste-

rior concentration for Bayesian regression trees and forests.

Annals of Statistics, 48(4):2108–2131, August 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-4/Posterior-concentration-
for-Bayesian-regression-trees-and-forests/10.1214/19-
AOS1879.full.

Ray:2020:SBC

[RvdV20]  Kolyan Ray and Aad van der Vaart.  Semiparametric

Bayesian causal inference. Annals of Statistics, 48(5):2999– 3020, October 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-5/
Semiparametric-Bayesian-causal-inference/10.1214/19-
AOS1919.full.

Reiss:2020:NUB

[RW20a] Markus Reiß and Martin Wahl. Nonasymptotic upper bounds for the reconstruction error of PCA. Annals of Statistics, 48 (2):1098–1123, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/euclid.aos/1590480047.

Royset:2020:VAC

[RW20b] Johannes O. Royset and Roger J-B Wets. Variational analysis of constrained M-estimators. Annals of Statistics, 48(5): 2759–2790, October 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-5/
Variational-analysis-of-constrained-M-estimators/10.
1214/19-AOS1905.full.

Reiss:2023:IMR

[RW23] Markus Reiss and Lars Winkelmann. Inference on the maximal rank of time-varying covariance matrices using highfrequency data. Annals of Statistics, 51(2):791–815, April 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-51/issue-2/Inference-on-
the-maximal-rank-of-time-varying-covariance-matrices/
10.1214/23-AOS2273.full.

Rao:2021:RGW

[RY21] Suhasini Subba Rao and Junho Yang. Reconciling the Gaussian and Whittle likelihood with an application to estimation in the frequency domain. Annals of Statistics, 49(5): 2774–2802, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-5/
Reconciling-the-Gaussian-and-Whittle-likelihood-with-
an-application-to/10.1214/21-AOS2055.full.

Savje:2021:ATE

[SAH21] Fredrik Sävje, Peter M. Aronow, and Michael G. Hudgens. Average treatment effects in the presence of unknown interference. Annals of Statistics, 49(2):673–701, April 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-2/Average-treatment-effects-
in-the-presence-of-unknown-interference/10.1214/20-AOS1973.
full.

Sell:2024:NCM

[SBC24] Torben Sell, Thomas B. Berrett, and Timothy I. Cannings. Nonparametric classification with missing data. Annals of Statistics, 52(3):1178–1200, June 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-3/
Nonparametric-classification-with-missing-data/10.1214/
24-AOS2389.full.

Silin:2022:CTN

[SF22] Igor Silin and Jianqing Fan. Canonical thresholding for nonsparse high-dimensional linear regression. Annals of Statistics, 50(1): 460–486, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-1/
Canonical-thresholding-for-nonsparse-high-dimensional-
linear-regression/10.1214/21-AOS2116.full.

Song:2024:CAD

[SF24] Yanglei Song and Georgios Fellouris. Change acceleration and detection. Annals of Statistics, 52(3):1050–1075, June 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-

of-statistics/volume-52/issue-3/Change-acceleration-
and-detection/10.1214/24-AOS2382.full.

Saha:2020:NML

[SG20] Sujayam Saha and Adityanand Guntuboyina. On the nonparametric maximum likelihood estimator for Gaussian location mixture densities with application to Gaussian denoising. Annals of Statistics, 48(2):738–762, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https:/

/projecteuclid.org/euclid.aos/1590480032.

Shen:2020:OEV

[SGWH20] Yandi Shen, Chao Gao, Daniela Witten, and Fang Han. Optimal estimation of variance in nonparametric regression with random design. Annals of Statistics, 48(6):3589– 3618, December 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-6/
Optimal-estimation-of-variance-in-nonparametric-regression-
with-random-design/10.1214/20-AOS1944.full.

Schmidt-Hieber:2020:NRU

[SH20a] Johannes Schmidt-Hieber. Nonparametric regression using deep neural networks with ReLU activation function. Annals of Statistics, 48(4):1875–1897, August 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-4/Nonparametric-regression-
using-deep-neural-networks-with-ReLU-activation-function/

10.1214/19-AOS1875.full. See discussion [Kut20] and correction [SHV24].

Shen:2020:HFE

[SH20b] Jinqi Shen and Tailen Hsing. Hurst function estimation. Annals of Statistics, 48(2):838–862, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1590480036.

Shi:2022:UCF

[SHDH22] Hongjian Shi, Marc Hallin, Mathias Drton, and Fang Han. On universally consistent and fully distribution-free rank tests of vector independence. Annals of Statistics, 50(4):1933– 1959, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-4/
On-universally-consistent-and-fully-distribution-free-
rank-tests-of/10.1214/21-AOS2151.full.

Schmidt-Hieber:2021:PAB

[SHSS^{+}21] Johannes Schmidt-Hieber, Laura Fee Schneider, Thomas Staudt, Andrea Krajina, Timo Aspelmeier, and Axel Munk. Posterior analysis of n in the binomial. Annals of Statistics, 49(6): 3534–3558, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-6/
Posterior-analysis-of-n-in-the-binomial-np-problem-with/
10.1214/21-AOS2096.full.

Schmidt-Hieber:2024:CNR

[SHV24] Johannes Schmidt-Hieber and Don Vu. Correction to “Nonparametric regression using deep neural networks with ReLU activation function”. Annals of Statistics, 52(1):413–414, February 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-52/issue-1/Correction-to-
Nonparametric-regression-using-deep-neural-networks-
with-ReLU/10.1214/24-AOS2351.full. See [SH20a].

Spector:2022:PKM

[SJ22] Asher Spector and Lucas Janson. Powerful knockoffs via minimizing reconstructability. Annals of Statistics, 50(1):252– 276, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-1/
Powerful-knockoffs-via-minimizing-reconstructability/
10.1214/21-AOS2104.full.

Steinberger:2023:CPI

[SL23] Lukas Steinberger and Hannes Leeb. Conditional predictive inference for stable algorithms. Annals of Statistics, 51(1): 290–311, February 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-1/
Conditional-predictive-inference-for-stable-algorithms/
10.1214/22-AOS2250.full.

Shao:2022:IRF

[SLY22] Lingxuan Shao, Zhenhua Lin, and Fang Yao. Intrinsic Riemannian functional data analysis for sparse longitudinal observations. Annals of Statistics, 50(3):1696–1721, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-3/Intrinsic-Riemannian-
functional-data-analysis-for-sparse-longitudinal-observations/
10.1214/22-AOS2172.full.

Syring:2020:RRO

[SM20] Nicholas Syring and Ryan Martin. Robust and rate-optimal Gibbs posterior inference on the boundary of a noisy image. Annals of Statistics, 48(3):1498–1513, June 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-3/Robust-and-rate-optimal-
Gibbs-posterior-inference-on-the-boundary/10.1214/19-
AOS1856.full.

Schoen:2023:OAO

[SM23] Eric D. Schoen and Robert W. Mee. Order-of-addition orthogonal arrays to study the effect of treatment ordering. Annals of Statistics, 51(4):1877–1894, August 2023. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-4/Order-of-addition-orthogonal-
arrays-to-study-the-effect-of/10.1214/23-AOS2317.full.

Stoltenberg:2022:CSD

[SMZ22] Emil A. Stoltenberg, Per A. Mykland, and Lan Zhang. A CLT for second difference estimators with an application to volatility and intensity. Annals of Statistics, 50(4):2072– 2095, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-4/
A-CLT-for-second-difference-estimators-with-an-application-
to/10.1214/22-AOS2176.full.

Schell:2023:NIC

[SO23] Alexander Schell and Harald Oberhauser. Nonlinear independent component analysis for discrete-time and continuoustime signals. Annals of Statistics, 51(2):487–518, April 2023.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-2/Nonlinear-independent-
component-analysis-for-discrete-time-and-continuous-
time/10.1214/23-AOS2256.full.

Shah:2020:HCI

[SP20] Rajen D. Shah and Jonas Peters. The hardness of conditional independence testing and the generalised covariance measure. Annals of Statistics, 48(3):1514–1538, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-48/issue-3/The-hardness-
of-conditional-independence-testing-and-the-generalised-
covariance/10.1214/19-AOS1857.full.

Sanders:2020:CBM

[SPY20] Jaron Sanders, Alexandre Proutière, and Se-Young Yun. Clustering in Block Markov Chains. Annals of Statistics, 48(6): 3488–3512, December 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-6/
Clustering-in-Block-Markov-Chains/10.1214/19-AOS1939.
full.

Schweinberger:2020:CCR

[SS20] Michael Schweinberger and Jonathan Stewart. Concentration and consistency results for canonical and curved exponentialfamily models of random graphs. Annals of Statistics, 48(1): 374–396, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/euclid.aos/1581930139.

Sriperumbudur:2022:AKP

[SS22] Bharath K. Sriperumbudur and Nicholas Sterge. Approximate kernel PCA: Computational versus statistical trade-off. Annals of Statistics, 50(5):2713–2736, October 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-5/Approximate-kernel-PCA-
Computational-versus-statistical-trade-off/10.1214/22-
AOS2204.full.

Spencer:2023:PSL

[SS23] Neil A. Spencer and Cosma Rohilla Shalizi. Projective, sparse and learnable latent position network models. Annals of Statistics, 51(6):2506–2525, December 2023. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-6/Projective-sparse-and-
learnable-latent-position-network-models/10.1214/23-
AOS2340.full.

Shi:2021:CVI

[SSL21] Chengchun Shi, Rui Song, and Wenbin Lu. Concordance and value information criteria for optimal treatment decision. Annals of Statistics, 49(1):49–75, February 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-1/Concordance-and-value-
information-criteria-for-optimal-treatment-decision/
10.1214/19-AOS1908.full.

Stankewitz:2024:ES

[Sta24] Bernhard Stankewitz. Early stopping for L^{2}-boosting in high-dimensional linear models. Annals of Statistics, 52(2): 491–518, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-2/
Early-stopping-for-L2-boosting-in-high-dimensional-linear-
models/10.1214/24-AOS2356.full.

Sherlock:2021:EDA

[STG21] Chris Sherlock, Alexandre H. Thiery, and Andrew Golightly. Efficiency of delayed-acceptance random walk Metropolis algorithms. Annals of Statistics, 49(5):2972–2990, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-49/issue-5/Efficiency-of-
delayed-acceptance-random-walk-Metropolis-algorithms/
10.1214/21-AOS2068.full.

Strawderman:2021:CSC

[Str21] William E. Strawderman. On Charles Stein’s contributions to (in)admissibility. Annals of Statistics, 49(4):1823– 1835, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-4/
On-Charles-Steins-contributions-to-inadmissibility/10.
1214/21-AOS2108.full.

Szabo:2023:OHD

[SVvZ23] Botond Szabó, Lasse Vuursteen, and Harry van Zanten. Optimal high-dimensional and nonparametric distributed testing under communication constraints. Annals of Statistics, 51(3):909–934, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-51/issue-3/Optimal-high-
dimensional-and-nonparametric-distributed-testing-under-
communication-constraints/10.1214/23-AOS2269.full.

Szabo:2020:ADM

[SvZ20] Botond Szabó and Harry van Zanten. Adaptive distributed methods under communication constraints. Annals of Statistics, 48 (4):2347–2380, August 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-4/
Adaptive-distributed-methods-under-communication-constraints/
10.1214/19-AOS1890.full.

Shao:2007:AST

[SW07] Xiaofeng Shao and Wei Biao Wu. Asymptotic spectral theory for nonlinear time series. Annals of Statistics, 35(4):1773–1801, August 2007. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1188405630. See correction [ZSW22].

Schramm:2022:CBE

[SW22] Tselil Schramm and Alexander S. Wein. Computational barriers to estimation from low-degree polynomials. Annals of Statistics, 50(3):1833–1858, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-3/
Computational-barriers-to-estimation-from-low-degree-
polynomials/10.1214/22-AOS2179.full.

Schmidt:2021:ATC

[SWFD21] Sara K. Schmidt, Max Wornowizki, Roland Fried, and Herold Dehling. An asymptotic test for constancy of the variance under short-range dependence. Annals of Statistics, 49(6):3460– 3481, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-6/
An-asymptotic-test-for-constancy-of-the-variance-under-
short/10.1214/21-AOS2092.full.

She:2021:AGB

[SWJ21] Yiyuan She, Zhifeng Wang, and Jiuwu Jin. Analysis of generalized Bregman surrogate algorithms for nonsmooth nonconvex statistical learning. Annals of Statistics, 49(6):3434– 3459, December 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-6/
Analysis-of-generalized-Bregman-surrogate-algorithms-
for-nonsmooth-nonconvex-statistical/10.1214/21-AOS2090.
full.

Sourisseau:2022:AAS

[SWZ22] Matt Sourisseau, Hau-Tieng Wu, and Zhou Zhou. Asymptotic analysis of synchrosqueezing transform-toward statistical inference with nonlinear-type time-frequency analysis. Annals of Statistics, 50(5):2694–2712, October 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-5/Asymptotic-analysis-
of-synchrosqueezing-transformtoward-statistical-inference-
with-nonlinear-type/10.1214/22-AOS2203.full.

Soloff:2024:EDC

[SXF24] Jake A. Soloff, Daniel Xiang, and William Fithian. The edge of discovery: Controlling the local false discovery rate at the margin. Annals of Statistics, 52(2):580–601, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-52/issue-2/The-
edge-of-discovery--Controlling-the-local-false-discovery/
10.1214/24-AOS2359.full.

Tan:2020:MAI

[Tan20] Zhiqiang Tan. Model-assisted inference for treatment effects using regularized calibrated estimation with high-dimensional data. Annals of Statistics, 48(2):811–837, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

https://projecteuclid.org/euclid.aos/1590480035.

Telschow:2023:EEE

[TCPS23] Fabian J. E. Telschow, Dan Cheng, Pratyush Pranav, and Armin Schwartzman. Estimation of expected Euler characteristic curves of nonstationary smooth random fields. Annals of Statistics, 51 (5):2272–2297, October 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-5/
Estimation-of-expected-Euler-characteristic-curves-of-
nonstationary-smooth-random/10.1214/23-AOS2337.full.

Tian:2020:PEA

[Tia20] Xiaoying Tian. Prediction error after model search. Annals of Statistics, 48(2):763–784, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https:/

/projecteuclid.org/euclid.aos/1590480033.

Tan:2020:SSO

[TSY20] Kai Tan, Lei Shi, and Zhou Yu. Sparse SIR: Optimal rates and adaptive estimation. Annals of Statistics, 48(1):64–85, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/euclid.aos/
1581930126.

Thepaut:2024:OES

[TV24] Solène Thépaut and Nicolas Verzelen. Optimal estimation of Schatten norms of a rectangular matrix. Annals of Statistics, 52 (4):1334–1359, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-4/
Optimal-estimation-of-Schatten-norms-of-a-rectangular-
matrix/10.1214/24-AOS2374.full.

Tang:2021:WEF

[TX21] Yu Tang and Hongquan Xu. Wordlength enumerator for fractional factorial designs. Annals of Statistics, 49(1):255– 271, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-1/
Wordlength-enumerator-for-fractional-factorial-designs/
10.1214/20-AOS1955.full.

Tang:2023:MRD

[TY23] Rong Tang and Yun Yang. Minimax rate of distribution estimation on unknown submanifolds under adversarial losses. Annals of Statistics, 51(3):1282–1308, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-3/Minimax-rate-of-distribution-
estimation-on-unknown-submanifolds-under-adversarial/
10.1214/23-AOS2291.full.

Tang:2020:PGE

[TYZ20] Niansheng Tang, Xiaodong Yan, and Xingqiu Zhao. Penalized generalized empirical likelihood with a diverging number of general estimating equations for censored data. Annals of Statistics, 48(1):607–627, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https:/

/projecteuclid.org/euclid.aos/1581930149.

vanDelft:2024:GFQ

[vDD24] Anne van Delft and Holger Dette. A general framework to quantify deviations from structural assumptions in the analysis of nonstationary function-valued processes. Annals of Statistics, 52(2):550–579, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-2/
A-general-framework-to-quantify-deviations-from-structural-
assumptions-in/10.1214/24-AOS2358.full.

vandeGeer:2021:WVZ

[vdGK21] Sara van de Geer and Chris A. J. Klaassen. Willem van Zwet, teacher and thesis advisor. Annals of Statistics, 49(5): 2448–2456, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-5/
Willem-van-Zwet-teacher-and-thesis-advisor/10.1214/21-
AOS2069.full.

vanderVaart:2021:SEN

[vdVW21] A. W. van der Vaart and J. A. Wellner. Stein 1956: Efficient nonparametric testing and estimation. Annals of Statistics, 49 (4):1836–1849, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-4/

Stein-1956-Efficient-nonparametric-testing-and-estimation/
10.1214/21-AOS2056.full.

Velasco:2022:ETS

[Vel22] Carlos Velasco. Estimation of time series models using residuals dependence measures. Annals of Statistics, 50(5):3039– 3063, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-5/
Estimation-of-time-series-models-using-residuals-dependence-
measures/10.1214/22-AOS2220.full.

Verzelen:2023:OCP

[VFLRB23] Nicolas Verzelen, Magalie Fromont, Matthieu Lerasle, and Patricia Reynaud-Bouret. Optimal change-point detection and localization. Annals of Statistics, 51(4):1586–1610, August 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-4/Optimal-change-point-
detection-and-localization/10.1214/23-AOS2297.full.

Vovk:2021:VCC

[VW21] Vladimir Vovk and Ruodu Wang. E-values: Calibration, combination and applications. Annals of Statistics, 49(3): 1736–1754, June 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-3/
E-values-Calibration-combination-and-applications/10.
1214/20-AOS2020.full.

Vovk:2022:AWM

[VWW22] Vladimir Vovk, Bin Wang, and Ruodu Wang. Admissible ways of merging p-values under arbitrary dependence. Annals of Statistics, 50(1):351–375, February 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-1/Admissible-ways-of-merging-
p-values-under-arbitrary-dependence/10.1214/21-AOS2109.
full.

Westling:2020:USN

[WC20] Ted Westling and Marco Carone. A unified study of nonparametric inference for monotone functions. Annals of Statistics, 48 (2):1001–1024, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/euclid.aos/1590480043.

Wang:2023:CCI

[WG23] Kang Wang and Subhashis Ghosal. Coverage of credible intervals in Bayesian multivariate isotonic regression. Annals of Statistics, 51(3):1376–1400, June 2023. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-3/Coverage-of-credible-
intervals-in-Bayesian-multivariate-isotonic-regression/
10.1214/23-AOS2298.full.

Wang:2022:DMR

[WGS^{+}22] Jingshu Wang, Lin Gui, Weijie J. Su, Chiara Sabatti, and Art B. Owen. Detecting multiple replicating signals using adaptive filtering procedures. Annals of Statistics, 50(4): 1890–1909, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-4/
Detecting-multiple-replicating-signals-using-adaptive-
filtering-procedures/10.1214/21-AOS2139.full.

Wang:2022:RDC

[WL22] Yuhao Wang and Xinran Li. Rerandomization with diminishing covariate imbalance and diverging number of covariates. Annals of Statistics, 50(6):3439–3465, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-6/Rerandomization-with-
diminishing-covariate-imbalance-and-diverging-number-
of-covariates/10.1214/22-AOS2235.full.

Wang:2021:TLP

[WM21] Chunyan Wang and Robert W. Mee. Two-level parallel flats designs. Annals of Statistics, 49(5):3015–3042, October 2021. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-49/issue-5/Two-
level-parallel-flats-designs/10.1214/21-AOS2071.full.

Wei:2022:CFM

[WN22] Yun Wei and XuanLong Nguyen. Convergence of de Finetti’s mixing measure in latent structure models for observed exchangeable sequences. Annals of Statistics, 50(4):1859– 1889, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-4/
Convergence-of-de-Finettis-mixing-measure-in-latent-
structure-models/10.1214/21-AOS2120.full.

Waghmare:2022:CCK

[WP22] Kartik G. Waghmare and Victor M. Panaretos. The completion of covariance kernels. Annals of Statistics, 50(6):3281–3306, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-50/issue-6/The-
completion-of-covariance-kernels/10.1214/22-AOS2228.
full.

Wang:2023:PSA

[WQW23] Jiayi Wang, Zhengling Qi, and Raymond K. W. Wong. Projected state-action balancing weights for offline reinforcement learning. Annals of Statistics, 51(4):1639–1665, August 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-51/issue-4/Projected-state-action-
balancing-weights-for-offline-reinforcement-learning/
10.1214/23-AOS2302.full.

Wang:2020:HTH

[WS20] Runmin Wang and Xiaofeng Shao. Hypothesis testing for high-dimensional time series via self-normalization. Annals of Statistics, 48(5):2728–2758, October 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-5/Hypothesis-testing-for-
high-dimensional-time-series-via-self-normalization/
10.1214/19-AOS1904.full.

Weinstein:2023:PAM

[WSB^{+}23] Asaf Weinstein, Weijie J. Su, Małgorzata Bogdan, Rina Foygel Barber, and Emmanuel J. Candès. A power analysis for model-X knockoffs with ℓ_{p}-regularized statistics. Annals of Statistics, 51 (3):1005–1029, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-3/
A-power-analysis-for-model-X-knockoffs-with-%e2%84%93p-
regularized/10.1214/23-AOS2274.full.

Wang:2023:ROR

[WT23] Di Wang and Ruey S. Tsay. Rate-optimal robust estimation of high-dimensional vector autoregressive models. Annals of Statistics, 51(2):846–877, April 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
51/issue-2/Rate-optimal-robust-estimation-of-high-dimensional-
vector-autoregressive-models/10.1214/23-AOS2278.full.

Wang:2020:WBE

[WWM20] Shuaiwen Wang, Haolei Weng, and Arian Maleki. Which bridge estimator is the best for variable selection? Annals of Statistics, 48(5):2791–2823, October 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-5/Which-bridge-estimator-
is-the-best-for-variable-selection/10.1214/19-AOS1906.
full.

Wang:2024:SHF

[WWZ^{+}24] Wen Wang, Shihao Wu, Ziwei Zhu, Ling Zhou, and Peter X.- K. Song. Supervised homogeneity fusion: a combinatorial approach. Annals of Statistics, 52(1):285–310, February 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-1/Supervised-homogeneity-
fusion-A-combinatorial-approach/10.1214/23-AOS2347.full.

Wu:2020:OEG

[WY20] Yihong Wu and Pengkun Yang. Optimal estimation of Gaussian mixtures via denoised method of moments. Annals of Statistics, 48(4):1981–2007, August 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-48/issue-4/Optimal-estimation-of-
Gaussian-mixtures-via-denoised-method-of-moments/10.
1214/19-AOS1873.full.

Wang:2021:OCP

[WYR21] Daren Wang, Yi Yu, and Alessandro Rinaldo. Optimal change point detection and localization in sparse dynamic networks. Annals of Statistics, 49(1):203–232, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-1/Optimal-change-point-
detection-and-localization-in-sparse-dynamic-networks/
10.1214/20-AOS1953.full.

Wong:2022:SLC

[WZL22] Kin Yau Wong, Donglin Zeng, and D. Y. Lin. Semiparametric latent-class models for multivariate longitudinal and survival data. Annals of Statistics, 50(1):487– 510, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-1/
Semiparametric-latent-class-models-for-multivariate-
longitudinal-and-survival-data/10.1214/21-AOS2117.full.

Wang:2022:ICP

[WZVS22] Runmin Wang, Changbo Zhu, Stanislav Volgushev, and Xiaofeng Shao. Inference for change points in high-dimensional data via selfnormalization. Annals of Statistics, 50(2):781– 806, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-2/
Inference-for-change-points-in-high-dimensional-data-
via-selfnormalization/10.1214/21-AOS2127.full.

Xu:2020:ORC

[XJL20] Min Xu, Varun Jog, and Po-Ling Loh. Optimal rates for community estimation in the weighted stochastic block model. Annals of Statistics, 48(1):183–204, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/euclid.aos/1581930131.

Xu:2023:IIM

[XM23] Yuanzhe Xu and Sumit Mukherjee. Inference in Ising models on dense regular graphs. Annals of Statistics, 51(3): 1183–1206, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-3/
Inference-in-Ising-models-on-dense-regular-graphs/10.
1214/23-AOS2286.full.

Xu:2021:HDN

[XS21] Min Xu and Richard J. Samworth. High-dimensional nonparametric density estimation via symmetry and shape constraints. Annals of Statistics, 49(2):650–672, April 2021. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-2/High-dimensional-nonparametric-
density-estimation-via-symmetry-and-shape-constraints/
10.1214/20-AOS1972.full.

Xu:2024:CPI

[XWZY24] Haotian Xu, Daren Wang, Zifeng Zhao, and Yi Yu. Changepoint inference in high-dimensional regression models under temporal dependence. Annals of Statistics, 52(3):999–1026, June 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-52/issue-3/Change-point-
inference-in-high-dimensional-regression-models-under-
temporal/10.1214/24-AOS2380.full.

Xue:2020:DCF

[XY20] Kaijie Xue and Fang Yao. Distribution and correlation-free twosample test of high-dimensional means. Annals of Statistics, 48 (3):1304–1328, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-3/
Distribution-and-correlation-free-two-sample-test-of-
high-dimensional/10.1214/19-AOS1848.full.

Xi:2020:CEE

[XYY20] Haokai Xi, Fan Yang, and Jun Yin. Convergence of eigenvector empirical spectral distribution of sample covariance matrices. Annals of Statistics, 48(2):953–982, April 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

https://projecteuclid.org/euclid.aos/1590480041.

Xia:2021:SOC

[XYZ21] Dong Xia, Ming Yuan, and Cun-Hui Zhang. Statistically optimal and computationally efficient low rank tensor completion from noisy entries. Annals of Statistics, 49(1):76– 99, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-1/
Statistically-optimal-and-computationally-efficient-
low-rank-tensor-completion-from/10.1214/20-AOS1942.full.

Xia:2022:ILR

[XZZ22] Dong Xia, Anru R. Zhang, and Yuchen Zhou. Inference for low-rank tensors — no need to debias. Annals of Statistics, 50 (2):1220–1245, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-2/
Inference-for-low-rank-tensorsno-need-to-debias/10.1214/
21-AOS2146.full.

Yan:2024:IHP

[YCF24] Yuling Yan, Yuxin Chen, and Jianqing Fan. Inference for heteroskedastic PCA with missing data. Annals of Statistics, 52 (2):729–756, April 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-2/
Inference-for-heteroskedastic-PCA-with-missing-data/
10.1214/24-AOS2366.full.

Yu:2024:BEL

[YKN24] Haihan Yu, Mark S. Kaiser, and Daniel J. Nordman. A blockwise empirical likelihood method for time series in frequency domain inference. Annals of Statistics, 52(3):1152– 1177, June 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-3/
A-blockwise-empirical-likelihood-method-for-time-series-
in-frequency/10.1214/24-AOS2388.full.

Yuan:2022:TCS

[YLFS22] Mingao Yuan, Ruiqi Liu, Yang Feng, and Zuofeng Shang. Testing community structure for hypergraphs. Annals of Statistics, 50 (1):147–169, February 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-1/
Testing-community-structure-for-hypergraphs/10.1214/
21-AOS2099.full.

Yadlowsky:2022:BCA

[YNB^{+}22] Steve Yadlowsky, Hongseok Namkoong, Sanjay Basu, John Duchi, and Lu Tian. Bounds on the conditional and average treatment effect with unobserved confounding factors. Annals of Statistics, 50(5):2587–2615, October 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-5/Bounds-on-the-conditional-
and-average-treatment-effect-with-unobserved/10.1214/
22-AOS2195.full.

Yuan:2021:CDD

[YQ21] Yubai Yuan and Annie Qu. Community detection with dependent connectivity. Annals of Statistics, 49(4):2378– 2428, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-4/
Community-detection-with-dependent-connectivity/10.1214/
20-AOS2042.full.

Ye:2021:DIV

[YSK21] Ting Ye, Jun Shao, and Hyunseung Kang. Debiased inversevariance weighted estimator in two-sample summary-data Mendelian randomization. Annals of Statistics, 49(4):2079– 2100, August 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-4/
Debiased-inverse-variance-weighted-estimator-in-two-
sample-summary-data/10.1214/20-AOS2027.full.

Yan:2024:LGM

[YWR24] Yuling Yan, Kaizheng Wang, and Philippe Rigollet. Learning Gaussian mixtures using the Wasserstein–Fisher–Rao gradient flow. Annals of Statistics, 52(4):1774–1795, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-4/Learning-Gaussian-mixtures-
using-the-WassersteinFisherRao-gradient-flow/10.1214/
24-AOS2416.full.

Yang:2022:TTU

[YZZ22] Wenhao Yang, Liangyu Zhang, and Zhihua Zhang. Toward theoretical understandings of robust Markov decision processes: Sample complexity and asymptotics. Annals of Statistics, 50(6): 3223–3248, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-6/
Toward-theoretical-understandings-of-robust-Markov-decision-
processes--Sample/10.1214/22-AOS2225.full.

Zhang:2023:ACS

[ZA23] Linfan Zhang and Arash A. Amini. Adjusted chi-square test for degree-corrected block models. Annals of Statistics, 51(6): 2366–2385, December 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-6/
Adjusted-chi-square-test-for-degree-corrected-block-
models/10.1214/23-AOS2329.full.

Zhou:2023:CAB

[ZC23] Quan Zhou and Hyunwoong Chang. Complexity analysis of Bayesian learning of high-dimensional DAG models and their equivalence classes. Annals of Statistics, 51(3): 1058–1085, June 2023. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-51/issue-3/
Complexity-analysis-of-Bayesian-learning-of-high-dimensional-
DAG-models/10.1214/23-AOS2280.full.

Zhang:2022:HPA

[ZCW22] Anru R. Zhang, T. Tony Cai, and Yihong Wu. Heteroskedastic PCA: Algorithm, optimality, and applications. Annals of Statistics, 50(1):53–80, February 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-1/Heteroskedastic-PCA-
Algorithm-optimality-and-applications/10.1214/21-AOS2074.
full.

Zhao:2022:RDB

[ZD22] Anqi Zhao and Peng Ding. Reconciling design-based and modelbased causal inferences for split-plot experiments. Annals of Statistics, 50(2):1170–1192, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://

projecteuclid.org/journals/annals-of-statistics/volume-
50/issue-2/Reconciling-design-based-and-model-based-
causal-inferences-for-split/10.1214/21-AOS2144.full.

Zrnic:2024:LSI

[ZF24] Tijana Zrnic and William Fithian. Locally simultaneous inference. Annals of Statistics, 52(3):1227–1253, June 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-52/issue-3/Locally-simultaneous-
inference/10.1214/24-AOS2391.full.

Zhang:2020:CRV

[ZG20] Fengshuo Zhang and Chao Gao. Convergence rates of variational posterior distributions. Annals of Statistics, 48(4): 2180–2207, August 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-4/
Convergence-rates-of-variational-posterior-distributions/
10.1214/19-AOS1883.full.

Zhang:2024:GWD

[ZGMS24] Zhengxin Zhang, Ziv Goldfeld, Youssef Mroueh, and Bharath K. Sriperumbudur. Gromov–Wasserstein distances: Entropic regularization, duality and sample complexity. Annals of Statistics, 52 (4):1616–1645, August 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-4/
GromovWasserstein-distances-Entropic-regularization-
duality-and-sample-complexity/10.1214/24-AOS2406.full.

Zhao:2022:PCA

[ZGV22] Yue Zhao, Irène Gijbels, and Ingrid Van Keilegom. Parametric copula adjusted for non- and semiparametric regression. Annals of Statistics, 50(2):754–780, April 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-2/Parametric-copula-adjusted-
for-non--and-semiparametric-regression/10.1214/21-AOS2126.
full.

Zhilova:2020:NBE

[Zhi20] Mayya Zhilova. Nonclassical Berry–Esseen inequalities and accuracy of the bootstrap. Annals of Statistics, 48(4):1922–1939, August 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-48/issue-4/Nonclassical-

BerryEsseen-inequalities-and-accuracy-of-the-bootstrap/
10.1214/18-AOS1802.full.

Zhilova:2022:NET

[Zhi22] Mayya Zhilova. New Edgeworth-type expansions with finite sample guarantees. Annals of Statistics, 50(5):2545–2561, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-50/issue-5/New-Edgeworth-
type-expansions-with-finite-sample-guarantees/10.1214/
22-AOS2192.full.

Zrnic:2023:PSI

[ZJ23] Tijana Zrnic and Michael I. Jordan. Post-selection inference via algorithmic stability. Annals of Statistics, 51(4):1666–1691, August 2023. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-51/issue-4/Post-
selection-inference-via-algorithmic-stability/10.1214/
23-AOS2303.full.

Zhu:2022:GRR

[ZJS22] Banghua Zhu, Jiantao Jiao, and Jacob Steinhardt. Generalized resilience and robust statistics. Annals of Statistics, 50(4): 2256–2283, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-4/
Generalized-resilience-and-robust-statistics/10.1214/
22-AOS2186.full.

Zhang:2024:SCV

[ZL24] Lu Zhang and Junwei Lu. StarTrek: Combinatorial variable selection with false discovery rate control. Annals of Statistics, 52 (1):78–102, February 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-52/issue-1/
StarTrek-Combinatorial-variable-selection-with-false-
discovery-rate-control/10.1214/23-AOS2296.full.

Zhang:2022:OBR

[ZLN22] Qihao Zhang, Soumendra N. Lahiri, and Daniel J. Nordman. On optimal block resampling for Gaussian-subordinated longrange dependent processes. Annals of Statistics, 50(6):3619– 3646, December 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-6/
On-optimal-block-resampling-for-Gaussian-subordinated-
long-range-dependent/10.1214/22-AOS2242.full.

Zhao:2022:EFM

[ZLSK22] Anqi Zhao, Youjin Lee, Dylan S. Small, and Bikram Karmakar. Evidence factors from multiple, possibly invalid, instrumental variables. Annals of Statistics, 50(3):1266– 1296, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-3/
Evidence-factors-from-multiple-possibly-invalid-instrumental-
variables/10.1214/21-AOS2148.full.

Zhong:2022:DLP

[ZMW22] Qixian Zhong, Jonas Mueller, and Jane-Ling Wang. Deep learning for the partially linear Cox model. Annals of Statistics, 50 (3):1348–1375, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-3/
Deep-learning-for-the-partially-linear-Cox-model/10.
1214/21-AOS2153.full.

Zhang:2022:RRR

[ZP22] Yunyi Zhang and Dimitris N. Politis. Ridge regression revisited: Debiasing, thresholding and bootstrap. Annals of Statistics, 50 (3):1401–1422, June 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-3/
Ridge-regression-revisited-Debiasing-thresholding-and-
bootstrap/10.1214/21-AOS2156.full.

Zhu:2020:EBS

[ZS20] Guangyu Zhu and Zhihua Su. Envelope-based sparse partial least squares. Annals of Statistics, 48(1):161–182, February 2020. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL https://projecteuclid.org/euclid.aos/1581930130.

Zhang:2022:SSG

[ZS22] Tianyu Zhang and Noah Simon. A sieve stochastic gradient descent estimator for online nonparametric regression in Sobolev ellipsoids. Annals of Statistics, 50(5):2848– 2871, October 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-50/issue-5/
A-sieve-stochastic-gradient-descent-estimator-for-online-
nonparametric-regression/10.1214/22-AOS2212.full.

Zhang:2022:CNS

[ZSW22] Yi Zhang, Xiaofeng Shao, and Weibiao Wu. Correction note: Asymptotic spectral theory for nonlinear time series. Annals of Statistics, 50(5):3088–3089, October 2022. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-5/Correction-note-Asymptotic-
spectral-theory-for-nonlinear-time-series/10.1214/22-
AOS2206.full. See [SW07].

Zou:2021:MBS

[ZVB21] Nan Zou, Stanislav Volgushev, and Axel Bücher. Multiple block sizes and overlapping blocks for multivariate time series extremes. Annals of Statistics, 49(1):295–320, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-49/issue-1/Multiple-block-sizes-
and-overlapping-blocks-for-multivariate-time-series/
10.1214/20-AOS1957.full.

Zhang:2021:CCS

[ZW21] Danna Zhang and Wei Biao Wu. Convergence of covariance and spectral density estimates for high-dimensional locally stationary processes. Annals of Statistics, 49(1):233– 254, February 2021. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-49/issue-1/
Convergence-of-covariance-and-spectral-density-estimates-
for-high-dimensional/10.1214/20-AOS1954.full.

Zhao:2020:SIT

[ZWH^{+}20] Qingyuan Zhao, Jingshu Wang, Gibran Hemani, Jack Bowden, and Dylan S. Small. Statistical inference in twosample summary-data Mendelian randomization using robust adjusted profile score. Annals of Statistics, 48(3): 1742–1769, June 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-3/
Statistical-inference-in-two-sample-summary-data-Mendelian-
randomization-using/10.1214/19-AOS1866.full.

Zou:2020:CSN

[ZWL20] Changliang Zou, Guanghui Wang, and Runze Li. Consistent selection of the number of change-points via sample-splitting. Annals of Statistics, 48(1):413–439, February 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

https://projecteuclid.org/euclid.aos/1581930141.

Zhang:2022:EEN

[ZX22] Yuan Zhang and Dong Xia. Edgeworth expansions for network moments. Annals of Statistics, 50(2):726–753, April 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-2/Edgeworth-expansions-
for-network-moments/10.1214/21-AOS2125.full.

Zhou:2024:RBI

[ZXZL24] Yeqing Zhou, Kai Xu, Liping Zhu, and Runze Li. Rankbased indices for testing independence between two highdimensional vectors. Annals of Statistics, 52(1):184–206, February 2024. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL https://projecteuclid.org/
journals/annals-of-statistics/volume-52/issue-1/Rank-
based-indices-for-testing-independence-between-two-high-
dimensional/10.1214/23-AOS2339.full.

Zhang:2020:TCG

[ZZ20] Anderson Y. Zhang and Harrison H. Zhou. Theoretical and computational guarantees of mean field variational inference for community detection. Annals of Statistics, 48(5):2575– 2598, October 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-5/
Theoretical-and-computational-guarantees-of-mean-field-
variational-inference-for/10.1214/19-AOS1898.full.

Zhang:2022:AIS

[ZZPZ22] Zhixiang Zhang, Shurong Zheng, Guangming Pan, and Ping-Shou Zhong. Asymptotic independence of spiked eigenvalues and linear spectral statistics for large sample covariance matrices. Annals of Statistics, 50(4):2205–2230, August 2022. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL https://projecteuclid.org/journals/annals-
of-statistics/volume-50/issue-4/Asymptotic-independence-
of-spiked-eigenvalues-and-linear-spectral-statistics-
for/10.1214/22-AOS2183.full.

Zhu:2020:DBR

[ZZYS20] Changbo Zhu, Xianyang Zhang, Shun Yao, and Xiaofeng Shao. Distance-based and RKHS-based dependence metrics in high dimension. Annals of Statistics, 48(6):3366– 3394, December 2020. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL https://projecteuclid.

org/journals/annals-of-statistics/volume-48/issue-6/
Distance-based-and-RKHS-based-dependence-metrics-in-
high-dimension/10.1214/19-AOS1934.full.

Zhao:2024:BRP

[ZZZ24] Bingxin Zhao, Shurong Zheng, and Hongtu Zhu. On blockwise and reference panel-based estimators for genetic data prediction in high dimensions. Annals of Statistics, 52(3):948–965, June 2024. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL https://projecteuclid.org/journals/
annals-of-statistics/volume-52/issue-3/On-blockwise-
and-reference-panel-based-estimators-for-genetic-data/
10.1214/24-AOS2378.full.